### v1

# Aggregation Models with Optimal Weights for Distributed Gaussian Processes

Abstract
GP models have received increasingly attentions in recent years due to their superb prediction accuracy and modeling flexibility. To address the computational burdens of GP models for large-scale datasets, distributed learning for GPs are often adopted. Current aggregation models for distributed GPs are not time-efficient when incorporating correlations between GP experts. In this work, we propose a novel approach for aggregated prediction in distributed GPs. The technique is suitable for both the exact and sparse variational GPs. The proposed method incorporates correlations among experts, leading to better prediction accuracy with manageable computational requirements. As demonstrated by empirical studies, the proposed approach results in more stable predictions in less time than state-of-the-art consistent aggregation models.
1 Introduction
Gaussian processes (GPs) (Rasmussen,, 2003) are a powerful tool for modeling and inference in various areas of machine learning, such as regression (O’Hagan,, 1978; Bishop et al.,, 1995; MacKay et al.,, 2003), classification (Kuss et al.,, 2005; Nickisch and Rasmussen,, 2008; Hensman et al.,, 2015), forecasting (Girard et al.,, 2002; Roberts et al.,, 2013), signal processing (Liutkus et al.,, 2011; Sarkka et al.,, 2013), Bayesian optimization (Snoek et al.,, 2012; Frazier,, 2018), and robotics and control (Deisenroth et al.,, 2013; Mukadam et al.,, 2016). Despite their strengths, GPs face significant challenges when applied to large-scale datasets due to the computational burden of inverting large covariance matrices.
To address these challenges, numerous efficient algorithms have been proposed. Sparse approximations using $m$ inducing points (Quinonero-Candela and Rasmussen,, 2005; Titsias,, 2009; Hensman et al.,, 2013) can reduce the computational complexity from $\mathcal{O}(n^{3})$ to $\mathcal{O}(m^{2}n)$ for a dataset of size $n$. Wilson and Nickisch, (2015) employed structured kernel interpolation combined with Kronecker and Toeplitz algebra to handle a large number of inducing points. Hartikainen and Särkkä, (2010); Grigorievskiy et al., (2017); Nickisch et al., (2018) reformulated GP regression as linear-Gaussian state space models, solving them with classical Kalman filtering theory. Katzfuss and Guinness, (2021); Katzfuss et al., (2022) presented a general Vecchia framework for GP predictions linear computational complexity in the total number of datasets. Chen et al., (2022); Chen and Tuo, (2022); Ding et al., (2024) represented the kernels via a sparse linear transformation to help reduce complexity. However, those approximation methods require at least $\mathcal{O}(n)$ time, which makes them impractical for very large $n$.
Another direction to address the computational issue is distributed learning, which involves distributing computations across multiple units, often referred to as nodes or experts. In distributed learning (Dean et al.,, 2012), the dataset is partitioned and processed in parallel, with results aggregated to form the final model. By applying the previously mentioned GP approximation methods in distributed learning setup across $M$ units, the time complexity can be reduced to $\mathcal{O}(n/M)$. Recent aggregation models for distributed GPs (Tresp,, 2000; Hinton,, 2002; Cao and Fleet,, 2014; Deisenroth and Ng,, 2015) rely on independence assumptions and cannot offer consistent predictions, meaning the aggregated predictive distribution does not converge to the true underlying predictive distribution as the training size $n$ increases to infinity. To overcome this inconsistency, Rullière et al., (2018) introduced the nested point-wise aggregation of experts (NPAE), which incorporates covariances between the experts’ predictions but is very time-consuming. Liu et al., (2018) proposed generalized robust Bayesian committee machine (grBCM), an efficient and consistent algorithm for distributed GPs.
In this work, we introduce a novel aggregation model for distributed GP to improve time complexity. Our methodology is based on GP regressions using the optimized combination technique (OptiCom) introduced in Section 3. The proposed method considers the correlations between the experts to ensure the consistency. It is also less time-consuming than NPAE and grBCM when the number of GP experts $M$ is not very large. Empirical studies show that the proposed algorithm outperforms state-of-the-art aggregation models in computational efficiency while providing stable and consistent predictions.
The remainder of this paper is structured as follows: Section 2 reviews the background and related work on GPs and distributed GPs. Section 3 presents the algorithm for extending OptiCom to GP regressions. Section 4 details our proposed algorithm for distributed GPs. Section 5 presents experimental results and comparisons. Finally, Section 6 concludes the paper with discussions.
2 Background
This section presents the background of this paper. We begin by introducing GP regression, GP training and sparse variational GP in Section 2.1. Subsequently, we outline the methods of the training and aggregated prediction for distributed GPs in Section 2.2.
2.1 GPs
2.1.1 GP regression
A Gaussian process (GP) is a collection of random variables, any finite number of which has a multivariate normal distribution. A GP is completely defined by a mean function $\mu(\cdot)$ and a covariance (or kernel) function $k(\cdot,\cdot^{\prime})$:
$f(\cdot)\sim\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime})),$ (1)
where $\mu(\cdot)=\mathbb{E}[f(\cdot)]$ and $k(\cdot,\cdot^{\prime})=\mathbb{E}[(f(\cdot)-\mu(\cdot))(f(\cdot^{\prime})-\mu(\cdot^{\prime}))]=\mathrm{cov}(f(\cdot),f(\cdot^{\prime}))$.
Suppose we have a set of training points $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ and the observations $\mathbf{y}=(y_{1},\ldots,y_{n})^{\top}$ where $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ with the i.i.d. noise $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a latent function. The GP regression imposes a GP prior over the latent function as $f(\cdot)\sim\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime}))$. By the definition of GP, we would obtain a $n$-dimensional multivariate Gaussian random variable $\mathbf{f}$ by evaluating the GP $f(\cdot)$ at the set of points $\mathbf{X}=\{\mathbf{x}\}_{i=1}^{n}$,
$\mathbf{f}\sim\mathcal{N}(\bm{\mu}_{\mathbf{f}},\mathbf{K}_{\mathbf{f}\mathbf{f}}),$ (2)
where $\mathbf{f}=\big{(}f(\mathbf{x}_{1}),\cdots,f(\mathbf{x}_{n})\big{)}^{\top}$, $\bm{\mu}_{\mathbf{f}}=\big{(}\mu(\mathbf{x}_{1}),\cdots,\mu(\mathbf{x}_{n})\big{)}^{\top}$ and $\mathbf{K}_{\mathbf{f}\mathbf{f}}=\big{[}k(\mathbf{x}_{i},\mathbf{x}_{j})\big{]}_{i,j=1}^{n}$ are the mean and covariance matrix obtained by evaluating at the location points $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$.
In this work, we assume $f$ is a zero-mean GP, i.e., $\mu(\cdot)=0$, then the posterior distribution takes the form
$p(f|\mathbf{y})=\mathcal{N}\big{(}\mu_{f|\mathbf{y}},k_{f|\mathbf{y}}\big{)},$ (3)
with
$\mu_{f|\mathbf{y}}(\cdot)=\mathbf{K}_{(\cdot)\mathbf{f}}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\mathbf{y},$ (4a)
$k_{f|\mathbf{y}}(\cdot,\cdot^{\prime})=k(\cdot,\cdot^{\prime})-\mathbf{K}_{(\cdot)\mathbf{f}}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\mathbf{K}_{\mathbf{f}(\cdot^{\prime})},$ (4b)
where $\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}=\mathbf{K}_{\mathbf{f}\mathbf{f}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n}$ is the covariance matrix of all data points $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ with diagonal observation noise, $\sigma_{\epsilon}^{2}>0$ is the noise variance, $\mathbf{I}_{n}$ is a $n\times n$ identity matrix, $\mathbf{K}_{(\cdot)\mathbf{f}}=k(\cdot,\mathbf{X})=\big{[}k(\cdot,\mathbf{x}_{i})\big{]}_{i=1}^{i=n}=k(\mathbf{X},\cdot)^{\top}=\mathbf{K}_{\mathbf{f}(\cdot)}^{\top}$ is the cross-covariance matrix.
2.1.2 GP training
The hyperparameters $\bm{\theta}$ of GPs, which may include GP variance, kernel lengthscale, are commonly learned by maximizing the log-marginal likelihood (Jones et al.,, 1998) given by
$\mathcal{L}(\bm{\theta})$ $=\log p(\mathbf{y}|\mathbf{f};\bm{\theta})=\log\mathcal{N}(\mathbf{y}|\mathbf{f},\sigma_{\epsilon}^{2}\mathbf{I}_{n})$ (5)
$=-\frac{1}{2}\Big{(}\mathbf{y}^{\top}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\mathbf{y}+\log\det(\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}})+n\log(2\pi)\Big{)}.$
A standard way of obtaining the optimized hyperparameters is by taking the derivative of each $\theta\in\bm{\theta}$ as follows:
$\frac{\partial}{\partial\theta}\mathcal{L}(\bm{\theta})=\frac{1}{2}\mathbf{y}^{\top}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\frac{\partial\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}}{\partial\theta}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\mathbf{y}-\frac{1}{2}\text{tr}\Big{(}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\frac{\partial\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}}{\partial\theta}\Big{)}.$ (6)
2.1.3 Sparse variational GP (SVGP)
Exact GP regression and training requires $\mathcal{O}(n^{3})$ time complexity due to the computation of the terms $\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}$, $\log\det(\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1})$, and $\text{tr}\Big{(}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\frac{\partial\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}}{\partial\theta}\Big{)}$ in eqs. 4, 5 and 6, which is prohibitive when $n$ is large. To solve this problem, Titsias, (2009) introduced sparse variational Gaussian process (SVGP), a variational framework for sparse GPs that approximates $n$ actual observations with $m$ inducing variables ($m\ll n$) by maximizing a lower bound of the true log marginal likelihood, which can reduce the time complexity of GP regression and training to $\mathcal{O}(nm^{2})$.
SVGP consists of $m$ inducing variables $\mathbf{u}=(u_{1},\ldots,u_{m})^{\top}\in\mathbb{R}^{m}$. These latent variables are values of the GP, corresponding to $m$ inducing inputs $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$, $\mathbf{z}_{i}\in\mathbb{R}^{d}$. We specify a Gaussian distribution over the inducing variables $q(\mathbf{u})=\mathcal{N}(\mathbf{m}_{\mathbf{u}},\mathbf{S}_{\mathbf{u}\mathbf{u}})$, then integrating out $\mathbf{u}$ with $q(\mathbf{u})$ yields
[…]
where
$\mu_{f}^{\mathbf{u}}(\cdot)=\mathbf{K}_{(\cdot)\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{m}_{\mathbf{u}},$ (8a)
$k_{f}^{\mathbf{u}}(\cdot,\cdot^{\prime})=k(\cdot,\cdot^{\prime})-\mathbf{K}_{(\cdot)\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}(\mathbf{K}_{\mathbf{u}\mathbf{u}}-\mathbf{S}_{\mathbf{u}\mathbf{u}})\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}(\cdot^{\prime})}.$ (8b)
Here $\mathbf{K}_{(\cdot)\mathbf{u}}=\big{[}k(\cdot,\mathbf{z}_{i})\big{]}_{i=1}^{m}=\mathbf{K}_{\mathbf{u}(\cdot)}^{\top}$, and $\mathbf{K}_{\mathbf{u}\mathbf{u}}=\big{[}k(\mathbf{z}_{i},\mathbf{z}_{j})\big{]}_{i,j=1}^{m}$. The goal is to approximate the exact posterior $p(\mathbf{f},\mathbf{u}|\mathbf{y})$ by an variational distribution $q(\mathbf{f},\mathbf{u})$, so we minimize the Kullback-Leibler (KL) divergence $\text{KL}\big{[}q(\mathbf{f},\mathbf{u})\|p(\mathbf{f},\mathbf{u}|\mathbf{y})\big{]}$, which is equivalent to maximize the evidence lower bound (ELBO) defined as follows:
$\mathcal{L}_{f}:=\mathbb{E}_{q(\mathbf{f},\mathbf{u})}\left[\log\frac{p(\mathbf{y},\mathbf{f},\mathbf{u})}{q(\mathbf{f},\mathbf{u})}\right]=\log p(\mathbf{y})-\text{KL}\big{[}q(\mathbf{f},\mathbf{u})\|p(\mathbf{f},\mathbf{u}|\mathbf{y})\big{]}.$ (9)
Under a Gaussian likelihood of the form $p(\mathbf{y}|\mathbf{f})=\mathcal{N}(\mathbf{y}|\mathbf{f},\sigma_{\epsilon}^{2}\mathbf{I}_{n})$, the optimal variational distribution $\hat{q}(\mathbf{u})=\mathcal{N}(\hat{\mathbf{m}}_{\mathbf{u}},\hat{\mathbf{S}}_{\mathbf{u}\mathbf{u}})$ can be obtained by
$\hat{\mathbf{m}}_{\mathbf{u}}=\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{u}}\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}}\mathbf{y},$ (10a)
$\hat{\mathbf{S}}_{\mathbf{u}\mathbf{u}}=\mathbf{K}_{\mathbf{u}\mathbf{u}}\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{u}},$ (10b)
where $\mathbf{K}_{\mathbf{u}\mathbf{f}}=\big{[}k(\mathbf{z}_{i},\mathbf{x}_{j})\big{]}_{i,j=1}^{m,n}=\mathbf{K}_{\mathbf{f}\mathbf{u}}^{\top}$, and $\mathbf{M}_{\mathbf{u}\mathbf{u}}=\big{[}\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}}\big{]}$. The corresponding optimal ELBO then becomes
^$\hat{\mathcal{L}}_{f}=\log\mathcal{N}(\mathbf{y}|\bm{\mu}_{\mathbf{f}},\mathbf{K}_{\mathbf{f}\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n})-\frac{1}{2}\sigma_{\epsilon}^{-2}\text{tr}(\mathbf{K}_{\mathbf{f}\mathbf{f}}-\mathbf{K}_{\mathbf{f}\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}}).$ (11)
Therefore, eq. 10 can lead to the predictive distribution $\hat{q}(f)=\mathcal{N}(\hat{\mu}_{f}^{\mathbf{u}},\hat{k}_{f}^{\mathbf{u}})$ with
$\hat{\mu}_{f}^{\mathbf{u}}(\cdot)=\sigma_{\epsilon}^{-2}\mathbf{K}_{(\cdot)\mathbf{u}}\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}}\mathbf{y},$ (12a)
$\hat{k}_{f}^{\mathbf{u}}(\cdot,\cdot^{\prime})=k(\cdot,\cdot^{\prime})-\mathbf{K}_{(\cdot)\mathbf{u}}(\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}-\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1})\mathbf{K}_{\mathbf{u}(\cdot^{\prime})}.$ (12b)
2.2 Distributed GP
2.2.1 Distributed GP training
Distibuted GPs (Ng and Deisenroth,, 2014; Deisenroth and Ng,, 2015; Liu et al.,, 2018; Yin et al.,, 2020) enable the parallel training of GPs across different experts to harness greater computing power and accelerate the training process for large-scale datasets. Assuming the same model as in Section 2.1.1, where $f(\cdot)\sim\mathcal{GP}(0,k(\cdot,\cdot^{\prime}))$ and $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ with the i.i.d. noise $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$. We first partition the global training dataset $\mathcal{D}:=\{\mathbf{X},\mathbf{y}\}$ of size $n$ into $M$ local datasets $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$, each of size $n_{i}$ for $i=1,\ldots,M$. Each local dataset $\mathcal{D}_{i}$ corresponds to an expert $\mathcal{M}_{i}$. Clearly, $n=\sum_{i=1}^{M}n_{i}$. We suppose that all local datasets share the same hyperparameters $\bm{\theta}$, as defined in Section 2.1.2. The global objective for distribued GP training is to maximize the global marginal log-likelihood $\log p(\mathbf{y}|\mathbf{f},\bm{\theta})$, where $\mathbf{f}=f(\mathbf{X})$. Assuming independence among all the experts $\{\mathcal{M}_{i}\}_{i=1}^{M}$, the global marginal log-likelihood is approximated as follows
$\log p(\mathbf{y}|\mathbf{f};\bm{\theta})$ $\approx\sum_{i=1}^{M}\log p_{i}(\mathbf{y}_{i}|\mathbf{f}_{i};\bm{\theta})=-\frac{1}{2}\sum_{i=1}^{M}\Big{(}\mathbf{y}_{i}^{\top}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{y}_{i}+\log\det(\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}})+n\log(2\pi)\Big{)}$ (13)
where $p_{i}(\mathbf{y}_{i}|\mathbf{f}_{i},\bm{\theta})\sim\mathcal{N}(\mathbf{0},\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}})$ is the local marginal likelihood of the $i$-th model $\mathcal{M}_{i}$ with $\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}=\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{i}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n_{i}}$ and $\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{i}}=k(\mathbf{X}_{i},\mathbf{X}_{i})\in\mathbb{R}^{n_{i}\times n_{i}}$, $\mathbf{f}_{i}=f(\mathbf{X}_{i})\in\mathbb{R}^{n_{i}}$.
The factorized approximation of the log-likilihood in eq. 13 degenerates the full covariance matrix $\mathbf{K}_{\mathbf{f}\mathbf{f}}=k(\mathbf{X},\mathbf{X})$ into a diagonal block matrix $\text{diag}[\mathbf{K}_{\mathbf{f}_{1},\mathbf{f}_{1}},\ldots,\mathbf{K}_{\mathbf{f}_{M},\mathbf{f}_{M}}]$, reducing the time complexity to $\mathcal{O}(n_{i}^{3})$. This factorized training method (Ng and Deisenroth,, 2014; Deisenroth and Ng,, 2015), termed as FACT, is one common approach to distributed GP training. In addition to FACT, various optimization methods have been proposed. Federated Averaging (FedAvg) (McMahan et al.,, 2017) is a popular aggregation strategy that trains a global model by averaging updates from multiple local datasets, FedProx (Li et al.,, 2020) generalizes FedAvg by adding a proximal term to address the heterogeneity in federated networks. The Alternating Direction of Multipliers Method (ADMM) (Boyd et al.,, 2011) reformulates the optimization problem as a nonconvex consensus problem with newly introduced local hyperparameters and the global hyperparameter. Proximal ADMM (pxADMM) (Hong et al.,, 2016) aims to reduce communication overhead and computational time, and Xie et al., (2019) derive a closed-form solution for distributed GP training based on pxADMM.
2.2.2 Aggregated prediction
Common aggregation methods for GP experts include product-of-experts (PoE) (Hinton,, 2002), generalised product-of-experts (gPoE) (Cao and Fleet,, 2014), Bayesian committee machine (BCM) (Tresp,, 2000), robust Bayesian committee machine (rBCM) (Deisenroth and Ng,, 2015), generalized robust Bayesian committee machine (grBCM) (Liu et al.,, 2018), and nested pointwise aggregation of experts (NPAE) (Rullière et al.,, 2018). The joint mean $\mu_{\mathcal{A}}$ and joint precision $\sigma_{\mathcal{A}}^{-2}$ for PoE, gPoE, BCM, rBCM are given by
$\mu_{\mathcal{A}}(\cdot)=\sigma_{\mathcal{A}}^{2}(\cdot)\sum_{i=1}^{M}\beta_{i}\sigma_{i}^{-2}(\cdot)\mu_{i}(\cdot),$ (14a)
$\sigma_{\mathcal{A}}^{-2}(\cdot)=\sum_{i=1}^{M}\beta_{i}\sigma_{i}^{-2}(\cdot)+(1-\sum_{i=1}^{M}\beta_{i})\sigma_{**}^{-2}(\cdot),$ (14b)
where $\mu_{i}(\cdot)$ and $\sigma_{i}^{-2}(\cdot)$ represent the local mean and local precision of expert $\mathcal{M}_{i}$, while $\sigma_{**}^{-2}(\cdot)=k(\cdot,\cdot)+\sigma_{\epsilon}^{2}$ is the prior variance, serving as a correction term for the BCM family. The weights, $\beta_{i}$, are defined as follows: $\beta_{i}=1$ for PoE and BCM, $\beta_{i}=1/M$ for gPoE, and $\beta_{i}=0.5(\log\sigma_{**}^{2}(\cdot)-\log\sigma_{i}^{2}(\cdot))$ for rBCM.
For grBCM, the $M$ experts are divided into two groups: the global expert $\mathcal{M}_{c}$, trained on dataset $\mathcal{D}_{c}=\mathcal{D}_{1}$ of size $n_{c}$, and the local experts $\{\mathcal{M}_{i}\}_{i=2}^{M}$, trained on datasets $\{\mathcal{D}_{i}\}_{i=2}^{M}$. The augmented dataset $\mathcal{D}_{+i}=\{\mathcal{D}_{c},\mathcal{D}_{i}\}$ leads to a new expert $\mathcal{M}_{+i}$ for $i=2,\ldots,M$. The joint mean and joint precision are given by:
$\mu_{\mathcal{A}}(\cdot)=\sigma_{\mathcal{A}}^{2}(\cdot)\left(\sum_{i=2}^{M}\beta_{i}\sigma_{+i}^{-2}(\cdot)\mu_{+i}(\cdot)-\Big{(}\sum_{i=2}^{M}\beta_{i}-1\Big{)}\sigma_{c}^{-2}(\cdot)\mu_{c}(\cdot)\right),$ (15a)
$\sigma_{\mathcal{A}}^{-2}(\cdot)=\sum_{i=2}^{M}\beta_{i}\sigma_{+i}^{-2}(\cdot)-\Big{(}\sum_{i=2}^{M}\beta_{i}-1\Big{)}\sigma_{c}^{-2}(\cdot),$ (15b)
where $\mu_{c}$ and $\sigma_{c}^{-2}$ are the mean and precision of expert $\mathcal{M}_{c}$, and $\mu_{+i}$ and $\sigma_{+i}^{-2}$ are the mean and precision of expert $\mathcal{M}_{+i}$. The weights are defined as $\beta_{2}=1$, and $\beta_{i}=0.5(\log\sigma_{c}^{2}(\cdot)-\log\sigma_{+i}^{2}(\cdot))$ for $i=3,\ldots,M$.
NPAE leverages the covariances between experts to ensure consistent predictions. Consequently, the joint mean and joint variance are computed as follows:
$\mu_{\mathcal{A}}(\cdot)=\mathbf{K}_{(\cdot)\mathcal{A}}\mathbf{K}_{\mathcal{A}\mathcal{A}}^{-1}\bm{\mu}_{\mathcal{A}},$ (16a)
$\sigma_{\mathcal{A}}^{2}(\cdot)=k(\cdot,\cdot)-\mathbf{K}_{(\cdot)\mathcal{A}}\mathbf{K}_{\mathcal{A}\mathcal{A}}^{-1}\mathbf{K}_{\mathcal{A}(\cdot)}+\sigma_{\epsilon}^{2},$ (16b)
with
$\mathrm{cov}[\mu_{i}(\cdot),y_{*}(\cdot)]=\mathbf{K}_{(\cdot),i}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}(\cdot)},$ (17a)
$\mathrm{cov}[\mu_{i}(\cdot),\mu_{j}(\cdot)]=\begin{cases}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{j}}\widetilde{\mathbf{K}}_{\mathbf{f}_{j}\mathbf{f}_{j}}^{-1}\mathbf{K}_{\mathbf{f}_{j}(\cdot)},&i\neq j,\\ \mathbf{K}_{(\cdot),i}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}(\cdot)},&i=j,\end{cases}$ (17b)
where $\mu_{i}(\cdot)$ represents the local mean of expert $\mathcal{M}_{i}$, $\mathbf{K}_{(\cdot)\mathbf{f}_{i}}=k(\cdot,\mathbf{X}_{i})=\mathbf{K}_{\mathbf{f}_{i}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{j}}=k(\mathbf{X}_{i},\mathbf{X}_{j})$, $\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}=\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{i}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n_{i}}$. $\mathbf{K}_{\mathcal{A}(\cdot)}=\big{[}\mathrm{cov}[\mu_{i}(\cdot),y_{*}(\cdot)]\big{]}_{i=1}^{M}=\mathbf{K}_{(\cdot)\mathcal{A}}^{\top}\in\mathbb{R}^{M}$, $\mathbf{K}_{\mathcal{A}\mathcal{A}}=\big{[}\mathrm{cov}[\mu_{i}(\cdot),\mu_{j}(\cdot)]\big{]}_{i,j=1}^{M}\in\mathbb{R}^{M\times M}$, $\bm{\mu}_{\mathcal{A}}=\big{[}\mu_{1}(\cdot),\ldots,\mu_{M}(\cdot)\big{]}^{\top}\in\mathbb{R}^{M}$.
3 GP regression with OptiCom
In this section, we detail the algorithms for GP regression using the optimized combination technique. First, we introduce the combination technique and the optimized combination technique in Section 3.1 and Section 3.2, respectively. Then, we present the algorithm for applying OptiCom to GP in Section 3.3.
3.1 Combination technique (CT)
Combination technique (CT) (Griebel et al.,, 1990; Hegland et al.,, 2007), first introduced in (Smolyak,, 1963), is an efficient tool for approximating the sparse grid spaces. When partial projection operators commute, as in interpolation with tensor product spaces, the combination technique provides the exact sparse grid solution (Griebel et al.,, 1990). Sparse grids of level $\eta$ and dimension $d$ are defined as (Garcke,, 2013)
$\Omega_{\eta}^{d}:=\bigcup_{|\underline{l}|_{1}=\eta+d-1}\Omega_{\underline{l}}=\bigcup_{|\underline{l}|_{1}=\eta+d-1}\Omega_{l_{1}}^{1}\times\cdots\times\Omega_{l_{d}}^{1}.$ (18)
Here, $\underline{l}=(l_{1},\ldots,l_{d})$, $|\underline{l}|_{1}=\sum_{j=1}^{d}l_{j}$, where $l_{j}\in\mathbb{N}^{+}$ for all $j=1,\ldots d$. $\Omega_{l_{1}}^{1}\times\cdots\times\Omega_{l_{d}}^{1}=\{(\omega_{l_{1}},\ldots,\omega_{l_{d}})|\omega_{l_{j}}\in\Omega_{l_{j}}^{1}\;\text{for}\;j\in\{1,\ldots,d\}\}$ denotes the n-ary Cartesian product over $d$ one-dimensional set $\Omega_{l_{j}}^{1}\subset\mathbb{R}$, $j=1,\ldots,d$. We suppose $\Omega_{l_{j}}^{1}$ consists of $2^{l_{j}-1}$ uniformly distributed points over a fixed interval, i.e., $\Omega_{l_{j}}^{1}$ over the interval $[0,1]$ is given by
$\Omega_{l_{j}}^{1}=\left\{\frac{i}{2^{l_{j}}}:i=1,\ldots,2^{l_{j}-1}\right\}.$ (19)
Each set of grids $\Omega_{\underline{l}}\subset\mathbb{R}^{d}$ is associated with a piecewise $d$-linear basis function $\bm{\phi}_{\underline{l},\underline{h}}(\cdot)$, which is defined as the product of the one-dimensional basis functions:
$\bm{\phi}_{\underline{l},\underline{h}}(\mathbf{x}):=\prod_{j=1}^{d}\phi_{l_{j},h_{j}}(x_{j}),\quad h_{j}=1,\ldots,2^{l_{j}-1}$ (20)
where $\mathbf{x}=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$ is a $d$-dimensional point, $\underline{h}=(h_{1},\ldots,h_{d})$, $\phi_{l_{j},h_{j}}(\cdot)$ is a piecewise linear hierarchical basis shown in Figure 7(a) in Appendix and defined as follows:
$\phi_{l_{j},h_{j}}(x)=\begin{cases}1-\lvert 2^{l_{j}}x-h_{j}\rvert,&x\in\big{[}\frac{h_{j}-1}{2^{l_{j}}},\frac{h_{j}+1}{2^{l_{j}}}\big{]},\\ 0,&\text{otherwise}.\end{cases}$ (21)
These basis functions can define function spaces $V_{\underline{l}}$ on the grids $\Omega_{\underline{l}}$,
$V_{\underline{l}}:=\text{span}\{\bm{\phi}_{\underline{l},\underline{h}}:h_{j}=1,\ldots,2^{l_{j}}-1,j=1,\ldots,d\}.$ (22)
Then the hierarchical increment spaces $W_{\underline{l}}$ can be defined by
$W_{\underline{l}}:=\text{span}\{\bm{\phi}_{\underline{l},\underline{h}}:\underline{h}\in\mathrm{B}_{\underline{l}}\},$ (23)
$\mathrm{B}_{\underline{l}}:=\{\underline{h}\in\mathbb{N}^{d}:h_{j}=1,\ldots,2^{l_{j}}-1,h_{j}\text{ is odd for all }j=1,\ldots,d\}.$ (24)
Therefore, the function spaces can be represented by the hierarchical increment spaces
$V_{\underline{l}}=\bigoplus_{\underline{k}\;\leq\;\underline{l}}W_{\underline{k}},$ (25)
and each function $f_{\underline{l}}\in V_{\underline{l}}$ can be uniquely represented by
$f_{\underline{l}}(\mathbf{x})=\sum_{|\underline{l}|_{1}\leq\eta+d-1}\sum_{\underline{h}\in\mathrm{B}_{\underline{l}}}\alpha_{\underline{k},\underline{h}}\bm{\phi}_{\underline{k},\underline{h}}(\mathbf{x})$ (26)
with hierarchical coefficients $\alpha_{\underline{l},\underline{h}}\in\mathbb{R}$. The combination technique linearly combines the discrete solutions $f_{\underline{l}}(\mathbf{x})$ from the partial grids $\Omega_{\underline{l}}$ according to the combination formula (see Figure 7(b) in Appendix for an example):
${\color[rgb]{0,0,1}{f_{\eta}^{c}(\mathbf{x})=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}(-1)^{\eta+d-1-|\underline{l}|_{1}}\binom{d-1}{|\underline{l}|_{1}-\eta}f_{\underline{l}}(\mathbf{x}),}}$ (27)
where the function $f_{\eta}^{c}$ is in the sparse gris space $V_{\eta}^{s}$ of level $\eta$ defined by
$V_{\eta}^{s}:=\bigoplus_{|\underline{l}|_{1}\leq\eta+d-1}W_{\underline{l}}.$ (28)
3.2 Optimized combination technique (OptiCom)
Optimized Combination Technique (OptiCom) (Hegland et al.,, 2002; Garcke,, 2006; Griebel et al.,, 2015) selects the “best possible” combination coefficients adaptively so that an optimal combination approximation is obtained. More specifically, we aim to minimize the functional
$J(c_{1},\ldots,c_{b})=\|\mathcal{P}_{\eta}^{s}f-\sum_{i=1}^{b}c_{i}\mathcal{P}_{i}f\|^{2},$ (29)
where $\mathcal{P}_{\eta}^{s}f$ denotes the projection into the sparse grid space $V_{\eta}^{s}$, $\mathcal{P}_{i}f$ denotes the projection into one of the spaces $V_{\underline{l}}$ in eq. 27, $b$ is the number of summed terms in combination technique formula eq. 27. Using simple expansions and formula $\langle\mathcal{P}_{\eta}^{s}f,\mathcal{P}_{i}f\rangle=\langle\mathcal{P}_{i}f,\mathcal{P}_{i}f\rangle$, we can obtain
$J(c_{1},\ldots,c_{b})=\sum_{i,j=1}^{b}c_{i}c_{j}\langle\mathcal{P}_{i}f,\mathcal{P}_{j}f\rangle-2\sum_{i=1}^{b}c_{i}\|\mathcal{P}_{i}f\|^{2}+\|\mathcal{P}_{\eta}^{s}f\|^{2}.$ (30)
By differentiating with respect to the combination coefficients $c_{i}$ and setting each of these derivatives to zero, we can derive the following linear system
$\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}&\cdots&\langle\mathcal{P}_{1}f,\mathcal{P}_{b}f\rangle\\ \langle\mathcal{P}_{2}f,\mathcal{P}_{1}f\rangle&\cdots&\langle\mathcal{P}_{2}f,\mathcal{P}_{b}f\rangle\\ \vdots&\ddots&\vdots\\ \langle\mathcal{P}_{b}f,\mathcal{P}_{1}f\rangle&\cdots&\|\mathcal{P}_{b}f\|^{2}\end{bmatrix}\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{b}\end{bmatrix}=\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}\\ \|\mathcal{P}_{2}f\|^{2}\\ \vdots\\ \|\mathcal{P}_{b}f\|^{2}\end{bmatrix}.$ (31)
Using the optimal coefficients $c_{i}$, the OptiCom formula for the sparse grid of level $\eta$ is then given by
${\color[rgb]{0,0,1}{f_{\eta}^{c}(\mathbf{x})=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}c_{\underline{l}}f_{\underline{l}}(\mathbf{x}).}}$ (32)
For the regression problem (Garcke,, 2006), we are seeking the solution of the optimization problem $f^{*}=\arg\min_{f\in V}R(f)$ with
$R(f):=\sum_{i=1}^{n}(f(\mathbf{x}_{i})-y_{i})^{2}+\lambda\|\mathcal{S}f\|_{2}^{2},$ (33)
where $\mathcal{S}$ is a linear operator, $\lambda$ is the tuning parameter that penalizes the flexibility of the model, $f^{*}\in V$ is the unknown function we aim to recover from the given dataset $\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, $\mathbf{x}_{i}\in\mathbb{R}^{d}$, $y_{i}\in\mathbb{R}$. The scalar product in the setting is then defined as
$\langle\mathcal{P}_{\underline{l}}f,\mathcal{P}_{\underline{k}}f\rangle_{\text{RLS}}=\sum_{i=1}^{n}f_{\underline{l}}(\mathbf{x}_{i})f_{\underline{k}}(\mathbf{x}_{i})+\lambda\langle\mathcal{S}f_{\underline{l}},\mathcal{S}f_{\underline{k}}\rangle_{2},$ (34)
so that the minimization eq. 33 is an orthogonal projection of $f^{*}$ into the space $V$, i.e. if $\|f-f^{*}\|_{\text{RLS}}^{2}\leq\|g-f^{*}\|_{\text{RLS}}^{2}$ then $R(f)\leq R(g)$.
3.3 GP with OptiCom
3.3.1 Posterior distribution
It’s easy to extend OptiCom to GP regression. Suppose the inducing inputs $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$ ($\mathbf{z}_{i}\in\mathbb{R}^{d}$) are sparse grids $\Omega_{\eta}^{d}$ of level $\eta$ and dimension $d$ defined in eq. 18, then the predictive mean and covariance in eq. 12 can be approximated in the form
$\hat{\mu}_{\eta}^{c}(\cdot)=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}c_{\underline{l}}\,\cdot\,\sigma_{\epsilon}^{-2}\mathbf{K}_{(\cdot)\mathbf{u}_{\underline{l}}}\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}^{-1}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{y},$ (35a)
$\hat{k}_{\eta}^{c}(\cdot,\cdot^{\prime})=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}(-1)^{\eta+d-1-|\underline{l}|_{1}}\binom{d-1}{|\underline{l}|_{1}-\eta}\Big{[}k(\cdot,\cdot^{\prime})-\mathbf{K}_{(\cdot)\mathbf{u}_{\underline{l}}}(\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}^{-1}-\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}^{-1})\mathbf{K}_{\mathbf{u}_{\underline{l}}(\cdot^{\prime})}\Big{]},$ (35b)
where $\mathbf{u}_{\underline{l}}$ is the vector of inducing variables corresponding to the inducing inputs $\Omega_{\underline{l}}$ defined in eq. 18, $\mathbf{K}_{(\cdot)\mathbf{u}_{\underline{l}}}=k(\cdot,\Omega_{\underline{l}})=\mathbf{K}_{\mathbf{u}_{\underline{l}}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}=k(\Omega_{\underline{l}},\Omega_{\underline{l}})$, and $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}=\big{[}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}_{\underline{l}}}\big{]}$.
3.3.2 Optimal coefficients
To compute the optimal coefficients $c_{\underline{l}}$ in GP regression, we need to compute the inner product with respect to the reproducing kernel Hilbert space (RKHS) (Aronszajn,, 1950). Let $\mathcal{H}$ be a RKHS endowed with an inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$, then the inner product between functions $f(\cdot)=\sum_{i=1}^{n}\alpha_{i}k(\cdot,\mathbf{x}_{i})$ and $g(\cdot)=\sum_{j=1}^{n^{\prime}}\alpha^{\prime}_{j}k(\cdot,\mathbf{x}^{\prime}_{j})$ is given by
$\langle f,g\rangle_{\mathcal{H}}=\sum_{i=1}^{n}\sum_{j=1}^{n^{\prime}}\alpha_{i}\alpha^{\prime}_{j}k(\mathbf{x}_{i},\mathbf{x}_{j}^{\prime}).$ (36)
In GP regression, the optimization problem in eq. 33 becomes
$R(f)=\sum_{i=1}^{n}(f(\mathbf{x}_{i})-y_{i})^{2}+\lambda\|f\|_{\mathcal{H}}^{2},$ (37)
where $\lambda=\sigma_{\epsilon}^{2}$ is the noise varaince. Here the projection into one of the spaces $V_{\underline{l}}$ has the form
$\mathcal{P}_{\underline{l}}f=\sigma_{\epsilon}^{-2}\mathbf{K}_{(\cdot)\mathbf{u}_{\underline{l}}}\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}^{-1}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{y}=\sigma_{\epsilon}^{-2}k(\cdot,\Omega_{\underline{l}})\bm{\alpha}_{\underline{l}}=\sigma_{\epsilon}^{-2}\sum_{i\in\mathbf{I}_{\underline{l}}}k(\cdot,\mathbf{z}_{i})\alpha_{i}:=f_{\underline{l}}(\cdot),$ (38)
where $\mathbf{I}_{\underline{l}}\subset\{1,\ldots,m\}$ is the set of indices of the inducing inputs $\{\mathbf{z}_{i}\}_{i=1}^{m}$ such that $\Omega_{\underline{l}}=\{\mathbf{z}_{i}:i\in\mathbf{I}_{\underline{l}}\}$, $\alpha_{i}$ is the $i$-th component of the vector $\bm{\alpha}_{\underline{l}}=\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}^{-1}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{y}$. The scalar product defined in eq. 34 becomes
$\langle\mathcal{P}_{\underline{l}}f,\mathcal{P}_{\underline{k}}f\rangle_{\text{RLS}}$ $=\sum_{i=1}^{n}f_{\underline{l}}(\mathbf{x}_{i})f_{\underline{k}}(\mathbf{x}_{i})+\lambda\langle f_{\underline{l}},f_{\underline{k}}\rangle_{\mathcal{H}}$
$=\sigma_{\epsilon}^{-4}\sum_{i=1}^{n}k(\mathbf{x}_{i},\Omega_{\underline{l}})\bm{\alpha}_{\underline{l}}k(\mathbf{x}_{i},\Omega_{\underline{k}})\bm{\alpha}_{\underline{k}}+\lambda\sigma_{\epsilon}^{-4}\;\bm{\alpha}_{\underline{l}}^{\top}k(\Omega_{\underline{l}},\Omega_{\underline{k}})\bm{\alpha}_{\underline{k}}$
$=\sigma_{\epsilon}^{-2}\bm{\alpha}_{\underline{l}}^{\top}\big{[}\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}_{\underline{k}}}+\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}\big{]}\bm{\alpha}_{\underline{k}}:=\sigma_{\epsilon}^{-2}\bm{\alpha}_{\underline{l}}^{\top}\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}\bm{\alpha}_{\underline{k}},$ (39)
where $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}=\big{[}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}_{\underline{k}}}\big{]}$. The details are outlined in Algorithm 1 and Algorithm 2 in Appendix.
Figure 1: Errors and time for posterior mean of zero mean and Matérn 3/2 GP with OptiCom and CT of dimension $d=2$ and level $\eta=1,2,\ldots,7$ tested on the Griewank function (Griewank,, 1981) over $n=2000$ training points and $n_{t}=1000$ test points. We denote GP with OptiCom by red dots, GP with CT by blue diamonds. Left: Logarithm of $l_{2}$ errors between GP posteriors and ground truth values averaged over 10 different lengthscales $[0.2,0.4,0.6,0.8,1.0,1.2,1.4,1.6,1.8,2.0]$. Right: Logarithm of time taken to compute the posterior mean averaged over $1000$ replications.
3.3.3 Complexity
From Lemma 6 in (Garcke,, 2013), we know that the number of summed terms in combination technique formula eq. 27 is $b=\mathcal{O}(\eta^{d-1})$, the size of each summed term is of order $m_{b}=\mathcal{O}(2^{(\eta+d-1)})$, and dimension of the sparse grid space $V_{\eta}^{s}$ defined in eq. 28, i.e., the number of grid points, is $|V_{\eta}^{s}|=\mathcal{O}(2^{\eta}\cdot\eta^{d-1})=\mathcal{O}\big{(}h_{\eta}^{-1}\cdot\log(h_{\eta}^{-1})^{d-1}\big{)}$, where $h_{\eta}=2^{-\eta}$ is the mesh size of the sparse grids $\Omega_{\eta}^{d}$. The additional operations compared to GP with CT include computing $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}$ when $\underline{l}\neq\underline{k}$ and solving a $b\times b$ linear system to obtain the optimal coefficients, which costs $\mathcal{O}(b^{2}m_{b}^{2}n)$ and $\mathcal{O}(b^{3})$ time respectively. Note that computation of $\{\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}\}_{\underline{l},\underline{k}\in\mathbb{I}(\eta,d)}$ and $\{\text{chol}(\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}})\}_{\underline{l}\in\mathbb{I}(\eta,d)}$ can be parallelized and reduced to $\mathcal{O}(m_{b}^{2}n)$ time and $\mathcal{O}(m_{b}^{3})$ time respectively, where $\mathbb{I}(\eta,d)=\{\underline{l}:\eta\leq|\underline{l}|_{1}\leq\eta+d-1\}$. Thus the computational complexity of the parallelized Algorithm 2 is $\mathcal{O}(m_{b}^{3}+b^{3}+m_{b}^{2}n)=\mathcal{O}(b^{3}+m_{b}^{2}n)$. Although the computational complexity of the parallelized GP with OptiCom is larger than that of parallelized GP with CT theoretically, which is $\mathcal{O}(m_{b}^{2}n)$, $b$ is very small in most cases, hence $\mathcal{O}(b^{3}+m_{b}^{2}n)\approx\mathcal{O}(m_{b}^{2}n)$. Figure 1 also demonstrates that GP with OptiCom achieves lower errors and smaller variance than GP with CT at nearly the same time cost.
4 Distributed GP with optimal weights
In this section, we extend the OptiCom to distributed GPs. Following the assumptions Section 2.2.1, we consider the global training dataset $\mathcal{D}=\{\mathbf{X},\mathbf{y}\}$ of size $n$, partitioned into $M$ local datasets $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$ of size $n_{i}$ for $i=1,\ldots,M$. All the local datasets are assumed to share the same hyperparameters $\bm{\theta}$.
4.1 Distributed SVGP
4.1.1 Optimal weights
To obtain the optimal weights, as shown the eq. 29 and eq. 30 in Section 3.2, the objective is to minimize
$J(\beta_{1},\ldots,\beta_{M})=\|f-\sum_{i=1}^{M}\beta_{i}\mathcal{P}_{i}f\|,$ (40)
where $f(\cdot)\sim\mathcal{GP}(0,k(\cdot,\cdot^{\prime}))$ as defined in Section 2.2.1, $\mathcal{P}_{i}f=f|\mathbf{y}_{i}:=f_{i}$ is the posterior conditioned on the local dataset $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$, and its distribution takes the form of Equation 12. Similar to Section 3.3, the optimal weights $\{\beta_{i}\}_{i=1}^{M}$ can be obtained by solving the following linear system:
$\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}&\cdots&\langle\mathcal{P}_{1}f,\mathcal{P}_{M}f\rangle\\ \langle\mathcal{P}_{2}f,\mathcal{P}_{1}f\rangle&\cdots&\langle\mathcal{P}_{2}f,\mathcal{P}_{M}f\rangle\\ \vdots&\ddots&\vdots\\ \langle\mathcal{P}_{M}f,\mathcal{P}_{1}f\rangle&\cdots&\|\mathcal{P}_{M}f\|^{2}\end{bmatrix}\begin{bmatrix}\beta_{1}\\ \beta_{2}\\ \vdots\\ \beta_{M}\end{bmatrix}=\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}\\ \|\mathcal{P}_{2}f\|^{2}\\ \vdots\\ \|\mathcal{P}_{M}f\|^{2}\end{bmatrix}.$ (41)
To reduce computational complexity, we randomly select a central dataset $\mathcal{D}_{c}:=\{\mathbf{X}_{c},\mathbf{y}_{c}\}\subset\mathcal{D}=\{\mathbf{X},\mathbf{y}\}$ of size $n_{c}\ll n$. In practice, we select one point from each local dataset $\mathcal{D}_{i}$ and combine them to form the central dataset, resulting in $n_{c}=M$. Consequently, the scalar product is approximated as follows:
$\langle\mathcal{P}_{l}f,\mathcal{P}_{k}f\rangle_{\text{RLS}}$ $=\sum_{\mathbf{x}_{i}\in\mathbf{X}}f_{l}(\mathbf{x}_{i})f_{k}(\mathbf{x}_{i})+\lambda\langle f_{l},f_{k}\rangle_{\mathcal{H}}$
$\approx\sum_{\mathbf{x}_{i}\in\mathbf{X}_{c}}f_{l}(\mathbf{x}_{i})f_{k}(\mathbf{x}_{i})+\lambda\langle f_{l},f_{k}\rangle_{\mathcal{H}}$
$=\sigma_{\epsilon}^{-2}\bm{\alpha}_{l}^{\top}\big{[}\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}_{c}}\mathbf{K}_{\mathbf{f}_{c}\mathbf{u}}+\mathbf{K}_{\mathbf{u}\mathbf{u}}\big{]}\bm{\alpha}_{k}:=\sigma_{\epsilon}^{-2}\bm{\alpha}_{l}^{\top}\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(c)}\bm{\alpha}_{k},$ (42)
where $\bm{\alpha}_{l}=[\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(l)}]^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}\mathbf{y}_{l}$, $\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(l)}=\big{[}\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}\mathbf{K}_{\mathbf{f}_{l}\mathbf{u}}\big{]}$, $\mathbf{K}_{\mathbf{u}\mathbf{u}}=k(\mathbf{Z},\mathbf{Z})$, $\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}=k(\mathbf{Z},\mathbf{X}_{l})=\mathbf{K}_{\mathbf{f}_{l}\mathbf{u}}^{\top}$, $\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(c)}=\big{[}\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}_{c}}\mathbf{K}_{\mathbf{f}_{c}\mathbf{u}}\big{]}$, $\mathbf{K}_{\mathbf{u}\mathbf{f}_{c}}=k(\mathbf{Z},\mathbf{X}_{c})=\mathbf{K}_{\mathbf{f}_{c}\mathbf{u}}^{\top}$, $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$ are the inducing inputs corresponding to the optimized inducing variables $\mathbf{u}\in\mathbb{R}^{m}$. The details are outlined in Algorithm 3 in Appendix.
4.1.2 Aggregated prediction
Instead of relying on PoE-based models, we directly use the optimal weights $\{\beta_{i}\}_{i=1}^{M}$ to predict the aggregated GP models after distributed training. Note that the correlations between the experts are already incorporated into the optimal weights. Therefore, we use the weighted local variances to approximate the aggregated variance, reducing computational complexity. The joint mean and joint variance are given as follows:
$\mu_{\mathcal{A}}(\cdot)=\sum_{i=1}^{M}\beta_{i}\mathbb{E}[f_{i}(\cdot)]=\sum_{i=1}^{M}\beta_{i}\sigma_{\epsilon}^{-2}\mathbf{K}_{(\cdot)\mathbf{u}}\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(i)-1}\mathbf{K}_{\mathbf{u}\mathbf{f}_{i}}\mathbf{y}_{i},$ (43a)
$\sigma_{\mathcal{A}}^{2}(\cdot)=\sum_{i=1}^{M}\beta_{i}^{2}\mathrm{Var}[f_{i}(\cdot),f_{i}(\cdot)]=\sum_{i=1}^{M}\beta_{i}^{2}\mathbf{K}_{(\cdot)\mathbf{u}}[\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(i)}]^{-1}\mathbf{K}_{\mathbf{u}(\cdot)},$ (43b)
The details are outlined in Algorithm 4 in Appendix.
4.1.3 Complexity
In Algorithm 3, the computation of $\{\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(i)}\}_{i=1}^{M}$ and $\{\text{chol}(\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(i)})\}_{i=1}^{M}$ requires $\mathcal{O}(m^{2}n_{i})$ time and $\mathcal{O}(m^{3})$ time, respectively, after parallelization, and solving the $M\times M$ linear system costs $\mathcal{O}(M^{3})$ time. Therefore, the time complexity of the parallelized Algorithm 3 is $\mathcal{O}(m^{2}n_{i}+m^{3}+M^{3})=\mathcal{O}(m^{2}n_{i}+M^{3})$. For the aggregated prediction, the computational complexity of the joint mean and joint variance is $\mathcal{O}(M)$. Thus, the computational complexity of the parallelized Algorithm 4 is $\mathcal{O}(m^{2}n_{i}+M^{3}+M)=\mathcal{O}(m^{2}n_{i}+M^{3})$, where $m$ is the size of shared inducing variables $\mathbf{u}$ across the experts $\{\mathcal{M}_{i}\}_{i=1}^{M}$, $n_{i}$ is the size of the local dataset $\mathcal{D}_{i}$ for expert $\mathcal{M}_{i}$, ($i=1,\ldots,M$), and $M$ is the number of experts. The comparison of the time complexity of the aggregation models and full SVGP is presented in the first row of Table 1.
Table 1: Complexity of the prediction for aggregation models and full GP. $m$ is the number of inducing points, $n$ is the size of the entire training dataset, $n_{i}$ is the size of the local dataset $\mathcal{D}_{i}$ corresponding to the expert $\mathcal{M}_{i}$, $i=1,\ldots.\,M$, $n_{c}$ is the size of the dataset $\mathcal{D}_{c}$ corresponding to the expert $\mathcal{M}_{c}$ for grBCM. $M$ is the number of experts.
Full PoE gPoE BCM rBCM grBCM NPAE opt (ours)
Time SVGP $\mathcal{O}(m^{2}n)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}(n_{i}+n_{c})+M)$ $\mathcal{O}(m^{2}n_{i}M+M^{3})$ $\mathcal{O}(m^{2}n_{i}+M^{3})$
ExactGP $\mathcal{O}(n^{3})$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}((n_{i}+n_{c})^{3}+M)$ $\mathcal{O}(n_{i}^{3}M+M^{3})$ $\mathcal{O}(n_{i}^{3}+M^{3})$
4.2 Distributed exact GP
The optimal weights method for distributed GPs is also applicable to exact GPs. In this case, the distribution of $\mathcal{P}_{i}f=f|\mathbf{y}_{i}:=f_{i}$ follows the form given in eq. 4. Consequently, the scalar product is expressed as:
$\langle\mathcal{P}_{l}f,\mathcal{P}_{k}f\rangle_{\text{RLS}}$ $\approx\sum_{\mathbf{x}_{i}\in\mathbf{X}_{c}}f_{l}(\mathbf{x}_{i})f_{k}(\mathbf{x}_{i})+\lambda\langle f_{l},f_{k}\rangle_{\mathcal{H}}$
$=\bm{\alpha}_{l}^{\top}\big{[}\mathbf{K}_{\mathbf{f}_{l}\mathbf{f}_{c}}\mathbf{K}_{\mathbf{f}_{c}\mathbf{f}_{k}}+\mathbf{K}_{\mathbf{f}_{l}\mathbf{f}_{k}}\big{]}\bm{\alpha}_{k}:=\bm{\alpha}_{l}^{\top}\mathbf{M}_{\mathbf{f}_{l}\mathbf{f}_{k}}^{(c)}\bm{\alpha}_{k},$ (44)
where $\bm{\alpha}_{l}=\widetilde{\mathbf{K}}_{\mathbf{f}_{l}\mathbf{f}_{l}}^{-1}\mathbf{y}_{l}$, $\widetilde{\mathbf{K}}_{\mathbf{f}_{l}\mathbf{f}_{l}}=\mathbf{K}_{\mathbf{f}_{l}\mathbf{f}_{l}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n_{l}}$, $\mathbf{M}_{\mathbf{f}_{l}\mathbf{f}_{k}}^{(c)}=\big{[}\mathbf{K}_{\mathbf{f}_{l}\mathbf{f}_{k}}+\mathbf{K}_{\mathbf{f}_{l}\mathbf{f}_{c}}\mathbf{K}_{\mathbf{f}_{c}\mathbf{f}_{k}}\big{]}$, $\mathbf{K}_{\mathbf{f}_{l}\mathbf{f}_{k}}=k(\mathbf{X}_{l},\mathbf{X}_{k})$. Therefore, the joint mean and joint variance are given by:
$\mu_{\mathcal{A}}(\cdot)=\sum_{i=1}^{M}\beta_{i}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{y}_{i},$ (45a)
$\sigma_{\mathcal{A}}^{2}(\cdot)=\sum_{i=1}^{M}\beta_{i}^{2}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}(\cdot)},$ (45b)
The detailed procedures are outlined in Algorithm 5 and Algorithm 6 in Appendix. The time complexity of the parallelized Algorithm 6 is $\mathcal{O}(n_{i}^{3}+M^{3})$, where $n_{i}$ is the size of the local dataset $\mathcal{D}_{i}$ for expert $\mathcal{M}_{i}$, and $M$ is the number of experts. A comparison of the time complexity of the aggregation models and the full exact GP is presented in the last row of Table 1.
5 Experiment
In this section, we evaluate the proposed algorithm against several baseline methods: full GP, PoE, gPoE, BCM, rBCM, and grBCM, using both synthetic data (Section 5.1) and UCI datasets (Asuncion et al.,, 2007) (Section 5.2). The size of the inducing variables is set to $m=128$. We use the radial basis function (RBF) kernel with separate lengthscales for each input dimension, defined by the following equation:
$k(\mathbf{x},\mathbf{x}^{\prime})=\sigma_{f}^{2}\exp\left(-\sum_{i=1}^{d}\frac{(\mathbf{x}_{i}-\mathbf{x}^{\prime}_{i})^{2}}{2l_{i}^{2}}\right),$ (46)
where $\mathbf{x}_{i},\mathbf{x}^{\prime}_{i}\in\mathbb{R}$ are the $i$-th components of the inputs $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}$, $l_{i}$ is the lengthscale along the $i$-th dimension, and $\sigma_{f}^{2}>0$ is the signal variance.
5.1 Synthetic data
5.1.1 Training
Figure 2: Hyperparameter estimates versus the training iterations on the $n=10^{4}$ dataset.
In this section, we assess the performance of four distributed training methods for SVGP: FACT, ADMM, FedAvg, and FedProx. We generate a training dataset of size $n=10^{4}$ and dimension $d=2$ using the RBF kernel with $(l_{1},l_{2},\sigma_{\epsilon},\sigma_{f})=(0.3,1.3,0.1,0.2)$. The dataset is distributed among $M=2,4,8,19,20,40,80,100$ experts, respectively. We set the initial values of the hyperparameters $(l_{1}^{(0)},l_{2}^{(0)},\sigma_{\epsilon}^{(0)},\sigma_{f}^{(0)})=(2.0,0.3,1.0,\sqrt{1.5})$, and use Adam optimizer (Kingma and Ba,, 2014) with a learning rate of $0.1$ for all methods. For ADMM, the regularization parameter is set to $\rho=500$. For FedProx, the proximal coefficient is set to $\mu=0.01$, with a sampling rate of 0.5 for the local updates. The hyperparameters of the SVGP $\bm{\theta}=(l_{1},l_{2},\sigma_{\epsilon},\sigma_{f},\mathbf{u})^{\top}$ are trained over 200 iterations, where $\mathbf{u}$ are the inducing variables of size $m=128$. The results are averaged over 10 different seeds for each setting of $M$.
Figure 2 illustrates the change of the parameters with the training iterations for different numbers of experts $M$. We observe that FACT and ADMM perform worse as the number of experts $M$ increases, while FedAvg performs the best overall, converging faster and more stably than FedProx. Figure 8 in Appendix shows boxplots of the hyperparameter estimates after 200 iterations, indicating that FedAvg converges closest to the ground truth with the least variability. Therefore, we use FedAvg for distributed training in the subsequent experiments with the UCI datasets.
5.1.2 Prediction
In this section, we evaluate the performance of aggregation models and full GP models without optimizing any hyperparameters. We use the Ackley function (Ackley,, 2012) defined over the interval $[-1,1]^{d}$, to benchmark the aggregation models:
$f(\mathbf{x})=-a\exp\left(-b\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}\right)-\exp\left(\frac{1}{d}\sum_{i=1}^{d}\cos(cx_{i})\right)+a+\exp(1),\quad\mathbf{x}\in\mathbb{R}^{d},$ (47)
with parameters $a=20$, $b=0.2$, $c=2\pi$. We employ the RBF kernel with fixed hyperparameters $(l_{i},\sigma_{\epsilon}^{2},\sigma_{f}^{2})=(0.3,0.025,1.0)$ for all $i=1,\ldots,d$. For SVGP, we use $m=128$ inducing points, randomly generated with a fixed seed across all aggregation methods.
Figure 3: SVGP posteriors with the different aggregation models on the Ackley function on $n=10^{4}$ training points and $n_{t}=50$ test points with the number of experts $M=4$, dimension $d=1$, and the the number of inducing variables $m=128$. We denote the aggregated predictions (aggreg), by red lines, the full SVGP posteriors (full) by blue lines, and the ground truth values (true) by black dots.
Figure 3 shows the aggregated predictions of SVGP on $n=10^{4}$ training points and $n_{t}=50$ test points, with dimension $d=1$ and $M=4$ experts. The proposed method (opt) and five baselines (PoE, gPoE, BCM, rBCM, grBCM) are compared to the full SVGP. All methods use the same inducing points. It’s observed that PoE, gPoE, grBCM and the proposed method (opt) provide better predictive means than BCM and rBCM. However, the predictive variances of PoE, BCM, and rBCM shrink to zero and is not consistent with the full SVGP. Our method, the distributed SVGP with optimal weights, delivers accurate predictions for both the mean and the variance.
(a) mean
(b) standard deviation
Figure 4: Comparison of the aggregation models for Exact GP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. The number of experts considered are $M=2,4,8,10,20,40,80,100$. The lower $x$-axis represents the size of the local training dataset $n_{i}=n/M$, and the upper $x$-axis represents the number of the experts $M$. Left: Logarithm of time for computing the aggregated predictions. Middle: RMSE between the aggregated predictions and the ground truth. Right: NLPD between the aggregated predictions and the ground truth. Top: Mean of the computational time, RMSE and NLPD. Bottom: Standard deviation of the computational time, RMSE and NLPD.
(a) mean
(b) standard deviation
Figure 5: Comparison of the aggregation models for SVGP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, $m=128$ inducing points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. The number of experts considered are $M=2,4,8,10,20,40,80,100$. The lower $x$-axis represents the size of the local training dataset $n_{i}=n/M$, and the upper $x$-axis represents the number of the experts $M$. Left: Logarithm of time for computing the aggregated predictions. Middle: RMSE between the aggregated predictions and the ground truth. Right: NLPD between the aggregated predictions and the ground truth. Top: Mean of the computational time, RMSE and NLPD. Bottom: Standard deviation of the computational time, RMSE and NLPD.
Next, we evaluate the aggregation models on $n=10^{4}$ training data of dimension $d=2$ and $n_{t}=50\times 50=2500$ test points, evenly spaced over the interval $[-1,1]^{2}$. The training dataset is divided among $M=2,4,8,10,20,40,80,100$ experts, respectively. To eliminate the impact of lengthscales and seeds, we average all the metrics over 10 different lengthscales $[0.2,0.4,0.6,0.8,1.0,1.2,1.4,1.6,1.8,2.0]$ and 10 different seeds $1,2,3,4,5,6,7,8,9,10$ for each setting of $M$.
Figure 4 compares three metrics for exact GPs: computational time, root mean squared error (RMSE), and negative log predictive density (NLPD) of all aggregation models as a function of the number of training points per GP expert. While grBCM preforms the best on the mean of RMSE and NLPD, it requires more computational time due to the larger size of the augmented local training datasets and has larger standard deviations of the metrics. Our method, GP with optimal weights, provides comparable means for these metrics with smaller standard deviations, indicating greater stability and less sensitivity compared to grBCM. The boxplots of RMSE and NLPD for exact GPs are presented in Figure 9 and Figure 10 in Appendix. These boxplots highlight the stability of our method, especially as the number of experts $M$ increases.
Figure 5 compares three metrics as a function of the number of training points per GP expert for SVGP across all aggregation models. For SVGP, PoE, gPoE, grBCM, significantly outperform the full SVGP, BCM, and rBCM. The boxplots of RMSE and NLPD for SVGP are shown in Figure 11 and Figure 12 in Appendix. These boxplots demonstrate that PoE, gPoE, grBCM, and our method all achieve the smallest means and variabilities for these metrics.
5.2 UCI regression
Figure 6: Comparison of the aggregation models for SVGP on the UCI datasets.
In this section, we evaluate the aggregation models on five real datasets from the UCI repository (Asuncion et al.,, 2007): PoleTele, Elevators, Kin40k, KEGG, and Protein. Each dataset is split into training and test sets in an 80:20 ratio. For the smaller datasets (PoleTele and Elevators, each around $15,000$ samples), the training datasets are divided into $M=4,10,20$ experts. For the larger datasets (Kin40k, KEGG, and Protein, each around $45,000$ samples), the training datasets are divided into $M=8,16,64$ experts. We use SVGP with $m=128$ inducing points and RBF kernel defined in eq. 46. All aggregation models utilize FedAvg with the Adam optimizer, using the same initial hyperparameters over 100 iterations for distributed training. Results are averaged over 10 different seeds. Figure 6 presents the performance metrics of the aggregation models on these five datasets. BCM performs the worst, particularly as the number of experts $M$ increases. While grBCM performs well for RMSE on most datasets, it has significantly worse NLPD on PoleTele. Although our method, distributed SVGP with optimal weights, does not always perform the best, it consistently provides stable and comparable predictions in less time than grBCM. Detailed metric values are reported in Table 2 and Table 3 in Appendix.
6 Conclusion
In this work, we introduced an algorithm to enhance GP regressions with inducing points on sparse grids using the optimized combination technique (OptiCom). Building on this foundation, we proposed a novel aggregated prediction algorithm suitable for both distributed exact GP and SVGP. This method leverages predictions from local models and optimally combines them using weights derived from solving a linear system. Additionally, it inherently incorporates correlations between experts through the computation of optimal weights. The proposed methods have a computational complexity requiring $\mathcal{O}(M^{3})$ additional time compared to PoE family and BCM family for $M$ GP experts. However, this overhead is minimal in practice, as the number of experts $M$ is typically small compared to the size of the local datasets. Through extensive experiments and comparisons with state-of-the-art aggregation models, we demonstrated the competitiveness and stability of our proposed method. Our algorithm consistently provides stable and accurate predictions, offering a robust solution for distributed GP regression tasks.
Acknowledgments and Disclosure of Funding
The authors acknowledge the generous support from the NSF grants DMS-2312173 and CNS-2328395. The experiments of this research were conducted on an A100 GPU provided by Texas A&M High Performance Research Computing.
Appendix A GP with OptiCom
A.1 Figures
(a) Black lines: Hierarchical basis of level $\eta=3$ in one dimension.
(b) Combination technique of level $\eta=3$ in $d=2$ dimensions: combine full grids $\Omega_{\underline{l}}$, $|\underline{l}|_{1}\in\{3,4\}$, to get a sparse grid $\Omega_{\eta=3}^{d=2}$ corresponding to $V_{3}^{s}$.
Figure 7: Basis functions and combination technique of level $\eta=3$.
A.2 Algorithms
Algorithm 1 Optimal Coefficients for GP with OptiCom
1:  Input: training inputs $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$, observations $\mathbf{y}=(y_{1},\ldots,y_{n})^{\top}$, noise variance $\sigma_{\epsilon}^{2}$, a GP denoted by $\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime}))$, sparse grids $\Omega_{\eta}^{d}$ of level $\eta$ and dimension $d$
2:  Output: coefficients $\mathbf{c}=\{c_{\underline{l}^{(i)}}\}_{i=1}^{b}$, kernel weights $\{\bm{\alpha}_{\underline{l}^{(i)}}\}_{i=1}^{b}$, Cholesky factors $\{\mathbf{L}_{\underline{l}^{(i)}}\}_{i=1}^{b}$, where $\{\underline{l}^{(i)}\}_{i=1}^{b}=\{\underline{l}:\eta\leq|\underline{l}|_{1}\leq\eta+d-1\}$
3:  $\mathbf{A}:=\text{zeros}(b,b)$
[…]
6:        $j:=i+t$
7:        $\mathbf{M}=\big{[}k(\Omega_{\underline{l}^{(i)}},\Omega_{\underline{l}^{(j)}})+\sigma_{\epsilon}^{-2}k(\Omega_{\underline{l}^{(i)}},\mathbf{X})k(\mathbf{X},\Omega_{\underline{l}^{(j)}})\big{]}$
[…]
9:           $\mathbf{L}_{\underline{l}^{(i)}}:=\text{chol}(\mathbf{M})$ s.t. $\mathbf{L}_{\underline{l}^{(i)}}\mathbf{L}_{\underline{l}^{(i)}}^{\top}=\mathbf{M}$
10:           $\mathbf{R}:=\mathbf{L}_{\underline{l}^{(i)}}\,\backslash\,\big{[}k(\Omega_{\underline{l}},\mathbf{X})\mathbf{y}\big{]}$
11:           $\bm{\alpha}_{\underline{l}^{(i)}}:=\mathbf{L}_{\underline{l}^{(i)}}^{\top}\,\backslash\,\mathbf{R}$
12:        end if
13:        $\mathbf{A}[i,j]=\mathbf{A}[j,i]=\sigma_{\epsilon}^{-2}\bm{\alpha}_{\underline{l}^{(i)}}^{\top}\mathbf{M}\bm{\alpha}_{\underline{l}^{(j)}}$
14:     end for
15:  end for
16:  $\mathbf{c}:=\mathbf{A}\,\backslash\,\text{diag}(\mathbf{A})$
17:  return $\mathbf{c}$, $\{\bm{\alpha}_{\underline{l}^{(i)}}\}_{i=1}^{b}$, $\{\mathbf{L}_{\underline{l}^{(i)}}\}_{i=1}^{b}$
Algorithm 2 GP Posterior with OptiCom
1:  Input: training inputs $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$, observations $\mathbf{y}=(y_{1},\ldots,y_{n})^{\top}$, noise variance $\sigma_{\epsilon}^{2}$, test inputs $\mathbf{X}^{*}=\{\mathbf{x}_{i}^{*}\}_{i=1}^{n_{t}}$, a GP denoted by $\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime}))$, sparse grids $\Omega_{\eta}^{d}$ of level $\eta$ and dimension $d$
2:  Output: posterior mean $\hat{\mu}_{\eta}^{c}(\mathbf{X}^{*})$ and posterior covariance $\hat{k}_{\eta}^{c}(\mathbf{X}^{*},\mathbf{X}^{*})$ in eq. 35
3:  $\mathbf{m}:=\text{zeros}(n_{t})$; $\mathbf{S}:=\text{zeros}(n_{t},n_{t})$
4:  get $\mathbf{c}=\{c_{\underline{l}^{(i)}}\}_{i=1}^{b}$, $\{\bm{\alpha}_{\underline{l}^{(i)}}\}_{i=1}^{b}$, $\{\mathbf{L}_{\underline{l}^{(i)}}\}_{i=1}^{b}$ from Algorithm 1
[…]
6:     $\mathbf{m}\leftarrow\mathbf{m}+c_{\underline{l}^{(i)}}k(\mathbf{X}^{*},\Omega_{\underline{l}^{(i)}})\bm{\alpha}_{\underline{l}^{(i)}}$
7:     $c_{\text{ct}}:=(-1)^{\eta+d-1-|\underline{l}^{(i)}|_{1}}\binom{d-1}{|\underline{l}^{(i)}|_{1}-\eta}$
8:     $\mathbf{P}:=k(\mathbf{X}^{*},\Omega_{\underline{l}^{(i)}})\big{[}k(\Omega_{\underline{l}^{(i)}},\Omega_{\underline{l}^{(i)}})\big{]}^{-1}k(\Omega_{\underline{l}^{(i)}},\mathbf{X}^{*})$
9:     $\mathbf{Q}:=\big{[}k(\mathbf{X}^{*},\Omega_{\underline{l}^{(i)}})\mathbf{L}_{\underline{l}^{(i)}}^{-\top}\big{]}\big{[}\mathbf{L}_{\underline{l}^{(i)}}^{-1}k(\Omega_{\underline{l}^{(i)}},\mathbf{X}^{*})\big{]}$
10:     $\mathbf{S}\leftarrow\mathbf{S}+c_{\text{ct}}\big{[}k(\mathbf{X}^{*},\mathbf{X}^{*})-\mathbf{P}+\mathbf{Q}\big{]}$
11:  end for
12:  $\hat{\mu}_{\eta}^{c}(\mathbf{X}^{*}):=\sigma_{\epsilon}^{-2}\mathbf{m}$
13:  $\hat{k}_{\eta}^{c}(\mathbf{X}^{*},\mathbf{X}^{*}):=\mathbf{S}$
14:  return $\hat{\mu}_{\eta}^{c}(\mathbf{X}^{*})$, $\hat{k}_{\eta}^{c}(\mathbf{X}^{*},\mathbf{X}^{*})$
Appendix B Distributed GP with optimal weights
B.1 Algorithms
Algorithm 3 Optimal Weights for Distributed SVGP
1:  Input: local datasets $\{\mathbf{X}_{i},\mathbf{y}_{i}\}_{i=1}^{M}$, noise variance $\sigma_{\epsilon}^{2}$, a GP denoted by $\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime}))$, inducing inputs $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$
2:  Output: optimal weights $\bm{\beta}=\{\beta_{i}\}_{i=1}^{M}$, kernel weights $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, Cholesky factors $\{\mathbf{L}_{i}\}_{i=1}^{M}$
3:  $\mathbf{X}_{c}:=\{\mathbf{x}_{i}^{(c)}\}_{i=1}^{M}$, $\mathbf{x}_{i}^{(c)}\in\mathbf{X}_{i}$ is randomly selected
4:  $\mathbf{M}^{(c)}=\big{[}k(\mathbf{Z},\mathbf{Z})+\sigma_{\epsilon}^{-2}k(\mathbf{Z},\mathbf{X}_{c})k(\mathbf{X}_{c},\mathbf{Z})\big{]}$
5:  $\mathbf{A}:=\text{zeros}(M,M)$
[…]
8:        $j:=i+t$
[…]
10:           $\mathbf{M}=\big{[}k(\mathbf{Z},\mathbf{Z})+\sigma_{\epsilon}^{-2}k(\mathbf{Z},\mathbf{X}_{i})k(\mathbf{X}_{i},\mathbf{Z})\big{]}$
11:           $\mathbf{L}_{i}:=\text{chol}(\mathbf{M})$ s.t. $\mathbf{L}_{i}\mathbf{L}_{i}^{\top}=\mathbf{M}$
12:           $\mathbf{R}:=\mathbf{L}_{i}\,\backslash\,\big{[}k(\mathbf{Z},\mathbf{X}_{i})\mathbf{y}_{i}\big{]}$
13:           $\bm{\alpha}_{i}:=\mathbf{L}_{i}^{\top}\,\backslash\,\mathbf{R}$
14:        end if
15:        $\mathbf{A}[i,j]=\mathbf{A}[j,i]=\sigma_{\epsilon}^{-2}\bm{\alpha}_{i}^{\top}\mathbf{M}^{(c)}\bm{\alpha}_{j}$
16:     end for
17:  end for
18:  $\bm{\beta}:=\mathbf{A}\,\backslash\,\text{diag}(\mathbf{A})$
19:  return $\bm{\beta}$, $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, $\{\mathbf{L}_{i}\}_{i=1}^{M}$
Algorithm 4 Aggregated Prediction for Distributed SVGP
1:  Input: local datasets $\{\mathbf{X}_{i},\mathbf{y}_{i}\}_{i=1}^{M}$, noise variance $\sigma_{\epsilon}^{2}$, test inputs $\mathbf{X}^{*}=\{\mathbf{x}_{i}^{*}\}_{i=1}^{n_{t}}$, a GP denoted by $\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime}))$, inducing inputs $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$
2:  Output: joint mean $\mu_{\mathcal{A}}(\mathbf{X}^{*})$ and joint variance $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*})$ in eq. 43
3:  get $\bm{\beta}=(\beta_{1},\ldots,\beta_{M})^{\top}$, $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, $\{\mathbf{L}_{i}\}_{i=1}^{M}$ from Algorithm 3
4:  $\mathbf{m}:=\text{zeros}(n_{t})$; $\mathbf{s}:=\text{zeros}(n_{t})$
[…]
6:     $\mathbf{m}\leftarrow\mathbf{m}+\beta_{i}k(\mathbf{X}^{*},\mathbf{Z})\bm{\alpha}_{i}$
7:     $\mathbf{R}:=\mathbf{L}_{i}\backslash K(\mathbf{Z},\mathbf{X}^{*})$
8:     $\mathbf{s}\leftarrow\mathbf{s}+\beta_{i}^{2}\mathbf{R}^{\top}\mathbf{R}$
9:  end for
10:  $\mu_{\mathcal{A}}(\mathbf{X}^{*}):=\sigma_{\epsilon}^{-2}\mathbf{m}$
11:  $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*}):=\mathbf{s}$
12:  return $\mu_{\mathcal{A}}(\mathbf{X}^{*})$, $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*})$
Algorithm 5 Optimal Weights for Distributed Exact GP
1:  Input: local datasets $\{\mathbf{X}_{i},\mathbf{y}_{i}\}_{i=1}^{M}$, noise variance $\sigma_{\epsilon}^{2}$, a GP denoted by $\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime}))$
2:  Output: optimal weights $\bm{\beta}=\{\beta_{i}\}_{i=1}^{M}$, kernel weights $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, Cholesky factors $\{\mathbf{L}_{i}\}_{i=1}^{M}$
3:  $\mathbf{X}_{c}:=\{\mathbf{x}_{i}^{(c)}\}_{i=1}^{M}$, $\mathbf{x}_{i}^{(c)}\in\mathbf{X}_{i}$ is randomly selected
4:  $\mathbf{A}:=\text{zeros}(M,M)$
[…]
7:        $j:=i+t$
8:        $\mathbf{M}^{(c)}=\big{[}k(\mathbf{X}_{i},\mathbf{X}_{j})+k(\mathbf{X}_{i},\mathbf{X}_{c})k(\mathbf{X}_{c},\mathbf{X}_{i})\big{]}$
[…]
10:           $\mathbf{L}_{i}:=\text{chol}(\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}})$ s.t. $\mathbf{L}_{i}\mathbf{L}_{i}^{\top}=\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}=k(\mathbf{X}_{i},\mathbf{X}_{i})+\sigma_{\epsilon}^{2}\mathbf{I}_{n_{i}}$
11:           $\mathbf{R}:=\mathbf{L}_{i}\,\backslash\,\mathbf{y}_{i}$
12:           $\bm{\alpha}_{i}:=\mathbf{L}_{i}^{\top}\,\backslash\,\mathbf{R}$
13:        end if
14:        $\mathbf{A}[i,j]=\mathbf{A}[j,i]=\bm{\alpha}_{i}^{\top}\mathbf{M}^{(c)}\bm{\alpha}_{j}$
15:     end for
16:  end for
17:  $\bm{\beta}:=\mathbf{A}\,\backslash\,\text{diag}(\mathbf{A})$
18:  return $\bm{\beta}$, $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, $\{\mathbf{L}_{i}\}_{i=1}^{M}$
Algorithm 6 Aggregated Prediction for Distributed Exact GP
1:  Input: local datasets $\{\mathbf{X}_{i},\mathbf{y}_{i}\}_{i=1}^{M}$, noise variance $\sigma_{\epsilon}^{2}$, test inputs $\mathbf{X}^{*}=\{\mathbf{x}_{i}^{*}\}_{i=1}^{n_{t}}$, a GP denoted by $\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime}))$
2:  Output: joint mean $\mu_{\mathcal{A}}(\mathbf{X}^{*})$ and joint variance $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*})$ in eq. 45
3:  get $\bm{\beta}=(\beta_{1},\ldots,\beta_{M})^{\top}$, $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, $\{\mathbf{L}_{i}\}_{i=1}^{M}$ from Algorithm 5
4:  $\mathbf{m}:=\text{zeros}(n_{t})$; $\mathbf{s}:=\text{zeros}(n_{t})$
[…]
6:     $\mathbf{m}\leftarrow\mathbf{m}+\beta_{i}k(\mathbf{X}^{*},\mathbf{X}_{i})\bm{\alpha}_{i}$
7:     $\mathbf{R}:=\mathbf{L}_{i}\backslash k(\mathbf{X}_{i},\mathbf{X}^{*})$
8:     $\mathbf{s}\leftarrow\mathbf{s}+\beta_{i}^{2}\mathbf{R}^{\top}\mathbf{R}$
9:  end for
10:  $\mu_{\mathcal{A}}(\mathbf{X}^{*}):=\mathbf{m}$
11:  $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*}):=\mathbf{s}$
12:  return $\mu_{\mathcal{A}}(\mathbf{X}^{*})$, $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*})$
Appendix C Experiment
C.1 Boxplots of the hyperparameter estimates
Figure 8: Boxplots of the optimized hyperparemeter estimates on the $n=10^{4}$ dataset.
C.2 Boxplots of the metrics
Figure 9: RMSE of the aggregation models for Exact GP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. Blue dotted line represents the mean, red solid line represents the median.
Figure 10: NLPD of the aggregation models for Exact GP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. Blue dotted line represents the mean, red solid line represents the median.
Figure 11: RMSE of the aggregation models for SVGP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, $m=128$ inducing points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. Blue dotted line represents the mean, red solid line represents the median.
Figure 12: NLPD of the aggregation models for SVGP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, $m=128$ inducing points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$. Blue dotted line represents the mean, red solid line represents the median.
C.3 Tables of the UCI datasets
Table 2: Results of the aggregation models for SVGP on the UCI datasets of size $\sim$15K. Best and worst performances are highlighted in yellow and red, respectively.
Data $M$ Metric SVGP PoE gPoE BCM rBCM grBCM opt (ours)
PoleTele 4 Time 1.588 $\pm$ 0.044 3.939 $\pm$ 0.099 3.939 $\pm$ 0.099 3.939 $\pm$ 0.099 3.939 $\pm$ 0.099 3.910 $\pm$ 0.091 3.939 $\pm$ 0.099
RMSE 16.416 $\pm$ 1.032 41.878 $\pm$ 0.315 41.878 $\pm$ 0.315 44.255 $\pm$ 0.877 41.709 $\pm$ 0.257 16.772 $\pm$ 2.289 35.596 $\pm$ 0.267
NLPD 22.096 $\pm$ 2.964 17.859 $\pm$ 0.245 7.403 $\pm$ 0.061 7.845 $\pm$ 0.174 7.374 $\pm$ 0.050 65.186 $\pm$ 17.279 6.293 $\pm$ 0.052
10 Time 1.588 $\pm$ 0.044 2.121 $\pm$ 0.005 2.121 $\pm$ 0.005 2.121 $\pm$ 0.005 2.121 $\pm$ 0.005 2.826 $\pm$ 0.008 2.121 $\pm$ 0.005
RMSE 48.678 $\pm$ 10.933 41.739 $\pm$ 0.277 41.739 $\pm$ 0.277 44.374 $\pm$ 1.066 41.709 $\pm$ 0.257 17.031 $\pm$ 1.645 30.156 $\pm$ 0.200
NLPD 189.683 $\pm$ 58.036 39.704 $\pm$ 0.529 7.455 $\pm$ 0.054 7.956 $\pm$ 0.210 7.451 $\pm$ 0.050 127.538 $\pm$ 24.461 5.387 $\pm$ 0.039
20 Time 1.588 $\pm$ 0.044 1.104 $\pm$ 0.018 1.104 $\pm$ 0.018 1.104 $\pm$ 0.018 1.104 $\pm$ 0.018 1.944 $\pm$ 0.031 1.104 $\pm$ 0.018
RMSE 48.678 $\pm$ 10.933 41.713 $\pm$ 0.263 41.713 $\pm$ 0.263 42.766 $\pm$ 1.073 41.709 $\pm$ 0.257 17.065 $\pm$ 2.667 25.617 $\pm$ 0.161
NLPD 189.683 $\pm$ 58.036 122.149 $\pm$ 1.430 9.972 $\pm$ 0.079 10.315 $\pm$ 0.347 9.972 $\pm$ 0.077 174.421 $\pm$ 54.154 6.124 $\pm$ 0.048
Elevators 4 Time 1.616 $\pm$ 0.036 4.893 $\pm$ 0.025 4.893 $\pm$ 0.025 4.893 $\pm$ 0.025 4.893 $\pm$ 0.025 4.551 $\pm$ 0.012 4.893 $\pm$ 0.025
RMSE 0.249 $\pm$ 0.004 0.253 $\pm$ 0.005 0.253 $\pm$ 0.005 0.254 $\pm$ 0.005 0.253 $\pm$ 0.005 0.255 $\pm$ 0.005 0.253 $\pm$ 0.005
NLPD 0.846 $\pm$ 0.002 0.792 $\pm$ 0.002 0.822 $\pm$ 0.002 0.822 $\pm$ 0.002 0.822 $\pm$ 0.002 0.815 $\pm$ 0.002 0.822 $\pm$ 0.002
10 Time 1.616 $\pm$ 0.036 2.263 $\pm$ 0.008 2.263 $\pm$ 0.008 2.263 $\pm$ 0.008 2.263 $\pm$ 0.008 3.058 $\pm$ 0.008 2.263 $\pm$ 0.008
RMSE 0.255 $\pm$ 0.006 0.253 $\pm$ 0.005 0.253 $\pm$ 0.005 0.324 $\pm$ 0.040 0.253 $\pm$ 0.005 0.259 $\pm$ 0.004 0.253 $\pm$ 0.005
NLPD 0.847 $\pm$ 0.002 0.786 $\pm$ 0.002 0.822 $\pm$ 0.002 0.851 $\pm$ 0.018 0.822 $\pm$ 0.002 0.805 $\pm$ 0.002 0.822 $\pm$ 0.002
20 Time 1.616 $\pm$ 0.036 1.171 $\pm$ 0.008 1.171 $\pm$ 0.008 1.171 $\pm$ 0.008 1.171 $\pm$ 0.008 2.057 $\pm$ 0.011 1.171 $\pm$ 0.008
RMSE 0.255 $\pm$ 0.006 0.254 $\pm$ 0.005 0.254 $\pm$ 0.005 0.554 $\pm$ 0.070 0.253 $\pm$ 0.005 0.261 $\pm$ 0.005 0.254 $\pm$ 0.005
NLPD 0.847 $\pm$ 0.002 0.784 $\pm$ 0.002 0.822 $\pm$ 0.001 0.986 $\pm$ 0.053 0.821 $\pm$ 0.001 0.799 $\pm$ 0.002 0.822 $\pm$ 0.001
Table 3: Results of the aggregation models for SVGP on the UCI datasets of size $\sim$45K. Best and worst performances are highlighted in yellow and red, respectively.
Data $M$ Metric SVGP PoE gPoE BCM rBCM grBCM opt (ours)
Kin40k 8 Time 1.821 ± 0.192 5.077 ± 0.110 5.077 ± 0.110 5.077 ± 0.110 5.077 ± 0.110 7.513 ± 0.158 5.077 ± 0.110
RMSE 0.391 ± 0.030 1.006 ± 0.007 1.006 ± 0.007 1.510 ± 0.128 0.997 ± 0.007 0.471 ± 0.016 0.906 ± 0.007
NLPD 0.979 ± 0.014 1.440 ± 0.009 1.469 ± 0.005 1.873 ± 0.120 1.463 ± 0.005 0.971 ± 0.011 1.322 ± 0.004
16 Time 1.821 ± 0.192 2.559 ± 0.014 2.559 ± 0.014 2.559 ± 0.014 2.559 ± 0.014 4.425 ± 0.019 2.559 ± 0.014
RMSE 1.233 ± 0.300 1.008 ± 0.009 1.008 ± 0.009 2.639 ± 0.384 0.997 ± 0.007 0.483 ± 0.010 0.816 ± 0.008
NLPD 1.735 ± 0.269 1.455 ± 0.013 1.452 ± 0.006 3.591 ± 0.765 1.444 ± 0.004 0.929 ± 0.007 1.176 ± 0.005
64 Time 1.821 ± 0.192 1.062 ± 0.012 1.062 ± 0.012 1.062 ± 0.012 1.062 ± 0.012 1.305 ± 0.009 1.062 ± 0.012
RMSE 1.233 ± 0.300 0.997 ± 0.007 0.997 ± 0.007 1.902 ± 0.554 0.997 ± 0.007 0.611 ± 0.038 0.727 ± 0.005
NLPD 1.735 ± 0.269 1.448 ± 0.010 1.469 ± 0.005 2.351 ± 0.715 1.469 ± 0.005 1.011 ± 0.033 1.071 ± 0.003
KEGG 8 Time 1.818 ± 0.256 6.211 ± 0.068 6.211 ± 0.068 6.211 ± 0.068 6.211 ± 0.068 9.227 ± 0.085 6.211 ± 0.068
RMSE 1.152 ± 0.162 1.408 ± 0.034 1.408 ± 0.034 2.527 ± 0.910 1.378 ± 0.008 1.096 ± 0.218 1.267 ± 0.031
NLPD 1.572 ± 0.151 1.969 ± 0.061 1.769 ± 0.021 2.900 ± 1.259 1.750 ± 0.008 1.567 ± 0.280 1.592 ± 0.019
16 Time 1.818 ± 0.256 3.449 ± 0.041 3.449 ± 0.041 3.449 ± 0.041 3.449 ± 0.041 5.407 ± 0.039 3.449 ± 0.041
RMSE 1.401 ± 0.062 1.455 ± 0.043 1.455 ± 0.043 7.284 ± 2.053 1.378 ± 0.008 1.355 ± 0.383 1.178 ± 0.035
NLPD 1.841 ± 0.071 2.147 ± 0.087 1.797 ± 0.027 14.708 ± 7.131 1.745 ± 0.004 2.072 ± 0.735 1.455 ± 0.022
64 Time 1.818 ± 0.256 1.095 ± 0.011 1.095 ± 0.011 1.095 ± 0.011 1.095 ± 0.011 2.075 ± 0.015 1.095 ± 0.011
RMSE 1.401 ± 0.062 1.386 ± 0.010 1.386 ± 0.010 9.475 ± 2.704 1.378 ± 0.008 1.250 ± 0.210 1.011 ± 0.007
NLPD 1.841 ± 0.071 2.088 ± 0.021 1.762 ± 0.004 21.055 ± 11.376 1.757 ± 0.006 1.882 ± 0.346 1.285 ± 0.003
Protein 8 Time 1.876 ± 0.181 6.113 ± 0.103 6.113 ± 0.103 6.113 ± 0.103 6.113 ± 0.103 8.637 ± 0.136 6.113 ± 0.103
RMSE 0.773 ± 0.003 0.773 ± 0.003 0.773 ± 0.003 0.783 ± 0.013 0.773 ± 0.003 0.773 ± 0.003 0.773 ± 0.003
NLPD 1.222 ± 0.002 1.176 ± 0.003 1.279 ± 0.002 1.285 ± 0.008 1.279 ± 0.002 1.206 ± 0.004 1.279 ± 0.002
16 Time 1.876 ± 0.181 3.484 ± 0.034 3.484 ± 0.034 3.484 ± 0.034 3.484 ± 0.034 5.427 ± 0.052 3.484 ± 0.034
RMSE 0.774 ± 0.003 0.773 ± 0.003 0.773 ± 0.003 0.795 ± 0.022 0.773 ± 0.003 0.774 ± 0.003 0.773 ± 0.003
NLPD 1.222 ± 0.002 1.171 ± 0.003 1.279 ± 0.002 1.293 ± 0.013 1.279 ± 0.002 1.188 ± 0.005 1.279 ± 0.002
64 Time 1.876 ± 0.181 1.120 ± 0.024 1.120 ± 0.024 1.120 ± 0.024 1.120 ± 0.024 2.155 ± 0.060 1.120 ± 0.024
RMSE 0.774 ± 0.003 0.773 ± 0.003 0.773 ± 0.003 0.893 ± 0.100 0.773 ± 0.003 0.776 ± 0.008 0.773 ± 0.003
NLPD 1.222 ± 0.002 1.168 ± 0.003 1.278 ± 0.002 1.360 ± 0.070 1.278 ± 0.002 1.175 ± 0.008 1.278 ± 0.002

### v2

@@
 Abstract
-GP models have received increasingly attentions in recent years due to their superb prediction accuracy and modeling flexibility.
+GP models have received increasing attention in recent years due to their superb prediction accuracy and modeling flexibility.
 To address the computational burdens of GP models for large-scale datasets, distributed learning for GPs are often adopted.
-Current aggregation models for distributed GPs are not time-efficient when incorporating correlations between GP experts.
+Current aggregation models for distributed GPs is not time-efficient when incorporating correlations between GP experts.
 In this work, we propose a novel approach for aggregated prediction in distributed GPs.
@@
 As demonstrated by empirical studies, the proposed approach results in more stable predictions in less time than state-of-the-art consistent aggregation models.
+Keywords: Distributed Gaussian processes; Optimized combination technique; Aggregation models; Inducing points.
 1 Introduction
-Gaussian processes (GPs) (Rasmussen,, 2003) are a powerful tool for modeling and inference in various areas of machine learning, such as regression (O’Hagan,, 1978; Bishop et al.,, 1995; MacKay et al.,, 2003), classification (Kuss et al.,, 2005; Nickisch and Rasmussen,, 2008; Hensman et al.,, 2015), forecasting (Girard et al.,, 2002; Roberts et al.,, 2013), signal processing (Liutkus et al.,, 2011; Sarkka et al.,, 2013), Bayesian optimization (Snoek et al.,, 2012; Frazier,, 2018), and robotics and control (Deisenroth et al.,, 2013; Mukadam et al.,, 2016).
+Gaussian processes (GPs) (Rasmussen, 2003) are a powerful tool for modeling and inference in various areas of machine learning, such as regression (O’Hagan, 1978; Bishop, 1995), classification (Kuss et al., 2005), signal processing (Liutkus et al., 2011), and robotics and control (Deisenroth et al., 2013).
 Despite their strengths, GPs face significant challenges when applied to large-scale datasets due to the computational burden of inverting large covariance matrices.
 To address these challenges, numerous efficient algorithms have been proposed.
-Sparse approximations using $m$ inducing points (Quinonero-Candela and Rasmussen,, 2005; Titsias,, 2009; Hensman et al.,, 2013) can reduce the computational complexity from $\mathcal{O}(n^{3})$ to $\mathcal{O}(m^{2}n)$ for a dataset of size $n$.
-Wilson and Nickisch, (2015) employed structured kernel interpolation combined with Kronecker and Toeplitz algebra to handle a large number of inducing points.
-Hartikainen and Särkkä, (2010); Grigorievskiy et al., (2017); Nickisch et al., (2018) reformulated GP regression as linear-Gaussian state space models, solving them with classical Kalman filtering theory.
-Katzfuss and Guinness, (2021); Katzfuss et al., (2022) presented a general Vecchia framework for GP predictions linear computational complexity in the total number of datasets.
-Chen et al., (2022); Chen and Tuo, (2022); Ding et al., (2024) represented the kernels via a sparse linear transformation to help reduce complexity.
+Sparse approximations using $m$ inducing points (Titsias, 2009) can reduce the computational complexity from $\mathcal{O}(n^{3})$ to $\mathcal{O}(m^{2}n)$ for a dataset of size $n$.
+Wilson and Nickisch (2015) employed structured kernel interpolation combined with Kronecker and Toeplitz algebra to handle a large number of inducing points.
+Hartikainen and Särkkä (2010) reformulated GP regression as linear-Gaussian state space models, solving them with classical Kalman filtering theory.
+Katzfuss and Guinness (2021) presented a general Vecchia framework for GP predictions linear computational complexity in the total number of datasets.
+Chen et al.
+(2022) represented the kernels via a sparse linear transformation to help reduce complexity.
+More recently, Lin et al.
+(2023, 2024) showed that a stochastic dual descent algorithm can efficiently yield competitive GP predictions.
 However, those approximation methods require at least $\mathcal{O}(n)$ time, which makes them impractical for very large $n$.
 Another direction to address the computational issue is distributed learning, which involves distributing computations across multiple units, often referred to as nodes or experts.
-In distributed learning (Dean et al.,, 2012), the dataset is partitioned and processed in parallel, with results aggregated to form the final model.
+In distributed learning (Dean et al., 2012), the dataset is partitioned and processed in parallel, with results aggregated to form the final model.
 By applying the previously mentioned GP approximation methods in distributed learning setup across $M$ units, the time complexity can be reduced to $\mathcal{O}(n/M)$.
-Recent aggregation models for distributed GPs (Tresp,, 2000; Hinton,, 2002; Cao and Fleet,, 2014; Deisenroth and Ng,, 2015) rely on independence assumptions and cannot offer consistent predictions, meaning the aggregated predictive distribution does not converge to the true underlying predictive distribution as the training size $n$ increases to infinity.
-To overcome this inconsistency, Rullière et al., (2018) introduced the nested point-wise aggregation of experts (NPAE), which incorporates covariances between the experts’ predictions but is very time-consuming.
-Liu et al., (2018) proposed generalized robust Bayesian committee machine (grBCM), an efficient and consistent algorithm for distributed GPs.
+Recent aggregation models for distributed GPs (Deisenroth and Ng, 2015) rely on independence assumptions and cannot offer consistent predictions, meaning the aggregated predictive distribution does not converge to the true underlying predictive distribution as the training size $n$ increases to infinity.
+To overcome this inconsistency, Rullière et al.
+(2018) introduced the nested point-wise aggregation of experts (NPAE), which incorporates covariances between the experts’ predictions but is very time-consuming.
+Liu et al.
+(2018) proposed generalized robust Bayesian committee machine (grBCM), an efficient and consistent algorithm for distributed GPs.
 In this work, we introduce a novel aggregation model for distributed GP to improve time complexity.
 Our methodology is based on GP regressions using the optimized combination technique (OptiCom) introduced in Section 3.
-The proposed method considers the correlations between the experts to ensure the consistency.
+The proposed method considers the correlations between the experts to improve the consistency of the aggregated predictions.
 It is also less time-consuming than NPAE and grBCM when the number of GP experts $M$ is not very large.
@@
 2 Background
-This section presents the background of this paper.
-We begin by introducing GP regression, GP training and sparse variational GP in Section 2.1.
-Subsequently, we outline the methods of the training and aggregated prediction for distributed GPs in Section 2.2.
+In this section, we provide the necessary background for understanding the proposed methodology.
+We begin with a review of Gaussian processes in Section 2.1, covering standard GP regression in Section 2.1.1, training procedures in Section 2.1.2, and sparse variational GP approximations in Section 2.1.3.
+We then introduce distributed GPs in Section 2.2, covering both training strategies in Section 2.2.1 and existing aggregation methods in Section 2.2.2.
 2.1 GPs
@@
 A Gaussian process (GP) is a collection of random variables, any finite number of which has a multivariate normal distribution.
-A GP is completely defined by a mean function $\mu(\cdot)$ and a covariance (or kernel) function $k(\cdot,\cdot^{\prime})$:
-$f(\cdot)\sim\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime})),$ (1)
-where $\mu(\cdot)=\mathbb{E}[f(\cdot)]$ and $k(\cdot,\cdot^{\prime})=\mathbb{E}[(f(\cdot)-\mu(\cdot))(f(\cdot^{\prime})-\mu(\cdot^{\prime}))]=\mathrm{cov}(f(\cdot),f(\cdot^{\prime}))$.
+It is defined by a mean function $\mu(\cdot)$ and a kernel function $k(\cdot,\cdot)$: $f(\cdot)\sim\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot))$.
 Suppose we have a set of training points $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ and the observations $\mathbf{y}=(y_{1},\ldots,y_{n})^{\top}$ where $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ with the i.i.d.
 noise $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a latent function.
-The GP regression imposes a GP prior over the latent function as $f(\cdot)\sim\mathcal{GP}(\mu(\cdot),k(\cdot,\cdot^{\prime}))$.
-By the definition of GP, we would obtain a $n$-dimensional multivariate Gaussian random variable $\mathbf{f}$ by evaluating the GP $f(\cdot)$ at the set of points $\mathbf{X}=\{\mathbf{x}\}_{i=1}^{n}$,
-$\mathbf{f}\sim\mathcal{N}(\bm{\mu}_{\mathbf{f}},\mathbf{K}_{\mathbf{f}\mathbf{f}}),$ (2)
-where $\mathbf{f}=\big{(}f(\mathbf{x}_{1}),\cdots,f(\mathbf{x}_{n})\big{)}^{\top}$, $\bm{\mu}_{\mathbf{f}}=\big{(}\mu(\mathbf{x}_{1}),\cdots,\mu(\mathbf{x}_{n})\big{)}^{\top}$ and $\mathbf{K}_{\mathbf{f}\mathbf{f}}=\big{[}k(\mathbf{x}_{i},\mathbf{x}_{j})\big{]}_{i,j=1}^{n}$ are the mean and covariance matrix obtained by evaluating at the location points $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$.
-In this work, we assume $f$ is a zero-mean GP, i.e., $\mu(\cdot)=0$, then the posterior distribution takes the form
-$p(f|\mathbf{y})=\mathcal{N}\big{(}\mu_{f|\mathbf{y}},k_{f|\mathbf{y}}\big{)},$ (3)
-with
+The GP regression imposes a GP prior over the latent function, and we have $\mathbf{f}=f(\mathbf{X})\sim\mathcal{N}(\bm{\mu}_{\mathbf{f}},\mathbf{K}_{\mathbf{f}\mathbf{f}})$ where $\bm{\mu}_{\mathbf{f}}=\big(\mu(\mathbf{x}_{1}),\cdots,\mu(\mathbf{x}_{n})\big)^{\top}$ and $\mathbf{K}_{\mathbf{f}\mathbf{f}}=\big[k(\mathbf{x}_{i},\mathbf{x}_{j})\big]_{i,j=1}^{n}$.
+In this work, we assume $f$ is a zero-mean GP, i.e., $\mu(\cdot)=0$, the posterior distribution $p(f|\mathbf{y})=\mathcal{N}\big(\mu_{f|\mathbf{y}},k_{f|\mathbf{y}}\big)$ takes the form
 $\mu_{f|\mathbf{y}}(\cdot)=\mathbf{K}_{(\cdot)\mathbf{f}}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\mathbf{y},$ (1a)
 $k_{f|\mathbf{y}}(\cdot,\cdot^{\prime})=k(\cdot,\cdot^{\prime})-\mathbf{K}_{(\cdot)\mathbf{f}}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\mathbf{K}_{\mathbf{f}(\cdot^{\prime})},$ (1b)
-where $\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}=\mathbf{K}_{\mathbf{f}\mathbf{f}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n}$ is the covariance matrix of all data points $\mathbf{X}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ with diagonal observation noise, $\sigma_{\epsilon}^{2}>0$ is the noise variance, $\mathbf{I}_{n}$ is a $n\times n$ identity matrix, $\mathbf{K}_{(\cdot)\mathbf{f}}=k(\cdot,\mathbf{X})=\big{[}k(\cdot,\mathbf{x}_{i})\big{]}_{i=1}^{i=n}=k(\mathbf{X},\cdot)^{\top}=\mathbf{K}_{\mathbf{f}(\cdot)}^{\top}$ is the cross-covariance matrix.
+where $\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}=\mathbf{K}_{\mathbf{f}\mathbf{f}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n}$ is the covariance matrix of all training points $\mathbf{X}$ with diagonal observation noise, $\sigma_{\epsilon}^{2}>0$ is the noise variance, $\mathbf{I}_{n}$ is a $n\times n$ identity matrix, $\mathbf{K}_{(\cdot)\mathbf{f}}=k(\cdot,\mathbf{X})=\big[k(\cdot,\mathbf{x}_{i})\big]_{i=1}^{i=n}=k(\mathbf{X},\cdot)^{\top}=\mathbf{K}_{\mathbf{f}(\cdot)}^{\top}$ is the cross-covariance matrix.
 2.1.2 GP training
-The hyperparameters $\bm{\theta}$ of GPs, which may include GP variance, kernel lengthscale, are commonly learned by maximizing the log-marginal likelihood (Jones et al.,, 1998) given by
-$\mathcal{L}(\bm{\theta})$ $=\log p(\mathbf{y}|\mathbf{f};\bm{\theta})=\log\mathcal{N}(\mathbf{y}|\mathbf{f},\sigma_{\epsilon}^{2}\mathbf{I}_{n})$ (5)
-$=-\frac{1}{2}\Big{(}\mathbf{y}^{\top}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\mathbf{y}+\log\det(\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}})+n\log(2\pi)\Big{)}.$
-A standard way of obtaining the optimized hyperparameters is by taking the derivative of each $\theta\in\bm{\theta}$ as follows:
-$\frac{\partial}{\partial\theta}\mathcal{L}(\bm{\theta})=\frac{1}{2}\mathbf{y}^{\top}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\frac{\partial\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}}{\partial\theta}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\mathbf{y}-\frac{1}{2}\text{tr}\Big{(}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\frac{\partial\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}}{\partial\theta}\Big{)}.$ (6)
+The hyperparameters $\bm{\theta}$ of GPs, which may include GP variance, kernel lengthscale, are commonly learned by maximizing the log-marginal likelihood $\mathcal{L}(\bm{\theta})=\log p(\mathbf{y}|\mathbf{f};\bm{\theta})$.
+A standard way of obtaining the optimized hyperparameters is by taking the derivative of each $\theta\in\bm{\theta}$.
 2.1.3 Sparse variational GP (SVGP)
-Exact GP regression and training requires $\mathcal{O}(n^{3})$ time complexity due to the computation of the terms $\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}$, $\log\det(\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1})$, and $\text{tr}\Big{(}\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\frac{\partial\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}}{\partial\theta}\Big{)}$ in eqs.
-4, 5 and 6, which is prohibitive when $n$ is large.
-To solve this problem, Titsias, (2009) introduced sparse variational Gaussian process (SVGP), a variational framework for sparse GPs that approximates $n$ actual observations with $m$ inducing variables ($m\ll n$) by maximizing a lower bound of the true log marginal likelihood, which can reduce the time complexity of GP regression and training to $\mathcal{O}(nm^{2})$.
-SVGP consists of $m$ inducing variables $\mathbf{u}=(u_{1},\ldots,u_{m})^{\top}\in\mathbb{R}^{m}$.
-These latent variables are values of the GP, corresponding to $m$ inducing inputs $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$, $\mathbf{z}_{i}\in\mathbb{R}^{d}$.
-We specify a Gaussian distribution over the inducing variables $q(\mathbf{u})=\mathcal{N}(\mathbf{m}_{\mathbf{u}},\mathbf{S}_{\mathbf{u}\mathbf{u}})$, then integrating out $\mathbf{u}$ with $q(\mathbf{u})$ yields
-[…]
-where
-$\mu_{f}^{\mathbf{u}}(\cdot)=\mathbf{K}_{(\cdot)\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{m}_{\mathbf{u}},$ (8a)
-$k_{f}^{\mathbf{u}}(\cdot,\cdot^{\prime})=k(\cdot,\cdot^{\prime})-\mathbf{K}_{(\cdot)\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}(\mathbf{K}_{\mathbf{u}\mathbf{u}}-\mathbf{S}_{\mathbf{u}\mathbf{u}})\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}(\cdot^{\prime})}.$ (8b)
-Here $\mathbf{K}_{(\cdot)\mathbf{u}}=\big{[}k(\cdot,\mathbf{z}_{i})\big{]}_{i=1}^{m}=\mathbf{K}_{\mathbf{u}(\cdot)}^{\top}$, and $\mathbf{K}_{\mathbf{u}\mathbf{u}}=\big{[}k(\mathbf{z}_{i},\mathbf{z}_{j})\big{]}_{i,j=1}^{m}$.
-The goal is to approximate the exact posterior $p(\mathbf{f},\mathbf{u}|\mathbf{y})$ by an variational distribution $q(\mathbf{f},\mathbf{u})$, so we minimize the Kullback-Leibler (KL) divergence $\text{KL}\big{[}q(\mathbf{f},\mathbf{u})\|p(\mathbf{f},\mathbf{u}|\mathbf{y})\big{]}$, which is equivalent to maximize the evidence lower bound (ELBO) defined as follows:
-$\mathcal{L}_{f}:=\mathbb{E}_{q(\mathbf{f},\mathbf{u})}\left[\log\frac{p(\mathbf{y},\mathbf{f},\mathbf{u})}{q(\mathbf{f},\mathbf{u})}\right]=\log p(\mathbf{y})-\text{KL}\big{[}q(\mathbf{f},\mathbf{u})\|p(\mathbf{f},\mathbf{u}|\mathbf{y})\big{]}.$ (9)
-Under a Gaussian likelihood of the form $p(\mathbf{y}|\mathbf{f})=\mathcal{N}(\mathbf{y}|\mathbf{f},\sigma_{\epsilon}^{2}\mathbf{I}_{n})$, the optimal variational distribution $\hat{q}(\mathbf{u})=\mathcal{N}(\hat{\mathbf{m}}_{\mathbf{u}},\hat{\mathbf{S}}_{\mathbf{u}\mathbf{u}})$ can be obtained by
-$\hat{\mathbf{m}}_{\mathbf{u}}=\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{u}}\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}}\mathbf{y},$ (10a)
-$\hat{\mathbf{S}}_{\mathbf{u}\mathbf{u}}=\mathbf{K}_{\mathbf{u}\mathbf{u}}\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{u}},$ (10b)
-where $\mathbf{K}_{\mathbf{u}\mathbf{f}}=\big{[}k(\mathbf{z}_{i},\mathbf{x}_{j})\big{]}_{i,j=1}^{m,n}=\mathbf{K}_{\mathbf{f}\mathbf{u}}^{\top}$, and $\mathbf{M}_{\mathbf{u}\mathbf{u}}=\big{[}\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}}\big{]}$.
-The corresponding optimal ELBO then becomes
-^$\hat{\mathcal{L}}_{f}=\log\mathcal{N}(\mathbf{y}|\bm{\mu}_{\mathbf{f}},\mathbf{K}_{\mathbf{f}\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n})-\frac{1}{2}\sigma_{\epsilon}^{-2}\text{tr}(\mathbf{K}_{\mathbf{f}\mathbf{f}}-\mathbf{K}_{\mathbf{f}\mathbf{u}}\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}}).$ (11)
-Therefore, eq. 10 can lead to the predictive distribution $\hat{q}(f)=\mathcal{N}(\hat{\mu}_{f}^{\mathbf{u}},\hat{k}_{f}^{\mathbf{u}})$ with
+Exact GP regression and training as described in Section 2.1.1 and Section 2.1.2 require $\mathcal{O}(n^{3})$ time due to the computation of the terms $\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}$, $\log\det(\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1})$, and $\text{tr}\Big(\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}^{-1}\frac{\partial\widetilde{\mathbf{K}}_{\mathbf{f}\mathbf{f}}}{\partial\theta}\Big)$, which is prohibitive when $n$ is large.
+To solve this problem, Titsias (2009) introduced sparse variational Gaussian process (SVGP), a variational framework for sparse GPs using $m\ll n$ inducing variables $\mathbf{u}=(u_{1},\ldots,u_{m})^{\top}$ with $m$ associated inducing inputs $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$, $\mathbf{z}_{i}\in\mathbb{R}^{d}$, which reduces the time complexity to $\mathcal{O}(nm^{2})$.
+The variational distribution $q(\mathbf{u})=\mathcal{N}(\mathbf{m}_{\mathbf{u}},\mathbf{S}_{\mathbf{u}\mathbf{u}})$ yields the predictive distribution $\hat{q}(f)=\mathcal{N}(\hat{\mu}_{f}^{\mathbf{u}},\hat{k}_{f}^{\mathbf{u}})$ as follows:
 $\hat{\mu}_{f}^{\mathbf{u}}(\cdot)=\sigma_{\epsilon}^{-2}\mathbf{K}_{(\cdot)\mathbf{u}}\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}}\mathbf{y},$ (2a)
-$\hat{k}_{f}^{\mathbf{u}}(\cdot,\cdot^{\prime})=k(\cdot,\cdot^{\prime})-\mathbf{K}_{(\cdot)\mathbf{u}}(\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}-\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1})\mathbf{K}_{\mathbf{u}(\cdot^{\prime})}.$ (12b)
+$\hat{k}_{f}^{\mathbf{u}}(\cdot,\cdot^{\prime})=k(\cdot,\cdot^{\prime})-\mathbf{K}_{(\cdot)\mathbf{u}}(\mathbf{K}_{\mathbf{u}\mathbf{u}}^{-1}-\mathbf{M}_{\mathbf{u}\mathbf{u}}^{-1})\mathbf{K}_{\mathbf{u}(\cdot^{\prime})},$ (2b)
+where $\mathbf{K}_{(\cdot)\mathbf{u}}=\big[k(\cdot,\mathbf{z}_{i})\big]_{i=1}^{m}=\mathbf{K}_{\mathbf{u}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{u}\mathbf{f}}=\big[k(\mathbf{z}_{i},\mathbf{x}_{j})\big]_{i,j=1}^{m,n}=\mathbf{K}_{\mathbf{f}\mathbf{u}}^{\top}$, and $\mathbf{M}_{\mathbf{u}\mathbf{u}}=\big[\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}}\big]$.
+The SVGP framework introduced here is essential for our distributed learning methodology in Section 4.
+By employing SVGP approximations at the local expert level before aggregation, we achieve a double reduction in computational cost: first, inducing points approximation reduces the cost from $\mathcal{O}(n^{3})$ to $\mathcal{O}(nm^{2})$; second, data parallelization across $M$ experts further reduces the time complexity to $\mathcal{O}(nm^{2}/M)$.
 2.2 Distributed GP
+While SVGP reduces complexity from $\mathcal{O}(n^{3})$ to $\mathcal{O}(nm^{2})$, this can still be impractical when $n$ is extremely large such as billions of observations.
+Distributed learning provides a natural solution by partitioning the data across $M$ computational units (experts) and processing them in parallel, further reducing per-unit complexity to $O(nm^{2}/M)$.
+This section introduces the distributed GP framework that motivates our main contributions.
+Section 2.2.1 describes how local experts are trained independently on partitioned local datasets.
+Section 2.2.2 reviews existing aggregation methods for distributed GPs and highlights their strengths and limitations.
 2.2.1 Distributed GP training
-Distibuted GPs (Ng and Deisenroth,, 2014; Deisenroth and Ng,, 2015; Liu et al.,, 2018; Yin et al.,, 2020) enable the parallel training of GPs across different experts to harness greater computing power and accelerate the training process for large-scale datasets.
-Assuming the same model as in Section 2.1.1, where $f(\cdot)\sim\mathcal{GP}(0,k(\cdot,\cdot^{\prime}))$ and $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ with the i.i.d.
-noise $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$.
-We first partition the global training dataset $\mathcal{D}:=\{\mathbf{X},\mathbf{y}\}$ of size $n$ into $M$ local datasets $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$, each of size $n_{i}$ for $i=1,\ldots,M$.
+Distributed GPs (Deisenroth and Ng, 2015) enable the parallel training of GPs across different experts to accelerate the computing process for large-scale datasets.
+Note that distributed GPs are not inherently private because they involve data sharing or centralized aggregation, whereas federated GPs preserve privacy by keeping raw data local and only exchanging model updates.
+Assuming the same model as in Section 2.1.1, we first partition the global training dataset $\mathcal{D}:=\{\mathbf{X},\mathbf{y}\}$ of size $n$ into $M$ local datasets $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$, each of size $n_{i}$ for $i=1,\ldots,M$.
 Each local dataset $\mathcal{D}_{i}$ corresponds to an expert $\mathcal{M}_{i}$.
 Clearly, $n=\sum_{i=1}^{M}n_{i}$.
-We suppose that all local datasets share the same hyperparameters $\bm{\theta}$, as defined in Section 2.1.2.
-The global objective for distribued GP training is to maximize the global marginal log-likelihood $\log p(\mathbf{y}|\mathbf{f},\bm{\theta})$, where $\mathbf{f}=f(\mathbf{X})$.
-Assuming independence among all the experts $\{\mathcal{M}_{i}\}_{i=1}^{M}$, the global marginal log-likelihood is approximated as follows
+We suppose that all local datasets share the same hyperparameters $\bm{\theta}$.
+The global objective for distributed GP training is to maximize the global marginal log-likelihood $\log p(\mathbf{y}|\mathbf{f},\bm{\theta})$, where $\mathbf{f}=f(\mathbf{X})$.
+Assuming independence among all the experts $\{\mathcal{M}_{i}\}_{i=1}^{M}$, the global marginal log-likelihood is approximated as:
 $\log p(\mathbf{y}|\mathbf{f};\bm{\theta})$ $\approx\sum_{i=1}^{M}\log p_{i}(\mathbf{y}_{i}|\mathbf{f}_{i};\bm{\theta})=-\frac{1}{2}\sum_{i=1}^{M}\Big(\mathbf{y}_{i}^{\top}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{y}_{i}+\log\det(\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}})+n\log(2\pi)\Big)$ (3)
 where $p_{i}(\mathbf{y}_{i}|\mathbf{f}_{i},\bm{\theta})\sim\mathcal{N}(\mathbf{0},\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}})$ is the local marginal likelihood of the $i$-th model $\mathcal{M}_{i}$ with $\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}=\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{i}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n_{i}}$ and $\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{i}}=k(\mathbf{X}_{i},\mathbf{X}_{i})\in\mathbb{R}^{n_{i}\times n_{i}}$, $\mathbf{f}_{i}=f(\mathbf{X}_{i})\in\mathbb{R}^{n_{i}}$.
-The factorized approximation of the log-likilihood in eq.
-13 degenerates the full covariance matrix $\mathbf{K}_{\mathbf{f}\mathbf{f}}=k(\mathbf{X},\mathbf{X})$ into a diagonal block matrix $\text{diag}[\mathbf{K}_{\mathbf{f}_{1},\mathbf{f}_{1}},\ldots,\mathbf{K}_{\mathbf{f}_{M},\mathbf{f}_{M}}]$, reducing the time complexity to $\mathcal{O}(n_{i}^{3})$.
-This factorized training method (Ng and Deisenroth,, 2014; Deisenroth and Ng,, 2015), termed as FACT, is one common approach to distributed GP training.
-In addition to FACT, various optimization methods have been proposed.
-Federated Averaging (FedAvg) (McMahan et al.,, 2017) is a popular aggregation strategy that trains a global model by averaging updates from multiple local datasets, FedProx (Li et al.,, 2020) generalizes FedAvg by adding a proximal term to address the heterogeneity in federated networks.
-The Alternating Direction of Multipliers Method (ADMM) (Boyd et al.,, 2011) reformulates the optimization problem as a nonconvex consensus problem with newly introduced local hyperparameters and the global hyperparameter.
-Proximal ADMM (pxADMM) (Hong et al.,, 2016) aims to reduce communication overhead and computational time, and Xie et al., (2019) derive a closed-form solution for distributed GP training based on pxADMM.
+The factorized approximation of the log-likelihood in eq.
+3 (FACT) approximates the full covariance matrix $\mathbf{K}_{\mathbf{f}\mathbf{f}}=k(\mathbf{X},\mathbf{X})$ by a diagonal block matrix $\text{diag}[\mathbf{K}_{\mathbf{f}_{1},\mathbf{f}_{1}},\ldots,\mathbf{K}_{\mathbf{f}_{M},\mathbf{f}_{M}}]$, reducing the time complexity to $\mathcal{O}(n_{i}^{3})$.
+Other optimization techniques include Federated Averaging (FedAvg) (McMahan et al., 2017), which averages local model updates; FedProx (Li et al., 2020), which improves FedAvg by addressing client heterogeneity; ADMM (Boyd et al., 2011), which solves a consensus optimization problem by introducing auxiliary variables and dual updates.
 2.2.2 Aggregated prediction
-Common aggregation methods for GP experts include product-of-experts (PoE) (Hinton,, 2002), generalised product-of-experts (gPoE) (Cao and Fleet,, 2014), Bayesian committee machine (BCM) (Tresp,, 2000), robust Bayesian committee machine (rBCM) (Deisenroth and Ng,, 2015), generalized robust Bayesian committee machine (grBCM) (Liu et al.,, 2018), and nested pointwise aggregation of experts (NPAE) (Rullière et al.,, 2018).
-The joint mean $\mu_{\mathcal{A}}$ and joint precision $\sigma_{\mathcal{A}}^{-2}$ for PoE, gPoE, BCM, rBCM are given by
+After training local experts, we aggregate local predictions to form a global prediction.
+Let $\mu_{i}(\cdot)$ and $\sigma_{i}^{2}(\cdot)$ denote the predictive mean and variance from expert $\mathcal{M}_{i}$.
+Common aggregation methods for GP experts include product-of-experts (PoE) (Hinton, 2002), generalised product-of-experts (gPoE) (Cao and Fleet, 2014), Bayesian committee machine (BCM) (Tresp, 2000), robust Bayesian committee machine (rBCM) (Deisenroth and Ng, 2015), generalized robust Bayesian committee machine (grBCM) (Liu et al., 2018), and nested pointwise aggregation of experts (NPAE) (Rullière et al., 2018).
+PoE, gPoE, BCM and rBCM produce a global prediction using a weighted average of local precisions, the joint mean $\mu_{\mathcal{A}}$ and joint precision $\sigma_{\mathcal{A}}^{-2}$ are given by
 $\mu_{\mathcal{A}}(\cdot)=\sigma_{\mathcal{A}}^{2}(\cdot)\sum_{i=1}^{M}\beta_{i}\sigma_{i}^{-2}(\cdot)\mu_{i}(\cdot),$ (4a)
 $\sigma_{\mathcal{A}}^{-2}(\cdot)=\sum_{i=1}^{M}\beta_{i}\sigma_{i}^{-2}(\cdot)+(1-\sum_{i=1}^{M}\beta_{i})\sigma_{**}^{-2}(\cdot),$ (4b)
-where $\mu_{i}(\cdot)$ and $\sigma_{i}^{-2}(\cdot)$ represent the local mean and local precision of expert $\mathcal{M}_{i}$, while $\sigma_{**}^{-2}(\cdot)=k(\cdot,\cdot)+\sigma_{\epsilon}^{2}$ is the prior variance, serving as a correction term for the BCM family.
-The weights, $\beta_{i}$, are defined as follows: $\beta_{i}=1$ for PoE and BCM, $\beta_{i}=1/M$ for gPoE, and $\beta_{i}=0.5(\log\sigma_{**}^{2}(\cdot)-\log\sigma_{i}^{2}(\cdot))$ for rBCM.
+where $\sigma_{**}^{-2}(\cdot)=k(\cdot,\cdot)+\sigma_{\epsilon}^{2}$ is the prior variance, serving as a correction term for the BCM family.
+The weights, $\beta_{i}$, are defined as follows: $\beta_{i}=1$ for PoE and BCM, and $\beta_{i}=0.5(\log\sigma_{**}^{2}(\cdot)-\log\sigma_{i}^{2}(\cdot))$ for gPoE and rBCM.
 For grBCM, the $M$ experts are divided into two groups: the global expert $\mathcal{M}_{c}$, trained on dataset $\mathcal{D}_{c}=\mathcal{D}_{1}$ of size $n_{c}$, and the local experts $\{\mathcal{M}_{i}\}_{i=2}^{M}$, trained on datasets $\{\mathcal{D}_{i}\}_{i=2}^{M}$.
@@
 The weights are defined as $\beta_{2}=1$, and $\beta_{i}=0.5(\log\sigma_{c}^{2}(\cdot)-\log\sigma_{+i}^{2}(\cdot))$ for $i=3,\ldots,M$.
-NPAE leverages the covariances between experts to ensure consistent predictions. Consequently, the joint mean and joint variance are computed as follows:
+NPAE leverages the covariances between experts to ensure consistent predictions. The joint mean and joint variance are computed as follows:
 $\mu_{\mathcal{A}}(\cdot)=\mathbf{K}_{(\cdot)\mathcal{A}}\mathbf{K}_{\mathcal{A}\mathcal{A}}^{-1}\bm{\mu}_{\mathcal{A}},$ (6a)
 $\sigma_{\mathcal{A}}^{2}(\cdot)=k(\cdot,\cdot)-\mathbf{K}_{(\cdot)\mathcal{A}}\mathbf{K}_{\mathcal{A}\mathcal{A}}^{-1}\mathbf{K}_{\mathcal{A}(\cdot)}+\sigma_{\epsilon}^{2},$ (6b)
-with
-$\mathrm{cov}[\mu_{i}(\cdot),y_{*}(\cdot)]=\mathbf{K}_{(\cdot),i}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}(\cdot)},$ (17a)
-$\mathrm{cov}[\mu_{i}(\cdot),\mu_{j}(\cdot)]=\begin{cases}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{j}}\widetilde{\mathbf{K}}_{\mathbf{f}_{j}\mathbf{f}_{j}}^{-1}\mathbf{K}_{\mathbf{f}_{j}(\cdot)},&i\neq j,\\ \mathbf{K}_{(\cdot),i}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}(\cdot)},&i=j,\end{cases}$ (17b)
-where $\mu_{i}(\cdot)$ represents the local mean of expert $\mathcal{M}_{i}$, $\mathbf{K}_{(\cdot)\mathbf{f}_{i}}=k(\cdot,\mathbf{X}_{i})=\mathbf{K}_{\mathbf{f}_{i}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{j}}=k(\mathbf{X}_{i},\mathbf{X}_{j})$, $\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}=\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{i}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n_{i}}$.
-$\mathbf{K}_{\mathcal{A}(\cdot)}=\big{[}\mathrm{cov}[\mu_{i}(\cdot),y_{*}(\cdot)]\big{]}_{i=1}^{M}=\mathbf{K}_{(\cdot)\mathcal{A}}^{\top}\in\mathbb{R}^{M}$, $\mathbf{K}_{\mathcal{A}\mathcal{A}}=\big{[}\mathrm{cov}[\mu_{i}(\cdot),\mu_{j}(\cdot)]\big{]}_{i,j=1}^{M}\in\mathbb{R}^{M\times M}$, $\bm{\mu}_{\mathcal{A}}=\big{[}\mu_{1}(\cdot),\ldots,\mu_{M}(\cdot)\big{]}^{\top}\in\mathbb{R}^{M}$.
+where $\mathbf{K}_{(\cdot)\mathbf{f}_{i}}=k(\cdot,\mathbf{X}_{i})=\mathbf{K}_{\mathbf{f}_{i}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{j}}=k(\mathbf{X}_{i},\mathbf{X}_{j})$, $\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}=\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{i}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n_{i}}$.
+$\mathbf{K}_{\mathcal{A}(\cdot)}=[\mathrm{cov}[\mu_{i}(\cdot),y_{*}(\cdot)]]_{i=1}^{M}$ == $\mathbf{K}_{(\cdot)\mathcal{A}}^{\top}\in\mathbb{R}^{M}$, $\mathbf{K}_{\mathcal{A}\mathcal{A}}=\big[\mathrm{cov}[\mu_{i}(\cdot),\mu_{j}(\cdot)]\big]_{i,j=1}^{M}\in\mathbb{R}^{M\times M}$, $\bm{\mu}_{\mathcal{A}}=\big[\mu_{1}(\cdot),\ldots,\mu_{M}(\cdot)\big]^{\top}\in\mathbb{R}^{M}$.
+Among these methods, PoE, gPoE, BCM, and rBCM are computationally efficient but assume independence between experts, leading to inconsistent predictions as the dataset size grows.
+The grBCM method addresses consistency by introducing a global expert but requires training augmented models $\mathcal{M}_{i}^{+}$ for $i=2,\ldots,M$, each on datasets of size $n_{c}+n_{i}$, significantly increasing computational cost.
+NPAE accounts for correlations through the full covariance matrix $\mathbf{K}_{\mathcal{A}\mathcal{A}}$ in eq.
+6, achieving consistency but at $O(M^{3})$ cost for inverting this matrix, which becomes prohibitive for large $M$.
+These limitations motivate our approach: we seek an aggregation method that (1) incorporates expert correlations for consistency, (2) maintains computational efficiency comparable to simple baselines, and (3) naturally applies to both exact GP and SVGP settings.
+Our key insight introduced in Section 4 is to leverage the optimized combination technique (OptiCom) from sparse grid theory, which is introduced in Section 3, to derive optimal aggregation weights that address these requirements.
 3 GP regression with OptiCom
-In this section, we detail the algorithms for GP regression using the optimized combination technique.
-First, we introduce the combination technique and the optimized combination technique in Section 3.1 and Section 3.2, respectively.
-Then, we present the algorithm for applying OptiCom to GP in Section 3.3.
+This section introduces GP regression using the optimized combination technique (OptiCom).
+We first review the standard combination technique (CT) in Section 3.1, followed by OptiCom in Section 3.2, and then show how it is applied to GP regression.
 3.1 Combination technique (CT)
-Combination technique (CT) (Griebel et al.,, 1990; Hegland et al.,, 2007), first introduced in (Smolyak,, 1963), is an efficient tool for approximating the sparse grid spaces.
-When partial projection operators commute, as in interpolation with tensor product spaces, the combination technique provides the exact sparse grid solution (Griebel et al.,, 1990).
-Sparse grids of level $\eta$ and dimension $d$ are defined as (Garcke,, 2013)
+Combination technique (CT) (Griebel et al., 1990; Hegland et al., 2007), first introduced in (Smolyak, 1963), is an efficient tool for approximating the sparse grid spaces.
+When partial projection operators commute, as in interpolation with tensor product spaces, the combination technique provides the exact sparse grid solution.
+Sparse grids of level $\eta$ and dimension $d$ are defined as (Garcke, 2013)
 $\Omega_{\eta}^{d}:=\bigcup_{|\underline{l}|_{1}=\eta+d-1}\Omega_{\underline{l}}=\bigcup_{|\underline{l}|_{1}=\eta+d-1}\Omega_{l_{1}}^{1}\times\cdots\times\Omega_{l_{d}}^{1}.$ (7)
@@
 $\Omega_{l_{1}}^{1}\times\cdots\times\Omega_{l_{d}}^{1}=\{(\omega_{l_{1}},\ldots,\omega_{l_{d}})|\omega_{l_{j}}\in\Omega_{l_{j}}^{1}\;\text{for}\;j\in\{1,\ldots,d\}\}$ denotes the n-ary Cartesian product over $d$ one-dimensional set $\Omega_{l_{j}}^{1}\subset\mathbb{R}$, $j=1,\ldots,d$.
-We suppose $\Omega_{l_{j}}^{1}$ consists of $2^{l_{j}-1}$ uniformly distributed points over a fixed interval, i.e., $\Omega_{l_{j}}^{1}$ over the interval $[0,1]$ is given by
-$\Omega_{l_{j}}^{1}=\left\{\frac{i}{2^{l_{j}}}:i=1,\ldots,2^{l_{j}-1}\right\}.$ (19)
-Each set of grids $\Omega_{\underline{l}}\subset\mathbb{R}^{d}$ is associated with a piecewise $d$-linear basis function $\bm{\phi}_{\underline{l},\underline{h}}(\cdot)$, which is defined as the product of the one-dimensional basis functions:
-$\bm{\phi}_{\underline{l},\underline{h}}(\mathbf{x}):=\prod_{j=1}^{d}\phi_{l_{j},h_{j}}(x_{j}),\quad h_{j}=1,\ldots,2^{l_{j}-1}$ (20)
-where $\mathbf{x}=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$ is a $d$-dimensional point, $\underline{h}=(h_{1},\ldots,h_{d})$, $\phi_{l_{j},h_{j}}(\cdot)$ is a piecewise linear hierarchical basis shown in Figure 7(a) in Appendix and defined as follows:
-$\phi_{l_{j},h_{j}}(x)=\begin{cases}1-\lvert 2^{l_{j}}x-h_{j}\rvert,&x\in\big{[}\frac{h_{j}-1}{2^{l_{j}}},\frac{h_{j}+1}{2^{l_{j}}}\big{]},\\ 0,&\text{otherwise}.\end{cases}$ (21)
-These basis functions can define function spaces $V_{\underline{l}}$ on the grids $\Omega_{\underline{l}}$,
-$V_{\underline{l}}:=\text{span}\{\bm{\phi}_{\underline{l},\underline{h}}:h_{j}=1,\ldots,2^{l_{j}}-1,j=1,\ldots,d\}.$ (22)
-Then the hierarchical increment spaces $W_{\underline{l}}$ can be defined by
+We suppose $\Omega_{l_{j}}^{1}$ consists of $2^{l_{j}-1}$ uniformly distributed points over a fixed interval, i.e., $\Omega_{l_{j}}^{1}$ over the interval $[0,1]$ is given by $\Omega_{l_{j}}^{1}=\left\{\frac{i}{2^{l_{j}}}:i=1,\ldots,2^{l_{j}-1}\right\}$.
+Each set of grids $\Omega_{\underline{l}}\subset\mathbb{R}^{d}$ is associated with a piecewise $d$-linear basis function $\bm{\phi}_{\underline{l},\underline{h}}(\cdot)$ defined as:
+$\bm{\phi}_{\underline{l},\underline{h}}(\mathbf{x}):=\prod_{j=1}^{d}\phi_{l_{j},h_{j}}(x_{j}),\quad h_{j}=1,\ldots,2^{l_{j}-1},\quad\phi_{l_{j},h_{j}}(x)=\begin{cases}1-\lvert 2^{l_{j}}x-h_{j}\rvert,&x\in\big[\frac{h_{j}-1}{2^{l_{j}}},\frac{h_{j}+1}{2^{l_{j}}}\big],\\ 0,&\text{otherwise},\end{cases}$ (8)
+where $\mathbf{x}=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$ is a $d$-dimensional point, $\underline{h}=(h_{1},\ldots,h_{d})$, $\phi_{l_{j},h_{j}}(\cdot)$ is a piecewise linear hierarchical basis shown in Figure 7(a).
+These basis functions can define function spaces $V_{\underline{l}}:=\text{span}\{\bm{\phi}_{\underline{l},\underline{h}}:h_{j}=1,\ldots,2^{l_{j}}-1,j=1,\ldots,d\}$ on the grids $\Omega_{\underline{l}}$, then the hierarchical increment spaces $W_{\underline{l}}$ can be defined by
 $W_{\underline{l}}:=\text{span}\{\bm{\phi}_{\underline{l},\underline{h}}:\underline{h}\in\mathrm{B}_{\underline{l}}\},$ (9)
 $\mathrm{B}_{\underline{l}}:=\{\underline{h}\in\mathbb{N}^{d}:h_{j}=1,\ldots,2^{l_{j}}-1,h_{j}\text{ is odd for all }j=1,\ldots,d\}.$ (10)
-Therefore, the function spaces can be represented by the hierarchical increment spaces
-$V_{\underline{l}}=\bigoplus_{\underline{k}\;\leq\;\underline{l}}W_{\underline{k}},$ (25)
-and each function $f_{\underline{l}}\in V_{\underline{l}}$ can be uniquely represented by
-$f_{\underline{l}}(\mathbf{x})=\sum_{|\underline{l}|_{1}\leq\eta+d-1}\sum_{\underline{h}\in\mathrm{B}_{\underline{l}}}\alpha_{\underline{k},\underline{h}}\bm{\phi}_{\underline{k},\underline{h}}(\mathbf{x})$ (26)
-with hierarchical coefficients $\alpha_{\underline{l},\underline{h}}\in\mathbb{R}$.
-The combination technique linearly combines the discrete solutions $f_{\underline{l}}(\mathbf{x})$ from the partial grids $\Omega_{\underline{l}}$ according to the combination formula (see Figure 7(b) in Appendix for an example):
-${\color[rgb]{0,0,1}{f_{\eta}^{c}(\mathbf{x})=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}(-1)^{\eta+d-1-|\underline{l}|_{1}}\binom{d-1}{|\underline{l}|_{1}-\eta}f_{\underline{l}}(\mathbf{x}),}}$ (27)
-where the function $f_{\eta}^{c}$ is in the sparse gris space $V_{\eta}^{s}$ of level $\eta$ defined by
-$V_{\eta}^{s}:=\bigoplus_{|\underline{l}|_{1}\leq\eta+d-1}W_{\underline{l}}.$ (28)
+Therefore, the function spaces can be represented by the hierarchical increment spaces $V_{\underline{l}}=\bigoplus_{\underline{k}\;\leq\;\underline{l}}W_{\underline{k}}$ and each function $f_{\underline{l}}\in V_{\underline{l}}$ can be uniquely represented by $f_{\underline{l}}(\mathbf{x})=\sum_{|\underline{l}|_{1}\leq\eta+d-1}\sum_{\underline{h}\in\mathrm{B}_{\underline{l}}}\alpha_{\underline{l},\underline{h}}\bm{\phi}_{\underline{l},\underline{h}}(\mathbf{x})$ with hierarchical coefficients $\alpha_{\underline{l},\underline{h}}\in\mathbb{R}$.
+The space $\oplus_{i=1}^{n}S_{i}$ is the direct sum of the spaces $S_{1},\ldots,S_{n}$ if and only if all nonzero vectors $s_{i}\in S_{i}$ (for $i=1,\ldots,n$) are linearly independent, and $\underline{k}\leq\underline{l}$ indicates that each component of $\underline{k}$ is less than or equal to the corresponding component of $\underline{l}$.
+The combination technique linearly combines the discrete solutions $f_{\underline{l}}(\mathbf{x})$ from the partial grids $\Omega_{\underline{l}}$ according to the combination formula (see Figure 7):
+$f_{\eta}^{c}(\mathbf{x})=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}(-1)^{\eta+d-1-|\underline{l}|_{1}}\binom{d-1}{|\underline{l}|_{1}-\eta}f_{\underline{l}}(\mathbf{x}),$ (11)
+where the function $f_{\eta}^{c}$ is in the sparse grid space $V_{\eta}^{s}:=\bigoplus_{|\underline{l}|_{1}\leq\eta+d-1}W_{\underline{l}}$.
 3.2 Optimized combination technique (OptiCom)
-Optimized Combination Technique (OptiCom) (Hegland et al.,, 2002; Garcke,, 2006; Griebel et al.,, 2015) selects the “best possible” combination coefficients adaptively so that an optimal combination approximation is obtained.
+Optimized Combination Technique (OptiCom) (Hegland, 2002; Garcke, 2006) selects the “best possible” combination coefficients adaptively so that an optimal combination approximation is obtained.
 More specifically, we aim to minimize the functional
@@
 where $\mathcal{P}_{\eta}^{s}f$ denotes the projection into the sparse grid space $V_{\eta}^{s}$, $\mathcal{P}_{i}f$ denotes the projection into one of the spaces $V_{\underline{l}}$ in eq.
-27, $b$ is the number of summed terms in combination technique formula eq.
-27.
+11, $b$ is the number of summed terms in combination technique formula eq.
+11.
 Using simple expansions and formula $\langle\mathcal{P}_{\eta}^{s}f,\mathcal{P}_{i}f\rangle=\langle\mathcal{P}_{i}f,\mathcal{P}_{i}f\rangle$, we can obtain
@@
 By differentiating with respect to the combination coefficients $c_{i}$ and setting each of these derivatives to zero, we can derive the following linear system
-$\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}&\cdots&\langle\mathcal{P}_{1}f,\mathcal{P}_{b}f\rangle\\ \langle\mathcal{P}_{2}f,\mathcal{P}_{1}f\rangle&\cdots&\langle\mathcal{P}_{2}f,\mathcal{P}_{b}f\rangle\\ \vdots&\ddots&\vdots\\ \langle\mathcal{P}_{b}f,\mathcal{P}_{1}f\rangle&\cdots&\|\mathcal{P}_{b}f\|^{2}\end{bmatrix}\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{b}\end{bmatrix}=\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}\\ \|\mathcal{P}_{2}f\|^{2}\\ \vdots\\ \|\mathcal{P}_{b}f\|^{2}\end{bmatrix}.$ (31)
+$\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}&\cdots&\langle\mathcal{P}_{1}f,\mathcal{P}_{b}f\rangle\\ \langle\mathcal{P}_{2}f,\mathcal{P}_{1}f\rangle&\cdots&\langle\mathcal{P}_{2}f,\mathcal{P}_{b}f\rangle\\ \vdots&\ddots&\vdots\\ \langle\mathcal{P}_{b}f,\mathcal{P}_{1}f\rangle&\cdots&\|\mathcal{P}_{b}f\|^{2}\end{bmatrix}\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{b}\end{bmatrix}=\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}\\ \|\mathcal{P}_{2}f\|^{2}\\ \vdots\\ \|\mathcal{P}_{b}f\|^{2}\end{bmatrix}$.
+(14)
 Using the optimal coefficients $c_{i}$, the OptiCom formula for the sparse grid of level $\eta$ is then given by
-${\color[rgb]{0,0,1}{f_{\eta}^{c}(\mathbf{x})=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}c_{\underline{l}}f_{\underline{l}}(\mathbf{x}).}}$ (32)
-For the regression problem (Garcke,, 2006), we are seeking the solution of the optimization problem $f^{*}=\arg\min_{f\in V}R(f)$ with
+$f_{\eta}^{c}(\mathbf{x})=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}c_{\underline{l}}f_{\underline{l}}(\mathbf{x}).$ (15)
+For the regression problem (Garcke, 2006), we are seeking the solution of the optimization problem $f^{*}=\arg\min_{f\in V}R(f)$ with
 $R(f):=\sum_{i=1}^{n}(f(\mathbf{x}_{i})-y_{i})^{2}+\lambda\|\mathcal{S}f\|_{2}^{2},$ (16)
@@
 Suppose the inducing inputs $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$ ($\mathbf{z}_{i}\in\mathbb{R}^{d}$) are sparse grids $\Omega_{\eta}^{d}$ of level $\eta$ and dimension $d$ defined in eq.
-18, then the predictive mean and covariance in eq.
+7, then the predictive mean and covariance in eq.
 2 can be approximated in the form
@@
 where $\mathbf{u}_{\underline{l}}$ is the vector of inducing variables corresponding to the inducing inputs $\Omega_{\underline{l}}$ defined in eq.
-18, $\mathbf{K}_{(\cdot)\mathbf{u}_{\underline{l}}}=k(\cdot,\Omega_{\underline{l}})=\mathbf{K}_{\mathbf{u}_{\underline{l}}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}=k(\Omega_{\underline{l}},\Omega_{\underline{l}})$, and $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}=\big{[}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}_{\underline{l}}}\big{]}$.
+7, $\mathbf{K}_{(\cdot)\mathbf{u}_{\underline{l}}}=k(\cdot,\Omega_{\underline{l}})=\mathbf{K}_{\mathbf{u}_{\underline{l}}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}=k(\Omega_{\underline{l}},\Omega_{\underline{l}})$, and $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}=\big[\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}_{\underline{l}}}\big]$.
+To ensure that the predictive covariance remains consistent with that of the full GP, we use the CT coefficients rather than those from OptiCom in the covariance computation.
 3.3.2 Optimal coefficients
-To compute the optimal coefficients $c_{\underline{l}}$ in GP regression, we need to compute the inner product with respect to the reproducing kernel Hilbert space (RKHS) (Aronszajn,, 1950).
-Let $\mathcal{H}$ be a RKHS endowed with an inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$, then the inner product between functions $f(\cdot)=\sum_{i=1}^{n}\alpha_{i}k(\cdot,\mathbf{x}_{i})$ and $g(\cdot)=\sum_{j=1}^{n^{\prime}}\alpha^{\prime}_{j}k(\cdot,\mathbf{x}^{\prime}_{j})$ is given by
-$\langle f,g\rangle_{\mathcal{H}}=\sum_{i=1}^{n}\sum_{j=1}^{n^{\prime}}\alpha_{i}\alpha^{\prime}_{j}k(\mathbf{x}_{i},\mathbf{x}_{j}^{\prime}).$ (36)
-In GP regression, the optimization problem in eq. 33 becomes
-$R(f)=\sum_{i=1}^{n}(f(\mathbf{x}_{i})-y_{i})^{2}+\lambda\|f\|_{\mathcal{H}}^{2},$ (37)
-where $\lambda=\sigma_{\epsilon}^{2}$ is the noise varaince. Here the projection into one of the spaces $V_{\underline{l}}$ has the form
+To compute the optimal coefficients $c_{\underline{l}}$ in GP regression, we need to compute the inner product with respect to the reproducing kernel Hilbert space (RKHS) (Aronszajn, 1950).
+Let $\mathcal{H}$ be a RKHS endowed with an inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$, then the inner product between functions $f(\cdot)=\sum_{i=1}^{n}\alpha_{i}k(\cdot,\mathbf{x}_{i})$ and $g(\cdot)=\sum_{j=1}^{n^{\prime}}\alpha^{\prime}_{j}k(\cdot,\mathbf{x}^{\prime}_{j})$ is given by $\langle f,g\rangle_{\mathcal{H}}=\sum_{i=1}^{n}\sum_{j=1}^{n^{\prime}}\alpha_{i}\alpha^{\prime}_{j}k(\mathbf{x}_{i},\mathbf{x}_{j}^{\prime})$.
+In GP regression, the optimization problem in eq.
+16 becomes $R(f)=\sum_{i=1}^{n}(f(\mathbf{x}_{i})-y_{i})^{2}+\lambda\|f\|_{\mathcal{H}}^{2}$, where $\lambda=\sigma_{\epsilon}^{2}$ is the noise variance.
+Here the projection into one of the spaces $V_{\underline{l}}$ has the form
 $\mathcal{P}_{\underline{l}}f=\sigma_{\epsilon}^{-2}\mathbf{K}_{(\cdot)\mathbf{u}_{\underline{l}}}\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{l}}}^{-1}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{y}=\sigma_{\epsilon}^{-2}k(\cdot,\Omega_{\underline{l}})\bm{\alpha}_{\underline{l}}=\sigma_{\epsilon}^{-2}\sum_{i\in\mathbf{I}_{\underline{l}}}k(\cdot,\mathbf{z}_{i})\alpha_{i}:=f_{\underline{l}}(\cdot),$ (19)
@@
 where $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}=\big[\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}_{\underline{l}}\mathbf{f}}\mathbf{K}_{\mathbf{f}\mathbf{u}_{\underline{k}}}\big]$.
-The details are outlined in Algorithm 1 and Algorithm 2 in Appendix.
-Figure 1: Errors and time for posterior mean of zero mean and Matérn 3/2 GP with OptiCom and CT of dimension $d=2$ and level $\eta=1,2,\ldots,7$ tested on the Griewank function (Griewank,, 1981) over $n=2000$ training points and $n_{t}=1000$ test points.
-We denote GP with OptiCom by red dots, GP with CT by blue diamonds.
+The details are outlined in Algorithm 1 and Algorithm 2.
+Figure 1: Errors and time for posterior mean of zero mean and Matérn 3/2 GP with OptiCom and CT of dimension $d=2$ and level $\eta=1,2,\ldots,7$ tested on the Griewank function (Griewank, 1981) over $n=2000$ training points and $n_{t}=1000$ test points.
+We denote GP with OptiCom by circles and GP with CT by diamonds.
 Left: Logarithm of $l_{2}$ errors between GP posteriors and ground truth values averaged over 10 different lengthscales $[0.2,0.4,0.6,0.8,1.0,1.2,1.4,1.6,1.8,2.0]$.
-Right: Logarithm of time taken to compute the posterior mean averaged over $1000$ replications.
+Right: Logarithm of time taken to compute the posterior mean averaged over 10001000 replications.
 3.3.3 Complexity
-From Lemma 6 in (Garcke,, 2013), we know that the number of summed terms in combination technique formula eq.
+From Lemma 6 in (Garcke, 2013), we know that the number of summed terms in combination technique formula eq.
 11 is $b=\mathcal{O}(\eta^{d-1})$, the size of each summed term is of order $m_{b}=\mathcal{O}(2^{(\eta+d-1)})$, and dimension of the sparse grid space $V_{\eta}^{s}$ defined in eq.
-28, i.e., the number of grid points, is $|V_{\eta}^{s}|=\mathcal{O}(2^{\eta}\cdot\eta^{d-1})=\mathcal{O}\big{(}h_{\eta}^{-1}\cdot\log(h_{\eta}^{-1})^{d-1}\big{)}$, where $h_{\eta}=2^{-\eta}$ is the mesh size of the sparse grids $\Omega_{\eta}^{d}$.
+11, i.e., the number of grid points, is $|V_{\eta}^{s}|=\mathcal{O}(2^{\eta}\cdot\eta^{d-1})=\mathcal{O}\big(h_{\eta}^{-1}\cdot\log(h_{\eta}^{-1})^{d-1}\big)$, where $h_{\eta}=2^{-\eta}$ is the mesh size of the sparse grids $\Omega_{\eta}^{d}$.
 The additional operations compared to GP with CT include computing $\mathbf{M}_{\mathbf{u}_{\underline{l}}\mathbf{u}_{\underline{k}}}$ when $\underline{l}\neq\underline{k}$ and solving a $b\times b$ linear system to obtain the optimal coefficients, which costs $\mathcal{O}(b^{2}m_{b}^{2}n)$ and $\mathcal{O}(b^{3})$ time respectively.
@@
 4 Distributed GP with optimal weights
-In this section, we extend the OptiCom to distributed GPs.
+In our previous formulation of GP with OptiCom, the inducing points were placed on fixed sparse grids rather than being learned through variational inference.
+While this design simplifies computation, it can limit performance when the sparse grid resolution is low, as the fixed locations may not adequately capture complex input structures.
+To address this limitation, we extend OptiCom to distributed GPs, which allows for greater flexibility and improved accuracy while maintaining computational efficiency and does not rely on sparse grid structures.
 Following the assumptions Section 2.2.1, we consider the global training dataset $\mathcal{D}=\{\mathbf{X},\mathbf{y}\}$ of size $n$, partitioned into $M$ local datasets $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$ of size $n_{i}$ for $i=1,\ldots,M$.
@@
 $J(\beta_{1},\ldots,\beta_{M})=\|f-\sum_{i=1}^{M}\beta_{i}\mathcal{P}_{i}f\|,$ (21)
-where $f(\cdot)\sim\mathcal{GP}(0,k(\cdot,\cdot^{\prime}))$ as defined in Section 2.2.1, $\mathcal{P}_{i}f=f|\mathbf{y}_{i}:=f_{i}$ is the posterior conditioned on the local dataset $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$, and its distribution takes the form of Equation 12.
+where $f(\cdot)\sim\mathcal{GP}(0,k(\cdot,\cdot))$ as defined in Section 2.2.1, $\mathcal{P}_{i}f=f|\mathbf{y}_{i}:=f_{i}$ is the posterior conditioned on the local dataset $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$, and its distribution takes the form of eq.
+2.
 Similar to Section 3.3, the optimal weights $\{\beta_{i}\}_{i=1}^{M}$ can be obtained by solving the following linear system:
-$\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}&\cdots&\langle\mathcal{P}_{1}f,\mathcal{P}_{M}f\rangle\\ \langle\mathcal{P}_{2}f,\mathcal{P}_{1}f\rangle&\cdots&\langle\mathcal{P}_{2}f,\mathcal{P}_{M}f\rangle\\ \vdots&\ddots&\vdots\\ \langle\mathcal{P}_{M}f,\mathcal{P}_{1}f\rangle&\cdots&\|\mathcal{P}_{M}f\|^{2}\end{bmatrix}\begin{bmatrix}\beta_{1}\\ \beta_{2}\\ \vdots\\ \beta_{M}\end{bmatrix}=\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}\\ \|\mathcal{P}_{2}f\|^{2}\\ \vdots\\ \|\mathcal{P}_{M}f\|^{2}\end{bmatrix}.$ (41)
+$\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}&\cdots&\langle\mathcal{P}_{1}f,\mathcal{P}_{M}f\rangle\\ \langle\mathcal{P}_{2}f,\mathcal{P}_{1}f\rangle&\cdots&\langle\mathcal{P}_{2}f,\mathcal{P}_{M}f\rangle\\ \vdots&\ddots&\vdots\\ \langle\mathcal{P}_{M}f,\mathcal{P}_{1}f\rangle&\cdots&\|\mathcal{P}_{M}f\|^{2}\end{bmatrix}\begin{bmatrix}\beta_{1}\\ \beta_{2}\\ \vdots\\ \beta_{M}\end{bmatrix}=\begin{bmatrix}\|\mathcal{P}_{1}f\|^{2}\\ \|\mathcal{P}_{2}f\|^{2}\\ \vdots\\ \|\mathcal{P}_{M}f\|^{2}\end{bmatrix}$.
+(22)
 To reduce computational complexity, we randomly select a central dataset $\mathcal{D}_{c}:=\{\mathbf{X}_{c},\mathbf{y}_{c}\}\subset\mathcal{D}=\{\mathbf{X},\mathbf{y}\}$ of size $n_{c}\ll n$.
@@
 where $\bm{\alpha}_{l}=[\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(l)}]^{-1}\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}\mathbf{y}_{l}$, $\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(l)}=\big[\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}\mathbf{K}_{\mathbf{f}_{l}\mathbf{u}}\big]$, $\mathbf{K}_{\mathbf{u}\mathbf{u}}=k(\mathbf{Z},\mathbf{Z})$, $\mathbf{K}_{\mathbf{u}\mathbf{f}_{l}}=k(\mathbf{Z},\mathbf{X}_{l})=\mathbf{K}_{\mathbf{f}_{l}\mathbf{u}}^{\top}$, $\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(c)}=\big[\mathbf{K}_{\mathbf{u}\mathbf{u}}+\sigma_{\epsilon}^{-2}\mathbf{K}_{\mathbf{u}\mathbf{f}_{c}}\mathbf{K}_{\mathbf{f}_{c}\mathbf{u}}\big]$, $\mathbf{K}_{\mathbf{u}\mathbf{f}_{c}}=k(\mathbf{Z},\mathbf{X}_{c})=\mathbf{K}_{\mathbf{f}_{c}\mathbf{u}}^{\top}$, $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{m}$ are the inducing inputs corresponding to the optimized inducing variables $\mathbf{u}\in\mathbb{R}^{m}$.
-The details are outlined in Algorithm 3 in Appendix.
+The details are outlined in Algorithm 3.
+In addition, the optimal weights computed in eq.
+22 inherently capture the correlations between local experts.
+This mechanism prevents redundant information from being overemphasized when experts are highly correlated, thus leading to a more balanced aggregation.
+As a result, the aggregated prediction becomes more consistent with that of a full GP.
+This integration of expert correlations is key to enhancing both the performance and stability of distributed GP.
 4.1.2 Aggregated prediction
@@
 $\sigma_{\mathcal{A}}^{2}(\cdot)=\sum_{i=1}^{M}\beta_{i}^{2}\mathrm{Var}[f_{i}(\cdot),f_{i}(\cdot)]=\sum_{i=1}^{M}\beta_{i}^{2}\mathbf{K}_{(\cdot)\mathbf{u}}[\mathbf{M}_{\mathbf{u}\mathbf{u}}^{(i)}]^{-1}\mathbf{K}_{\mathbf{u}(\cdot)},$ (24b)
-The details are outlined in Algorithm 4 in Appendix.
+The details are outlined in Algorithm 4.
 4.1.3 Complexity
@@
 The comparison of the time complexity of the aggregation models and full SVGP is presented in the first row of Table 1.
-Table 1: Complexity of the prediction for aggregation models and full GP.
+Table 1: Complexity of the prediction for aggregation models.
 $m$ is the number of inducing points, $n$ is the size of the entire training dataset, $n_{i}$ is the size of the local dataset $\mathcal{D}_{i}$ corresponding to the expert $\mathcal{M}_{i}$, $i=1,\ldots.\,M$, $n_{c}$ is the size of the dataset $\mathcal{D}_{c}$ corresponding to the expert $\mathcal{M}_{c}$ for grBCM.
 $M$ is the number of experts.
-Full PoE gPoE BCM rBCM grBCM NPAE opt (ours)
-Time SVGP $\mathcal{O}(m^{2}n)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}(n_{i}+n_{c})+M)$ $\mathcal{O}(m^{2}n_{i}M+M^{3})$ $\mathcal{O}(m^{2}n_{i}+M^{3})$
-ExactGP $\mathcal{O}(n^{3})$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}((n_{i}+n_{c})^{3}+M)$ $\mathcal{O}(n_{i}^{3}M+M^{3})$ $\mathcal{O}(n_{i}^{3}+M^{3})$
+PoE gPoE BCM rBCM grBCM NPAE opt (ours)
+Time SVGP $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}n_{i}+M)$ $\mathcal{O}(m^{2}(n_{i}+n_{c})+M)$ $\mathcal{O}(m^{2}n_{i}M+M^{3})$ $\mathcal{O}(m^{2}n_{i}+M^{3})$
+ExactGP $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}(n_{i}^{3}+M)$ $\mathcal{O}((n_{i}+n_{c})^{3}+M)$ $\mathcal{O}(n_{i}^{3}M+M^{3})$ $\mathcal{O}(n_{i}^{3}+M^{3})$
 4.2 Distributed exact GP
@@
 In this case, the distribution of $\mathcal{P}_{i}f=f|\mathbf{y}_{i}:=f_{i}$ follows the form given in eq.
-4.
+1.
 Consequently, the scalar product is expressed as:
@@
 Therefore, the joint mean and joint variance are given by:
-$\mu_{\mathcal{A}}(\cdot)=\sum_{i=1}^{M}\beta_{i}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{y}_{i},$ (45a)
-$\sigma_{\mathcal{A}}^{2}(\cdot)=\sum_{i=1}^{M}\beta_{i}^{2}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}(\cdot)},$ (45b)
-The detailed procedures are outlined in Algorithm 5 and Algorithm 6 in Appendix.
+$\mu_{\mathcal{A}}(\cdot)=\sum_{i=1}^{M}\beta_{i}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{y}_{i},\quad\sigma_{\mathcal{A}}^{2}(\cdot)=\sum_{i=1}^{M}\beta_{i}^{2}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}(\cdot)}.$ (26)
+The detailed procedures are outlined in Algorithm 5 and Algorithm 6.
 The time complexity of the parallelized Algorithm 6 is $\mathcal{O}(n_{i}^{3}+M^{3})$, where $n_{i}$ is the size of the local dataset $\mathcal{D}_{i}$ for expert $\mathcal{M}_{i}$, and $M$ is the number of experts.
 A comparison of the time complexity of the aggregation models and the full exact GP is presented in the last row of Table 1.
-5 Experiment
-In this section, we evaluate the proposed algorithm against several baseline methods: full GP, PoE, gPoE, BCM, rBCM, and grBCM, using both synthetic data (Section 5.1) and UCI datasets (Asuncion et al.,, 2007) (Section 5.2).
+4.3 Convergence
+In this section, we introduce the theoretical properties of the proposed aggregation method with optimal weights.
+In contrast to PoE and BCM families whose weights are chosen heuristically, our weights are defined as the minimizer of a function space residual objective in eq.
+21 and are computed from the resulting linear system in eq.
+22.
+This optimization-based method is what enables the consistency for the aggregated predictor.
+Setting.
+Let the global dataset $\mathcal{D}$ be randomly partitioned into local datasets $\{\mathcal{D}_{i}\}_{i=1}^{M_{n}}$ without replacement, where the number of experts $M_{n}$ increases with the global dataset size $n$ such that (i) $\lim_{n\to\infty}M_{n}=\infty$, and (ii) $\liminf_{n\to\infty}\frac{n}{M_{n}^{2}}>0$.
+Then the local sample sizes satisfy $n_{i}\asymp n/M_{n}\to\infty$.
+Assume each expert $\mathcal{M}_{i}$ is a zero-mean GP with the same kernel and Gaussian noise model as the full GP.
+For each GP expert, let $\mathcal{P}_{i}f$ denote the posterior operator conditioned on the local dataset $\mathcal{D}_{i}$ (as defined in eq.
+21).
+Let $\mu_{i}(\cdot)$ and $\sigma_{i}^{2}(\cdot)$ be the predictive mean and variance of expert $i$, where $i=1,\ldots,M_{n}$, and let $\mu_{\text{full}}(\cdot)$ and $\sigma_{\text{full}}^{2}(\cdot)$ denote the predictive mean and variance of the full GP trained on $\mathcal{D}$.
+Assumptions.
+  1. 1.
+    local predictive consistency.
+    We assume a standard GP predictive consistency property: as the local sample size $n_{i}\to\infty$, the predictive mean and variance of the $i$-th expert converge to the same limiting predictive quantities as the full GP, at any fixed test input $\mathbf{x}_{*}$.
+This is a classical result in GP posterior consistency theory (Choi, 2005), and it is also the auxiliary fact from proof of Proposition 2 in (Liu et al., 2018) to control the asymptotic behavior of each expert’s predictive mean and variance.
+Specifically, for any fixed $\mathbf{x}_{*}$,
+    $\lim_{n\to\infty}\max_{1\leq i\leq M_{n}}\big|\mu_{i}(\mathbf{x}_{*})-\mu_{\text{full}}(\mathbf{x}_{*})\big|=0,\qquad\lim_{n\to\infty}\max_{1\leq i\leq M_{n}}\big|\sigma_{i}^{2}(\mathbf{x}_{*})-\sigma_{\text{full}}^{2}(\mathbf{x}_{*})\big|=0.$ (27)
+  2. 2.
+    weight stability.
+    We assume the optimizer $\bm{\beta}_{n}=(\beta_{1,n},\ldots,\beta_{M_{n},n})^{\top}$ returned by the linear system in eq. 22 is stable, i.e.:
+    $\|\bm{\beta}_{n}\|_{\ell_{1}}\leq B\quad\text{for all sufficiently large $n$},$ (28)
+    for some constant $B$ independent of $n$.
+This assumption is standard for linear estimators: without controlling the magnitude of $\bm{\beta}_{n}$, the residual term $\sum_{i=1}^{M_{n}}\beta_{i,n}(f-\mathcal{P}_{i}f)$ below may fail to vanish even if each $(f-\mathcal{P}_{i}f)\to 0$.
+In practice, such stability can be enforced by adding a small jitter regularizer to the Gram matrix in eq.
+22.
+Aggregation prediction.
+The aggregated predictor of the proposed method has the form (see eqs. 24 and 26)
+$\mu_{\mathcal{A}}(\mathbf{x}_{*})=\sum_{i=1}^{M_{n}}\beta_{i,n}\,\mu_{i}(\mathbf{x}_{*}),\qquad\sigma_{\mathcal{A}}^{2}(\mathbf{x}_{*})=\sum_{i=1}^{M_{n}}\beta_{i,n}^{2}\,\sigma_{i}^{2}(\mathbf{x}_{*}).$ (29)
+Define the sum of weights $s_{n}:=\sum_{i=1}^{M_{n}}\beta_{i,n}$, then we have the decomposition as follows:
+$\mu_{\mathcal{A}}(\mathbf{x}_{*})=s_{n}\,\mu_{\text{full}}(\mathbf{x}_{*})+\sum_{i=1}^{M_{n}}\beta_{i,n}\big(\mu_{i}(\mathbf{x}_{*})-\mu_{\text{full}}(\mathbf{x}_{*})\big).$ (30)
+Thus, mean consistency reduces to proving (i) $s_{n}\to 1$, and (ii) the second term in eq. 30 converges to zero.
+Mean.
+  • Step 1: the objective value at the optimizer converges to zero.
+    Define the function space residual objective from eq. 21
+    $J_{n}(\bm{\beta})\;:=\;\Big\|f-\sum_{i=1}^{M_{n}}\beta_{i}\,\mathcal{P}_{i}f\Big\|.$ (31)
+    Let $\bm{\beta}_{n}$ be a minimizer of $J_{n}(\bm{\beta})$. Then for any $j\in\{1,\dots,M_{n}\}$,
+    $J_{n}(\bm{\beta}_{n})\;\leq\;J_{n}(\mathbf{e}_{j})\;=\;\|f-\mathcal{P}_{j}f\|,$ (32)
+    where $\mathbf{e}_{j}$ is the $j$-th standard basis vector. By Assumption 1 and $n_{j}\asymp n/M_{n}\to\infty$, we have
+    $\lim_{n\to\infty}\|f-\mathcal{P}_{j}f\|=0$ (33)
+    Therefore,
+    $\lim_{n\to\infty}J_{n}(\bm{\beta}_{n})=0$ (34)
+  • Step 2: derive $s_{n}=\sum_{i=1}^{M_{n}}\beta_{i,n}\to 1$.
+    Using the identity
+    $f-\sum_{i=1}^{M_{n}}\beta_{i,n}\mathcal{P}_{i}f=(1-s_{n})f+\sum_{i=1}^{M_{n}}\beta_{i,n}\big(f-\mathcal{P}_{i}f\big),$ (35)
+    we bound $|1-s_{n}|$ by applying the triangle inequality:
+    $|1-s_{n}|\cdot\|f\|$ $=\big\|(1-s_{n})f\big\|$
+    $\leq\Big\|f-\sum_{i=1}^{M_{n}}\beta_{i,n}\mathcal{P}_{i}f\Big\|+\Big\|\sum_{i=1}^{M_{n}}\beta_{i,n}(f-\mathcal{P}_{i}f)\Big\|$
+    $\leq J_{n}(\bm{\beta}_{n})+\sum_{i=1}^{M_{n}}|\beta_{i,n}|\cdot\|f-\mathcal{P}_{i}f\|$
+    $\leq J_{n}(\bm{\beta}_{n})+\|\bm{\beta}_{n}\|_{\ell_{1}}\,\max_{1\leq i\leq M_{n}}\|f-\mathcal{P}_{i}f\|.$ (36)
+    By eq.
+34, the first term $J_{n}(\bm{\beta}_{n})$ tends to 0.
+By Assumption 1, $\lim_{n\to\infty}\max_{i}\|f-\mathcal{P}_{i}f\|=0$, and by Assumption 2, $\|\bm{\beta}_{n}\|_{\ell_{1}}\leq B$.
+Hence the right hand side of section 4.3 converges to 0, which implies that
+    $\lim_{n\to\infty}\sum_{i=1}^{M_{n}}\beta_{i,n}=\lim_{n\to\infty}s_{n}=1.$ (37)
+  • Step 3: prove mean consistency. Return to eq. 30. The second term is bounded by
+    $\Big|\sum_{i=1}^{M_{n}}\beta_{i,n}\big(\mu_{i}(\mathbf{x}_{*})-\mu_{\text{full}}(\mathbf{x}_{*})\big)\Big|\leq\|\bm{\beta}_{n}\|_{\ell_{1}}\,\max_{1\leq i\leq M_{n}}\big|\mu_{i}(\mathbf{x}_{*})-\mu_{\text{full}}(\mathbf{x}_{*})\big|\xrightarrow[]{}0\quad\text{as}\quad n\to\infty$ (38)
+    using eqs. 28 and 27. Combining this with eq. 37 yields
+    $\lim_{n\to\infty}\mu_{\mathcal{A}}(\mathbf{x}_{*})$ $=\lim_{n\to\infty}\Big(s_{n}\,\mu_{\text{full}}(\mathbf{x}_{*})+\sum_{i=1}^{M_{n}}\beta_{i,n}\big(\mu_{i}(\mathbf{x}_{*})-\mu_{\text{full}}(\mathbf{x}_{*})\big)\Big)$ (39)
+    $=1\cdot\mu_{\text{full}}(\mathbf{x}_{*})+0$ (40)
+    $=\mu_{\text{full}}(\mathbf{x}_{*}).$ (41)
+    Therefore, the aggregated predictive mean is consistent.
+Variance.
+Using eq. 29 we get the following decomposition for aggregated predictive variance:
+$\sigma_{\mathcal{A}}^{2}(\mathbf{x}_{*})$ $=\sum_{i=1}^{M_{n}}\beta_{i,n}^{2}\sigma_{i}^{2}(\mathbf{x}_{*})$
+$=\Big(\sum_{i=1}^{M_{n}}\beta_{i,n}^{2}\Big)\sigma_{\text{full}}^{2}(\mathbf{x}_{*})+\sum_{i=1}^{M_{n}}\beta_{i,n}^{2}\big(\sigma_{i}^{2}(\mathbf{x}_{*})-\sigma_{\text{full}}^{2}(\mathbf{x}_{*})\big).$ (42)
+If $\|\bm{\beta}_{n}\|_{\ell_{2}}$ is uniformly bounded, then the second term in section 4.3 goes to zero by Assumption 1.
+However, to obtain $\sigma_{\mathcal{A}}^{2}(\mathbf{x}_{*})\to\sigma_{\text{full}}^{2}(\mathbf{x}_{*})$ one further needs the stronger condition $\sum_{i}\beta_{i,n}^{2}\to 1$, which we do not enforce.
+Consequently, our theoretical guarantee is primarily for the predictive mean, while variance consistency would require additional constraints on $\bm{\beta}_{n}$.
+Stability.
+Assumption 2 is required for a rigorous argument, i.e., the term $\sum_{i}\beta_{i,n}(f-\mathcal{P}_{i}f)$ in eq.
+35 may not vanish without controlling the magnitude of $\bm{\beta}_{n}$ even when each component $(f-\mathcal{P}_{i}f)\to 0$.
+In our method, $\bm{\beta}_{n}$ is obtained by solving the finite dimensional linear system eq.
+22 derived from the first-order optimality conditions of the function space objective $J_{n}(\bm{\beta})$.
+The solution is stable whenever the corresponding Gram matrix is well-conditioned.
+In practice, we ensure numerical stability by adding a small jitter to eq.
+22, which is standard in GP computations and typically yields bounded solutions.
+5 Experiments
+In this section, we evaluate the proposed algorithm against full GP and several baselines that aggregate predictions via weighted averaging: PoE, gPoE, BCM, rBCM, and grBCM, using both synthetic data (Section 5.1) and UCI datasets (Asuncion and Newman, 2007) (Section 5.2).
+We restrict baselines to aggregation methods because our contribution is an aggregation algorithm for distributed GP experts, scalable methods for centralized GP such as inducing points approximations, structured kernel interpolation, Vecchia framework, or stochastic dual descent solvers are complementary and can be used within each expert, but they do not provide a competing aggregation mechanism under the distributed learning setting.
 The size of the inducing variables is set to $m=128$.
-We use the radial basis function (RBF) kernel with separate lengthscales for each input dimension, defined by the following equation:
-$k(\mathbf{x},\mathbf{x}^{\prime})=\sigma_{f}^{2}\exp\left(-\sum_{i=1}^{d}\frac{(\mathbf{x}_{i}-\mathbf{x}^{\prime}_{i})^{2}}{2l_{i}^{2}}\right),$ (46)
-where $\mathbf{x}_{i},\mathbf{x}^{\prime}_{i}\in\mathbb{R}$ are the $i$-th components of the inputs $\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}$, $l_{i}$ is the lengthscale along the $i$-th dimension, and $\sigma_{f}^{2}>0$ is the signal variance.
+We use the radial basis function (RBF) kernel with separate lengthscales $l_{i}$ for each input dimension $i$, and $\sigma_{f}^{2}>0$ is the signal variance of the kernel.
+The code for the experiments is available at https://github.com/hchen19/optimal-distgp.
 5.1 Synthetic data
@@
 We generate a training dataset of size $n=10^{4}$ and dimension $d=2$ using the RBF kernel with $(l_{1},l_{2},\sigma_{\epsilon},\sigma_{f})=(0.3,1.3,0.1,0.2)$.
-The dataset is distributed among $M=2,4,8,19,20,40,80,100$ experts, respectively.
-We set the initial values of the hyperparameters $(l_{1}^{(0)},l_{2}^{(0)},\sigma_{\epsilon}^{(0)},\sigma_{f}^{(0)})=(2.0,0.3,1.0,\sqrt{1.5})$, and use Adam optimizer (Kingma and Ba,, 2014) with a learning rate of $0.1$ for all methods.
+The dataset is distributed among $M=2,4,8,10,20,40,80,100$ experts, respectively.
+We set the initial values of the hyperparameters $(l_{1}^{(0)},l_{2}^{(0)},\sigma_{\epsilon}^{(0)},\sigma_{f}^{(0)})=(2.0,0.3,1.0,\sqrt{1.5})$, and use Adam optimizer (Kingma and Ba, 2015) with a learning rate of $0.1$ for all methods.
 For ADMM, the regularization parameter is set to $\rho=500$.
@@
 We observe that FACT and ADMM perform worse as the number of experts $M$ increases, while FedAvg performs the best overall, converging faster and more stably than FedProx.
-Figure 8 in Appendix shows boxplots of the hyperparameter estimates after 200 iterations, indicating that FedAvg converges closest to the ground truth with the least variability.
-Therefore, we use FedAvg for distributed training in the subsequent experiments with the UCI datasets.
+Figure 8 shows boxplots of the hyperparameter estimates after 200 iterations, indicating that FedAvg converges closest to the ground truth with the least variability.
+Therefore, we use FedAvg for distributed training in the experiments with the UCI datasets.
+Figure 3: SVGP posteriors with the different aggregation models on $n=100$ training points in [1,1][-1,1] and $n_{t}=50$ test points in [1.2,1.2][-1.2,1.2] with the number of experts $M=4$, dimension $d=1$, and the number of inducing variables $m=128$.
+We denote the aggregated predictions (aggreg) by dashed lines, the full SVGP posteriors (full) by solid lines, and the observations by dots.
 5.1.2 Prediction
 In this section, we evaluate the performance of aggregation models and full GP models without optimizing any hyperparameters.
-We use the Ackley function (Ackley,, 2012) defined over the interval $[-1,1]^{d}$, to benchmark the aggregation models:
-$f(\mathbf{x})=-a\exp\left(-b\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}\right)-\exp\left(\frac{1}{d}\sum_{i=1}^{d}\cos(cx_{i})\right)+a+\exp(1),\quad\mathbf{x}\in\mathbb{R}^{d},$ (47)
-with parameters $a=20$, $b=0.2$, $c=2\pi$.
-We employ the RBF kernel with fixed hyperparameters $(l_{i},\sigma_{\epsilon}^{2},\sigma_{f}^{2})=(0.3,0.025,1.0)$ for all $i=1,\ldots,d$.
+We generate the ground truth dataset using RBF kernel with fixed hyperparameters $(l_{i},\sigma_{\epsilon}^{2},\sigma_{f}^{2})=(0.3,0.25,1.0)$ for all $i=1,\ldots,d$.
 For SVGP, we use $m=128$ inducing points, randomly generated with a fixed seed across all aggregation methods.
-Figure 3: SVGP posteriors with the different aggregation models on the Ackley function on $n=10^{4}$ training points and $n_{t}=50$ test points with the number of experts $M=4$, dimension $d=1$, and the the number of inducing variables $m=128$.
-We denote the aggregated predictions (aggreg), by red lines, the full SVGP posteriors (full) by blue lines, and the ground truth values (true) by black dots.
-Figure 3 shows the aggregated predictions of SVGP on $n=10^{4}$ training points and $n_{t}=50$ test points, with dimension $d=1$ and $M=4$ experts.
-The proposed method (opt) and five baselines (PoE, gPoE, BCM, rBCM, grBCM) are compared to the full SVGP.
-All methods use the same inducing points.
-It’s observed that PoE, gPoE, grBCM and the proposed method (opt) provide better predictive means than BCM and rBCM.
-However, the predictive variances of PoE, BCM, and rBCM shrink to zero and is not consistent with the full SVGP.
-Our method, the distributed SVGP with optimal weights, delivers accurate predictions for both the mean and the variance.
+Figure 3 shows the aggregated predictions of SVGP on $n=100$ training points uniformly sampled in [1,1][-1,1] and $n_{t}=50$ test points evenly spaced in [1.2,1.2][-1.2,1.2], with dimension $d=1$ and $M=4$ experts.
+We compare our method (opt) against five baselines (PoE, gPoE, BCM, rBCM, grBCM) and the full SVGP model using the same inducing points.
+The proposed method yields predictions that closely match the full SVGP, especially in the extrapolation region outside the training domain (i.e., in $1.2,-1]\cup[1,1.2]$).
+In contrast, most baselines show degraded performance in these regions.
+Furthermore, the predictive variances of our method and grBCM align more closely with the full SVGP, suggesting better consistency of the predictions compared to other methods.
 (a) mean
 (b) standard deviation
-Figure 4: Comparison of the aggregation models for Exact GP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$.
+Figure 4: Comparison of the aggregation models for Exact GP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points in $1,1]^{2}$ and $n_{t}=2500$ test points in $1.2,1.2]^{2}$.
 The number of experts considered are $M=2,4,8,10,20,40,80,100$.
@@
 Left: Logarithm of time for computing the aggregated predictions.
-Middle: RMSE between the aggregated predictions and the ground truth.
-Right: NLPD between the aggregated predictions and the ground truth.
-Top: Mean of the computational time, RMSE and NLPD.
-Bottom: Standard deviation of the computational time, RMSE and NLPD.
+Middle Left: RMSE between the aggregated predictions and the ground truth.
+Middle Right: NLPD between the aggregated predictions and the ground truth.
+Right: 2-Wasserstein distance to the full GP.
+Top: Mean of the computational time, RMSE, NLPD, and 2-Wasserstein distance.
+Bottom: Standard deviation of the computational time, RMSE, NLPD, and 2-Wasserstein distance.
 (a) mean
 (b) standard deviation
-Figure 5: Comparison of the aggregation models for SVGP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, $m=128$ inducing points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$.
+Figure 5: Comparison of the aggregation models for SVGP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points in $1,1]^{2}$, $m=128$ inducing points, and $n_{t}=2500$ test points in $1.2,1.2]^{2}$.
 The number of experts considered are $M=2,4,8,10,20,40,80,100$.
@@
 Left: Logarithm of time for computing the aggregated predictions.
-Middle: RMSE between the aggregated predictions and the ground truth.
-Right: NLPD between the aggregated predictions and the ground truth.
-Top: Mean of the computational time, RMSE and NLPD.
-Bottom: Standard deviation of the computational time, RMSE and NLPD.
-Next, we evaluate the aggregation models on $n=10^{4}$ training data of dimension $d=2$ and $n_{t}=50\times 50=2500$ test points, evenly spaced over the interval $[-1,1]^{2}$.
+Middle Left: RMSE between the aggregated predictions and the ground truth.
+Middle Right: NLPD between the aggregated predictions and the ground truth.
+Right: 2-Wasserstein distance to the full GP.
+Top: Mean of the computational time, RMSE, NLPD, and 2-Wasserstein distance.
+Bottom: Standard deviation of the computational time, RMSE, NLPD, and 2-Wasserstein distance.
+Next, we evaluate the aggregation models on $n=10^{4}$ training data of dimension $d=2$ in $1,1]^{2}$ and $n_{t}=50\times 50=2500$ test points, evenly spaced over the interval $1.2,1.2]^{2}$.
 The training dataset is divided among $M=2,4,8,10,20,40,80,100$ experts, respectively.
-To eliminate the impact of lengthscales and seeds, we average all the metrics over 10 different lengthscales $[0.2,0.4,0.6,0.8,1.0,1.2,1.4,1.6,1.8,2.0]$ and 10 different seeds $1,2,3,4,5,6,7,8,9,10$ for each setting of $M$.
-Figure 4 compares three metrics for exact GPs: computational time, root mean squared error (RMSE), and negative log predictive density (NLPD) of all aggregation models as a function of the number of training points per GP expert.
-While grBCM preforms the best on the mean of RMSE and NLPD, it requires more computational time due to the larger size of the augmented local training datasets and has larger standard deviations of the metrics.
-Our method, GP with optimal weights, provides comparable means for these metrics with smaller standard deviations, indicating greater stability and less sensitivity compared to grBCM.
-The boxplots of RMSE and NLPD for exact GPs are presented in Figure 9 and Figure 10 in Appendix.
-These boxplots highlight the stability of our method, especially as the number of experts $M$ increases.
-Figure 5 compares three metrics as a function of the number of training points per GP expert for SVGP across all aggregation models.
-For SVGP, PoE, gPoE, grBCM, significantly outperform the full SVGP, BCM, and rBCM.
-The boxplots of RMSE and NLPD for SVGP are shown in Figure 11 and Figure 12 in Appendix.
-These boxplots demonstrate that PoE, gPoE, grBCM, and our method all achieve the smallest means and variabilities for these metrics.
-5.2 UCI regression
-Figure 6: Comparison of the aggregation models for SVGP on the UCI datasets.
-In this section, we evaluate the aggregation models on five real datasets from the UCI repository (Asuncion et al.,, 2007): PoleTele, Elevators, Kin40k, KEGG, and Protein.
-Each dataset is split into training and test sets in an 80:20 ratio.
-For the smaller datasets (PoleTele and Elevators, each around $15,000$ samples), the training datasets are divided into $M=4,10,20$ experts.
-For the larger datasets (Kin40k, KEGG, and Protein, each around $45,000$ samples), the training datasets are divided into $M=8,16,64$ experts.
-We use SVGP with $m=128$ inducing points and RBF kernel defined in eq.
-46.
-All aggregation models utilize FedAvg with the Adam optimizer, using the same initial hyperparameters over 100 iterations for distributed training.
-Results are averaged over 10 different seeds.
-Figure 6 presents the performance metrics of the aggregation models on these five datasets.
-BCM performs the worst, particularly as the number of experts $M$ increases.
-While grBCM performs well for RMSE on most datasets, it has significantly worse NLPD on PoleTele.
-Although our method, distributed SVGP with optimal weights, does not always perform the best, it consistently provides stable and comparable predictions in less time than grBCM.
-Detailed metric values are reported in Table 2 and Table 3 in Appendix.
-6 Conclusion
+To eliminate the impact of seeds, we average all the metrics over 10 different seeds for each setting of $M$.
+Figure 4 compares the performance of various aggregation models for exact GPs using four metrics: computational time, root mean squared error (RMSE), negative log predictive density (NLPD), and 2-Wasserstein distance to the full exact GP (measuring how closely each aggregated model approximates the full GP).
+While grBCM performs the best in RMSE, NLPD, and 2-Wasserstein distance, it requires more computational time due to the larger size of the augmented local training datasets and has larger standard deviations of the metrics.
+Our method, GP with optimal weights, provides comparable predictions with significantly lower computational cost.
+Figure 5 also compares four metrics for SVGP across all aggregation models.
+It’s observed that our method and grBCM significantly outperform other methods for SVGP.
+5.2 UCI regression for temporal extrapolation
+In this section, we evaluate the aggregation models on three real-world time-series datasets from the UCI repository (Asuncion and Newman, 2007): Gas Turbine CO and NOx Emission (Kaya et al., 2019), Beijing PM2.5 (Liang et al., 2015), and Metro Interstate Traffic Volume (Hogue, 2019).
+Unlike standard random train-test splits, we employ temporal extrapolation splits where models are trained on earlier years and tested on future years, creating a more realistic and challenging evaluation scenario that tests the models’ ability to handle distribution shift and regime changes.
+Dataset descriptions and splits:
+  • Gas Turbine CO & NOx (36,733 samples, 10 features): Hourly sensor measurements from 2011–2015.
+Training on 2011–2013, testing on 2014–2015.
+This dataset contains two prediction targets: CO and NOx emissions, we run experiments independently for each target.
+The future years contain operational regimes not fully represented in training data.
+  • Beijing PM2.5 (43,824 samples, 13 features): Hourly air quality measurements from 2010–2014.
+Training on 2010–2013, testing on 2014.
+The test period includes extreme pollution events such as haze episodes that challenge model calibration.
+  • Metro Traffic (48,204 samples, 8 features): Hourly traffic volume measurements from 2012–2018.
+Training on 2012–2016, testing on 2017–2018.
+Traffic patterns exhibit regime shifts from construction, policy changes, and seasonal variations.
+Experimental setup:
+To create a challenging distributed setting where expert diversity and correlation become important, we partition each training dataset into $M=20,40,60$ heterogeneous experts by randomly shuffling and dividing the data.
+Specifically, we use non-shared hyperparameters across experts (i.e., each expert optimizes its own lengthscale, kernel variance, and noise variance independently), which creates the strong correlated local datasets that our method can handle.
+We employ SVGP with $m=128$ inducing points and RBF kernel.
+All aggregation models utilize FedAvg with the Adam optimizer, with 500 training iterations to ensure sufficient convergence for the extrapolation task.
+Results are averaged over 10 different random seeds.
+Figure 6 presents the performance metrics of the aggregation models on these four datasets.
+It’s observed that: (1) BCM and rBCM perform poorly, particularly for NOx and Beijing PM2.5, where it yields significantly degraded RMSE and NLPD due to overconfident but incorrect experts being equally weighted.
+(2) While grBCM shows competitive RMSE, it requires substantially more computation time due to the need to train additional models.
+(3) Our method, distributed SVGP with optimal weights (opt), consistently achieves the best or near-best RMSE and NLPD across all datasets while maintaining computational efficiency comparable to simpler baselines.
+Notably, on the challenging extrapolation tasks (NOx and Metro Traffic), our method’s ability to account for expert correlation and reliability provides clear advantages in both prediction accuracy and uncertainty quantification.
+Detailed metric values are reported in Table 2.
+Figure 6: Comparison of aggregation models for SVGP on temporal extrapolation tasks. Models are trained on earlier years and tested on future years to evaluate performance under distribution shift.
+Table 2: Results of aggregation models for SVGP on UCI datasets for temporal extrapolation tasks.
+Models are trained on earlier years and tested on future years.
+Best and worst performances are highlighted in yellow and red, respectively.
+Data $M$ Metric SVGP PoE gPoE BCM rBCM grBCM opt (ours)
+Gas Turbine CO 20 Time 35.405 $\pm$ 0.415 0.164 $\pm$ 0.002 0.164 $\pm$ 0.001 0.163 $\pm$ 0.001 0.164 $\pm$ 0.001 0.297 $\pm$ 0.001 0.163 $\pm$ 0.001
+RMSE 1.642 $\pm$ 0.053 1.518 $\pm$ 0.005 1.518 $\pm$ 0.005 2.828 $\pm$ 0.054 2.510 $\pm$ 0.045 1.615 $\pm$ 0.117 1.519 $\pm$ 0.005
+NLPD 1.955 $\pm$ 0.047 2.310 $\pm$ 0.009 2.157 $\pm$ 0.007 5.881 $\pm$ 0.195 4.350 $\pm$ 0.123 1.991 $\pm$ 0.199 1.830 $\pm$ 0.004
+40 Time 35.405 $\pm$ 0.415 0.100 $\pm$ 0.008 0.095 $\pm$ 0.005 0.094 $\pm$ 0.005 0.093 $\pm$ 0.005 0.157 $\pm$ 0.005 0.091 $\pm$ 0.005
+RMSE 1.656 $\pm$ 0.000 1.526 $\pm$ 0.002 1.525 $\pm$ 0.002 3.106 $\pm$ 0.020 2.939 $\pm$ 0.019 1.733 $\pm$ 0.214 1.527 $\pm$ 0.004
+NLPD 1.973 $\pm$ 0.000 2.366 $\pm$ 0.005 2.273 $\pm$ 0.006 7.269 $\pm$ 0.085 6.170 $\pm$ 0.071 2.244 $\pm$ 0.378 1.833 $\pm$ 0.004
+60 Time 35.405 $\pm$ 0.415 0.079 $\pm$ 0.007 0.074 $\pm$ 0.005 0.072 $\pm$ 0.002 0.072 $\pm$ 0.002 0.109 $\pm$ 0.001 0.069 $\pm$ 0.001
+RMSE 1.656 $\pm$ 0.000 1.550 $\pm$ 0.004 1.549 $\pm$ 0.005 3.187 $\pm$ 0.021 3.078 $\pm$ 0.020 1.789 $\pm$ 0.196 1.552 $\pm$ 0.005
+NLPD 1.973 $\pm$ 0.000 2.432 $\pm$ 0.010 2.360 $\pm$ 0.010 7.748 $\pm$ 0.094 6.935 $\pm$ 0.077 2.370 $\pm$ 0.261 1.849 $\pm$ 0.005
+Gas Turbine NOx 20 Time 36.239 $\pm$ 1.322 0.164 $\pm$ 0.003 0.165 $\pm$ 0.003 0.164 $\pm$ 0.003 0.165 $\pm$ 0.003 0.299 $\pm$ 0.008 0.164 $\pm$ 0.003
+RMSE 1.065 $\pm$ 0.041 1.070 $\pm$ 0.012 1.070 $\pm$ 0.012 1.964 $\pm$ 0.035 1.949 $\pm$ 0.032 1.095 $\pm$ 0.083 1.066 $\pm$ 0.012
+NLPD 4.103 $\pm$ 0.446 7.900 $\pm$ 0.171 7.898 $\pm$ 0.176 26.145 $\pm$ 0.892 25.703 $\pm$ 0.802 5.297 $\pm$ 1.285 4.032 $\pm$ 0.100
+40 Time 36.239 $\pm$ 1.322 0.104 $\pm$ 0.007 0.099 $\pm$ 0.006 0.095 $\pm$ 0.004 0.094 $\pm$ 0.004 0.162 $\pm$ 0.008 0.092 $\pm$ 0.004
+RMSE 1.115 $\pm$ 0.000 1.063 $\pm$ 0.008 1.063 $\pm$ 0.008 2.036 $\pm$ 0.017 2.021 $\pm$ 0.016 1.096 $\pm$ 0.057 1.062 $\pm$ 0.013
+NLPD 4.332 $\pm$ 0.000 7.990 $\pm$ 0.126 7.987 $\pm$ 0.122 28.948 $\pm$ 0.476 28.460 $\pm$ 0.443 5.350 $\pm$ 0.713 3.890 $\pm$ 0.101
+60 Time 36.240 $\pm$ 1.322 0.084 $\pm$ 0.003 0.076 $\pm$ 0.005 0.075 $\pm$ 0.004 0.075 $\pm$ 0.003 0.112 $\pm$ 0.003 0.072 $\pm$ 0.002
+RMSE 1.115 $\pm$ 0.000 1.066 $\pm$ 0.007 1.066 $\pm$ 0.007 2.085 $\pm$ 0.020 2.069 $\pm$ 0.020 1.043 $\pm$ 0.102 1.065 $\pm$ 0.007
+NLPD 4.332 $\pm$ 0.000 8.104 $\pm$ 0.112 8.100 $\pm$ 0.111 30.656 $\pm$ 0.594 30.121 $\pm$ 0.565 5.759 $\pm$ 1.683 3.860 $\pm$ 0.053
+Beijing PM2.5 20 Time 60.044 $\pm$ 1.202 0.242 $\pm$ 0.007 0.240 $\pm$ 0.007 0.240 $\pm$ 0.007 0.240 $\pm$ 0.007 0.422 $\pm$ 0.008 0.239 $\pm$ 0.007
+RMSE 7.780 $\pm$ 0.116 7.776 $\pm$ 0.016 7.775 $\pm$ 0.015 8.530 $\pm$ 0.048 8.363 $\pm$ 0.039 7.755 $\pm$ 0.094 7.775 $\pm$ 0.018
+NLPD 8.640 $\pm$ 0.652 35.010 $\pm$ 0.188 30.668 $\pm$ 0.299 38.362 $\pm$ 0.289 31.270 $\pm$ 0.280 9.147 $\pm$ 0.602 7.350 $\pm$ 0.100
+40 Time 60.044 $\pm$ 1.202 0.138 $\pm$ 0.008 0.127 $\pm$ 0.003 0.125 $\pm$ 0.002 0.125 $\pm$ 0.002 0.228 $\pm$ 0.002 0.124 $\pm$ 0.002
+RMSE 7.655 $\pm$ 0.000 7.830 $\pm$ 0.013 7.829 $\pm$ 0.013 8.823 $\pm$ 0.055 8.682 $\pm$ 0.054 7.786 $\pm$ 0.119 7.831 $\pm$ 0.010
+NLPD 9.020 $\pm$ 0.000 39.237 $\pm$ 0.137 36.061 $\pm$ 0.179 46.846 $\pm$ 0.554 40.320 $\pm$ 0.481 11.787 $\pm$ 2.199 7.232 $\pm$ 0.096
+60 Time 60.044 $\pm$ 1.202 0.123 $\pm$ 0.016 0.095 $\pm$ 0.003 0.094 $\pm$ 0.002 0.095 $\pm$ 0.003 0.159 $\pm$ 0.004 0.094 $\pm$ 0.002
+RMSE 7.655 $\pm$ 0.000 7.871 $\pm$ 0.012 7.869 $\pm$ 0.012 9.041 $\pm$ 0.055 8.917 $\pm$ 0.047 7.942 $\pm$ 0.157 7.873 $\pm$ 0.013
+NLPD 9.020 $\pm$ 0.000 41.091 $\pm$ 0.128 38.502 $\pm$ 0.151 51.722 $\pm$ 0.562 45.666 $\pm$ 0.351 12.596 $\pm$ 1.516 7.140 $\pm$ 0.095
+Metro Traffic 20 Time 57.739 $\pm$ 2.806 0.317 $\pm$ 0.051 0.262 $\pm$ 0.010 0.255 $\pm$ 0.010 0.254 $\pm$ 0.009 0.467 $\pm$ 0.018 0.253 $\pm$ 0.010
+RMSE 6.065 $\pm$ 0.701 5.612 $\pm$ 0.099 5.534 $\pm$ 0.060 58.979 $\pm$ 2.695 34.593 $\pm$ 0.416 6.512 $\pm$ 1.272 5.741 $\pm$ 0.199
+NLPD 7.874 $\pm$ 0.083 11.481 $\pm$ 0.157 7.960 $\pm$ 0.023 235.124 $\pm$ 9.921 30.309 $\pm$ 1.069 7.942 $\pm$ 0.230 7.841 $\pm$ 0.026
+40 Time 57.739 $\pm$ 2.806 0.419 $\pm$ 0.118 0.162 $\pm$ 0.002 0.151 $\pm$ 0.004 0.146 $\pm$ 0.001 0.236 $\pm$ 0.003 0.144 $\pm$ 0.001
+RMSE 6.236 $\pm$ 0.000 5.587 $\pm$ 0.027 5.569 $\pm$ 0.020 49.724 $\pm$ 0.474 41.100 $\pm$ 0.278 6.050 $\pm$ 0.663 5.648 $\pm$ 0.067
+NLPD 7.893 $\pm$ 0.001 16.096 $\pm$ 0.074 9.088 $\pm$ 0.020 354.159 $\pm$ 3.765 76.711 $\pm$ 0.896 7.872 $\pm$ 0.145 7.843 $\pm$ 0.012
+60 Time 57.740 $\pm$ 2.806 0.330 $\pm$ 0.128 0.137 $\pm$ 0.004 0.122 $\pm$ 0.002 0.122 $\pm$ 0.002 0.167 $\pm$ 0.006 0.118 $\pm$ 0.001
+RMSE 6.236 $\pm$ 0.000 5.651 $\pm$ 0.023 5.638 $\pm$ 0.020 44.833 $\pm$ 0.618 40.345 $\pm$ 0.422 5.884 $\pm$ 0.406 5.692 $\pm$ 0.069
+NLPD 7.893 $\pm$ 0.001 20.677 $\pm$ 0.092 10.498 $\pm$ 0.038 433.313 $\pm$ 6.898 113.021 $\pm$ 0.620 7.829 $\pm$ 0.143 7.856 $\pm$ 0.012
+6 Discussions
+Conclusions
 In this work, we introduced an algorithm to enhance GP regressions with inducing points on sparse grids using the optimized combination technique (OptiCom).
@@
 Through extensive experiments and comparisons with state-of-the-art aggregation models, we demonstrated the competitiveness and stability of our proposed method.
-Our algorithm consistently provides stable and accurate predictions, offering a robust solution for distributed GP regression tasks.
-Acknowledgments and Disclosure of Funding
+Our algorithm consistently provides stable and accurate predictions, offering a robust solution for distributed GP regression.
+Practical considerations
+Beyond model design, practical deployment involves key trade-offs in choosing the number of experts and inducing points.
+The choice of the number of experts $M$ involves a trade-off between efficiency and accuracy.
+Increasing $M$ improves computational efficiency, as each expert handles a smaller subset of the data and can be trained in parallel.
+However, if $M$ becomes too large, each expert receives too little data, degrading the quality of the local GP approximations.
+This effect is clearly illustrated in Figure 4.
+In practice, we recommend selecting $M$ based on empirical validation, such as cross-validation.
+A practical set of initial candidates can be defined as $M\in\{\left\lfloor\sqrt{n}/i\right\rfloor:i=4,5,6\}$, where $n$ is size of the global training data.
+Similarly, the number of inducing points affects both accuracy and computational cost.
+More inducing points improve local approximations, and consequently lead to more accurate aggregated predictions, they also increase computational cost.
+In practice, selecting 5–10% of the local training data size as the number of inducing points provides a good balance.
+Limitations
+While our method guarantees consistent predictive means in the limit, it does not ensure consistent predictive variances.
+Theoretically, variance consistency would require the aggregation weights to satisfy an additional $\ell_{2}$-norm condition, which we do not impose.
+As a result, variance estimates may be slightly less accurate compared to methods like grBCM, especially in uncertainty-sensitive applications.
+Future directions
+A direction for future work is to extend our framework such that it also ensures variance consistency, potentially by modifying the weight optimization process or incorporating additional constraints.
+Furthermore, integrating adaptive selection of expert sizes and inducing points could further enhance the scalability and flexibility of the method in large-scale settings.
+Acknowledgements
 The authors acknowledge the generous support from the NSF grants DMS-2312173 and CNS-2328395.
 The experiments of this research were conducted on an A100 GPU provided by Texas A&M High Performance Research Computing.
+Data availability statement
+The data that support the findings of this study are openly available in (Asuncion and Newman, 2007) at https://archive.ics.uci.edu/datasets.
 Appendix A GP with OptiCom
+In this section, we provide supporting figures for optimized combination technique (OptiCom) and detailed algorithms for GP with OptiCom in Section A.1 and Section A.2 respectively.
 A.1 Figures
@@
 2:Output: coefficients $\mathbf{c}=\{c_{\underline{l}^{(i)}}\}_{i=1}^{b}$, kernel weights $\{\bm{\alpha}_{\underline{l}^{(i)}}\}_{i=1}^{b}$, Cholesky factors $\{\mathbf{L}_{\underline{l}^{(i)}}\}_{i=1}^{b}$, where $\{\underline{l}^{(i)}\}_{i=1}^{b}=\{\underline{l}:\eta\leq|\underline{l}|_{1}\leq\eta+d-1\}$
-3:  $\mathbf{A}:=\text{zeros}(b,b)$
 […]
-[…]
+4:for $t=0$ to b1b-1 do # for loop over $t$ and $i$ can be parallelized
+5:  for $i=1$ to btb-t do # loop over the diagonal with offset =t=t
 6:   $j:=i+t$
 7:   $\mathbf{M}=\big[k(\Omega_{\underline{l}^{(i)}},\Omega_{\underline{l}^{(j)}})+\sigma_{\epsilon}^{-2}k(\Omega_{\underline{l}^{(i)}},\mathbf{X})k(\mathbf{X},\Omega_{\underline{l}^{(j)}})\big]$
-[…]
+8:   if $t=0$ then # $i=j$, $\mathbf{M}$ is symmetric and positive definite
 9:    $\mathbf{L}_{\underline{l}^{(i)}}:=\text{chol}(\mathbf{M})$ s.t. $\mathbf{L}_{\underline{l}^{(i)}}\mathbf{L}_{\underline{l}^{(i)}}^{\top}=\mathbf{M}$
@@
 15:end for
-16:  $\mathbf{c}:=\mathbf{A}\,\backslash\,\text{diag}(\mathbf{A})$
+[…]
 17:return $\mathbf{c}$, $\{\bm{\alpha}_{\underline{l}^{(i)}}\}_{i=1}^{b}$, $\{\mathbf{L}_{\underline{l}^{(i)}}\}_{i=1}^{b}$
@@
 2:Output: posterior mean $\hat{\mu}_{\eta}^{c}(\mathbf{X}^{*})$ and posterior covariance $\hat{k}_{\eta}^{c}(\mathbf{X}^{*},\mathbf{X}^{*})$ in eq. 18
-3:  $\mathbf{m}:=\text{zeros}(n_{t})$; $\mathbf{S}:=\text{zeros}(n_{t},n_{t})$
+[…]
 4: get $\mathbf{c}=\{c_{\underline{l}^{(i)}}\}_{i=1}^{b}$, $\{\bm{\alpha}_{\underline{l}^{(i)}}\}_{i=1}^{b}$, $\{\mathbf{L}_{\underline{l}^{(i)}}\}_{i=1}^{b}$ from Algorithm 1
-[…]
+5:for $i=1$ to $b$ do # for loop over $i$ can be parallelized
 6:  $\mathbf{m}\leftarrow\mathbf{m}+c_{\underline{l}^{(i)}}k(\mathbf{X}^{*},\Omega_{\underline{l}^{(i)}})\bm{\alpha}_{\underline{l}^{(i)}}$
@@
 11:end for
-12:  $\hat{\mu}_{\eta}^{c}(\mathbf{X}^{*}):=\sigma_{\epsilon}^{-2}\mathbf{m}$
-13:  $\hat{k}_{\eta}^{c}(\mathbf{X}^{*},\mathbf{X}^{*}):=\mathbf{S}$
+[…]
+[…]
 14:return $\hat{\mu}_{\eta}^{c}(\mathbf{X}^{*})$, $\hat{k}_{\eta}^{c}(\mathbf{X}^{*},\mathbf{X}^{*})$
 Appendix B Distributed GP with optimal weights
-B.1 Algorithms
+In this section, we outline the algorithms of the proposed aggregation model for distributed SVGP and distributed exact GP in Section B.1 and Section B.2 respectively.
+B.1 SVGP
 Algorithm 3 Optimal Weights for Distributed SVGP
@@
 2:Output: optimal weights $\bm{\beta}=\{\beta_{i}\}_{i=1}^{M}$, kernel weights $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, Cholesky factors $\{\mathbf{L}_{i}\}_{i=1}^{M}$
-3:  $\mathbf{X}_{c}:=\{\mathbf{x}_{i}^{(c)}\}_{i=1}^{M}$, $\mathbf{x}_{i}^{(c)}\in\mathbf{X}_{i}$ is randomly selected
-4:  $\mathbf{M}^{(c)}=\big{[}k(\mathbf{Z},\mathbf{Z})+\sigma_{\epsilon}^{-2}k(\mathbf{Z},\mathbf{X}_{c})k(\mathbf{X}_{c},\mathbf{Z})\big{]}$
-5:  $\mathbf{A}:=\text{zeros}(M,M)$
 […]
 […]
+[…]
+6:for $t=0$ to M1M-1 do # for loop over $t$ and $i$ can be parallelized
+7:  for $i=1$ to MtM-t do # loop over the diagonal with offset =t=t
 8:   $j:=i+t$
-[…]
+9:   if $t=0$ then # $i=j$
 10:    $\mathbf{M}=\big[k(\mathbf{Z},\mathbf{Z})+\sigma_{\epsilon}^{-2}k(\mathbf{Z},\mathbf{X}_{i})k(\mathbf{X}_{i},\mathbf{Z})\big]$
@@
 17:end for
-18:  $\bm{\beta}:=\mathbf{A}\,\backslash\,\text{diag}(\mathbf{A})$
+[…]
 19:return $\bm{\beta}$, $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, $\{\mathbf{L}_{i}\}_{i=1}^{M}$
@@
 3: get $\bm{\beta}=(\beta_{1},\ldots,\beta_{M})^{\top}$, $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, $\{\mathbf{L}_{i}\}_{i=1}^{M}$ from Algorithm 3
-4:  $\mathbf{m}:=\text{zeros}(n_{t})$; $\mathbf{s}:=\text{zeros}(n_{t})$
 […]
+5:for $i=1$ to $M$ do # for loop over $i$ can be parallelized
 6:  $\mathbf{m}\leftarrow\mathbf{m}+\beta_{i}k(\mathbf{X}^{*},\mathbf{Z})\bm{\alpha}_{i}$
-7:     $\mathbf{R}:=\mathbf{L}_{i}\backslash K(\mathbf{Z},\mathbf{X}^{*})$
+[…]
 8:  $\mathbf{s}\leftarrow\mathbf{s}+\beta_{i}^{2}\mathbf{R}^{\top}\mathbf{R}$
 9:end for
-10:  $\mu_{\mathcal{A}}(\mathbf{X}^{*}):=\sigma_{\epsilon}^{-2}\mathbf{m}$
-11:  $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*}):=\mathbf{s}$
+[…]
+[…]
 12:return $\mu_{\mathcal{A}}(\mathbf{X}^{*})$, $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*})$
+B.2 Exact GP
 Algorithm 5 Optimal Weights for Distributed Exact GP
@@
 2:Output: optimal weights $\bm{\beta}=\{\beta_{i}\}_{i=1}^{M}$, kernel weights $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, Cholesky factors $\{\mathbf{L}_{i}\}_{i=1}^{M}$
-3:  $\mathbf{X}_{c}:=\{\mathbf{x}_{i}^{(c)}\}_{i=1}^{M}$, $\mathbf{x}_{i}^{(c)}\in\mathbf{X}_{i}$ is randomly selected
-4:  $\mathbf{A}:=\text{zeros}(M,M)$
 […]
 […]
+5:for $t=0$ to M1M-1 do # for loop over $t$ and $i$ can be parallelized
+6:  for $i=1$ to MtM-t do # loop over the diagonal with offset =t=t
 7:   $j:=i+t$
 8:   $\mathbf{M}^{(c)}=\big[k(\mathbf{X}_{i},\mathbf{X}_{j})+k(\mathbf{X}_{i},\mathbf{X}_{c})k(\mathbf{X}_{c},\mathbf{X}_{i})\big]$
-[…]
+9:   if $t=0$ then # $i=j$
 10:    $\mathbf{L}_{i}:=\text{chol}(\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}})$ s.t.
@@
 16:end for
-17:  $\bm{\beta}:=\mathbf{A}\,\backslash\,\text{diag}(\mathbf{A})$
+[…]
 18:return $\bm{\beta}$, $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, $\{\mathbf{L}_{i}\}_{i=1}^{M}$
@@
 3: get $\bm{\beta}=(\beta_{1},\ldots,\beta_{M})^{\top}$, $\{\bm{\alpha}_{i}\}_{i=1}^{M}$, $\{\mathbf{L}_{i}\}_{i=1}^{M}$ from Algorithm 5
-4:  $\mathbf{m}:=\text{zeros}(n_{t})$; $\mathbf{s}:=\text{zeros}(n_{t})$
 […]
+5:for $i=1$ to $M$ do # for loop over $i$ can be parallelized
 6:  $\mathbf{m}\leftarrow\mathbf{m}+\beta_{i}k(\mathbf{X}^{*},\mathbf{X}_{i})\bm{\alpha}_{i}$
-7:     $\mathbf{R}:=\mathbf{L}_{i}\backslash k(\mathbf{X}_{i},\mathbf{X}^{*})$
+[…]
 8:  $\mathbf{s}\leftarrow\mathbf{s}+\beta_{i}^{2}\mathbf{R}^{\top}\mathbf{R}$
 9:end for
-10:  $\mu_{\mathcal{A}}(\mathbf{X}^{*}):=\mathbf{m}$
-11:  $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*}):=\mathbf{s}$
+[…]
+[…]
 12:return $\mu_{\mathcal{A}}(\mathbf{X}^{*})$, $\sigma_{\mathcal{A}}^{2}(\mathbf{X}^{*})$
 Appendix C Experiment
+In this section, we provide the supporting boxplots for the distributed training on synthetic data and tables for the aggregation models on UCI datasets in Section C.1.
 C.1 Boxplots of the hyperparameter estimates
-Figure 8: Boxplots of the optimized hyperparemeter estimates on the $n=10^{4}$ dataset.
-C.2 Boxplots of the metrics
-Figure 9: RMSE of the aggregation models for Exact GP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$.
-Blue dotted line represents the mean, red solid line represents the median.
-Figure 10: NLPD of the aggregation models for Exact GP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$.
-Blue dotted line represents the mean, red solid line represents the median.
-Figure 11: RMSE of the aggregation models for SVGP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, $m=128$ inducing points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$.
-Blue dotted line represents the mean, red solid line represents the median.
-Figure 12: NLPD of the aggregation models for SVGP with the RBF kernel in dimension $d=2$ on $n=10^{4}$ training points, $m=128$ inducing points, and $n_{t}=2500$ test points over the interval $[-1,1]^{2}$.
-Blue dotted line represents the mean, red solid line represents the median.
-C.3 Tables of the UCI datasets
-Table 2: Results of the aggregation models for SVGP on the UCI datasets of size $\sim$15K. Best and worst performances are highlighted in yellow and red, respectively.
-Data $M$ Metric SVGP PoE gPoE BCM rBCM grBCM opt (ours)
-PoleTele 4 Time 1.588 $\pm$ 0.044 3.939 $\pm$ 0.099 3.939 $\pm$ 0.099 3.939 $\pm$ 0.099 3.939 $\pm$ 0.099 3.910 $\pm$ 0.091 3.939 $\pm$ 0.099
-RMSE 16.416 $\pm$ 1.032 41.878 $\pm$ 0.315 41.878 $\pm$ 0.315 44.255 $\pm$ 0.877 41.709 $\pm$ 0.257 16.772 $\pm$ 2.289 35.596 $\pm$ 0.267
-NLPD 22.096 $\pm$ 2.964 17.859 $\pm$ 0.245 7.403 $\pm$ 0.061 7.845 $\pm$ 0.174 7.374 $\pm$ 0.050 65.186 $\pm$ 17.279 6.293 $\pm$ 0.052
-10 Time 1.588 $\pm$ 0.044 2.121 $\pm$ 0.005 2.121 $\pm$ 0.005 2.121 $\pm$ 0.005 2.121 $\pm$ 0.005 2.826 $\pm$ 0.008 2.121 $\pm$ 0.005
-RMSE 48.678 $\pm$ 10.933 41.739 $\pm$ 0.277 41.739 $\pm$ 0.277 44.374 $\pm$ 1.066 41.709 $\pm$ 0.257 17.031 $\pm$ 1.645 30.156 $\pm$ 0.200
-NLPD 189.683 $\pm$ 58.036 39.704 $\pm$ 0.529 7.455 $\pm$ 0.054 7.956 $\pm$ 0.210 7.451 $\pm$ 0.050 127.538 $\pm$ 24.461 5.387 $\pm$ 0.039
-20 Time 1.588 $\pm$ 0.044 1.104 $\pm$ 0.018 1.104 $\pm$ 0.018 1.104 $\pm$ 0.018 1.104 $\pm$ 0.018 1.944 $\pm$ 0.031 1.104 $\pm$ 0.018
-RMSE 48.678 $\pm$ 10.933 41.713 $\pm$ 0.263 41.713 $\pm$ 0.263 42.766 $\pm$ 1.073 41.709 $\pm$ 0.257 17.065 $\pm$ 2.667 25.617 $\pm$ 0.161
-NLPD 189.683 $\pm$ 58.036 122.149 $\pm$ 1.430 9.972 $\pm$ 0.079 10.315 $\pm$ 0.347 9.972 $\pm$ 0.077 174.421 $\pm$ 54.154 6.124 $\pm$ 0.048
-Elevators 4 Time 1.616 $\pm$ 0.036 4.893 $\pm$ 0.025 4.893 $\pm$ 0.025 4.893 $\pm$ 0.025 4.893 $\pm$ 0.025 4.551 $\pm$ 0.012 4.893 $\pm$ 0.025
-RMSE 0.249 $\pm$ 0.004 0.253 $\pm$ 0.005 0.253 $\pm$ 0.005 0.254 $\pm$ 0.005 0.253 $\pm$ 0.005 0.255 $\pm$ 0.005 0.253 $\pm$ 0.005
-NLPD 0.846 $\pm$ 0.002 0.792 $\pm$ 0.002 0.822 $\pm$ 0.002 0.822 $\pm$ 0.002 0.822 $\pm$ 0.002 0.815 $\pm$ 0.002 0.822 $\pm$ 0.002
-10 Time 1.616 $\pm$ 0.036 2.263 $\pm$ 0.008 2.263 $\pm$ 0.008 2.263 $\pm$ 0.008 2.263 $\pm$ 0.008 3.058 $\pm$ 0.008 2.263 $\pm$ 0.008
-RMSE 0.255 $\pm$ 0.006 0.253 $\pm$ 0.005 0.253 $\pm$ 0.005 0.324 $\pm$ 0.040 0.253 $\pm$ 0.005 0.259 $\pm$ 0.004 0.253 $\pm$ 0.005
-NLPD 0.847 $\pm$ 0.002 0.786 $\pm$ 0.002 0.822 $\pm$ 0.002 0.851 $\pm$ 0.018 0.822 $\pm$ 0.002 0.805 $\pm$ 0.002 0.822 $\pm$ 0.002
-20 Time 1.616 $\pm$ 0.036 1.171 $\pm$ 0.008 1.171 $\pm$ 0.008 1.171 $\pm$ 0.008 1.171 $\pm$ 0.008 2.057 $\pm$ 0.011 1.171 $\pm$ 0.008
-RMSE 0.255 $\pm$ 0.006 0.254 $\pm$ 0.005 0.254 $\pm$ 0.005 0.554 $\pm$ 0.070 0.253 $\pm$ 0.005 0.261 $\pm$ 0.005 0.254 $\pm$ 0.005
-NLPD 0.847 $\pm$ 0.002 0.784 $\pm$ 0.002 0.822 $\pm$ 0.001 0.986 $\pm$ 0.053 0.821 $\pm$ 0.001 0.799 $\pm$ 0.002 0.822 $\pm$ 0.001
-Table 3: Results of the aggregation models for SVGP on the UCI datasets of size $\sim$45K. Best and worst performances are highlighted in yellow and red, respectively.
-Data $M$ Metric SVGP PoE gPoE BCM rBCM grBCM opt (ours)
-Kin40k 8 Time 1.821 ± 0.192 5.077 ± 0.110 5.077 ± 0.110 5.077 ± 0.110 5.077 ± 0.110 7.513 ± 0.158 5.077 ± 0.110
-RMSE 0.391 ± 0.030 1.006 ± 0.007 1.006 ± 0.007 1.510 ± 0.128 0.997 ± 0.007 0.471 ± 0.016 0.906 ± 0.007
-NLPD 0.979 ± 0.014 1.440 ± 0.009 1.469 ± 0.005 1.873 ± 0.120 1.463 ± 0.005 0.971 ± 0.011 1.322 ± 0.004
-16 Time 1.821 ± 0.192 2.559 ± 0.014 2.559 ± 0.014 2.559 ± 0.014 2.559 ± 0.014 4.425 ± 0.019 2.559 ± 0.014
-RMSE 1.233 ± 0.300 1.008 ± 0.009 1.008 ± 0.009 2.639 ± 0.384 0.997 ± 0.007 0.483 ± 0.010 0.816 ± 0.008
-NLPD 1.735 ± 0.269 1.455 ± 0.013 1.452 ± 0.006 3.591 ± 0.765 1.444 ± 0.004 0.929 ± 0.007 1.176 ± 0.005
-64 Time 1.821 ± 0.192 1.062 ± 0.012 1.062 ± 0.012 1.062 ± 0.012 1.062 ± 0.012 1.305 ± 0.009 1.062 ± 0.012
-RMSE 1.233 ± 0.300 0.997 ± 0.007 0.997 ± 0.007 1.902 ± 0.554 0.997 ± 0.007 0.611 ± 0.038 0.727 ± 0.005
-NLPD 1.735 ± 0.269 1.448 ± 0.010 1.469 ± 0.005 2.351 ± 0.715 1.469 ± 0.005 1.011 ± 0.033 1.071 ± 0.003
-KEGG 8 Time 1.818 ± 0.256 6.211 ± 0.068 6.211 ± 0.068 6.211 ± 0.068 6.211 ± 0.068 9.227 ± 0.085 6.211 ± 0.068
-RMSE 1.152 ± 0.162 1.408 ± 0.034 1.408 ± 0.034 2.527 ± 0.910 1.378 ± 0.008 1.096 ± 0.218 1.267 ± 0.031
-NLPD 1.572 ± 0.151 1.969 ± 0.061 1.769 ± 0.021 2.900 ± 1.259 1.750 ± 0.008 1.567 ± 0.280 1.592 ± 0.019
-16 Time 1.818 ± 0.256 3.449 ± 0.041 3.449 ± 0.041 3.449 ± 0.041 3.449 ± 0.041 5.407 ± 0.039 3.449 ± 0.041
-RMSE 1.401 ± 0.062 1.455 ± 0.043 1.455 ± 0.043 7.284 ± 2.053 1.378 ± 0.008 1.355 ± 0.383 1.178 ± 0.035
-NLPD 1.841 ± 0.071 2.147 ± 0.087 1.797 ± 0.027 14.708 ± 7.131 1.745 ± 0.004 2.072 ± 0.735 1.455 ± 0.022
-64 Time 1.818 ± 0.256 1.095 ± 0.011 1.095 ± 0.011 1.095 ± 0.011 1.095 ± 0.011 2.075 ± 0.015 1.095 ± 0.011
-RMSE 1.401 ± 0.062 1.386 ± 0.010 1.386 ± 0.010 9.475 ± 2.704 1.378 ± 0.008 1.250 ± 0.210 1.011 ± 0.007
-NLPD 1.841 ± 0.071 2.088 ± 0.021 1.762 ± 0.004 21.055 ± 11.376 1.757 ± 0.006 1.882 ± 0.346 1.285 ± 0.003
-Protein 8 Time 1.876 ± 0.181 6.113 ± 0.103 6.113 ± 0.103 6.113 ± 0.103 6.113 ± 0.103 8.637 ± 0.136 6.113 ± 0.103
-RMSE 0.773 ± 0.003 0.773 ± 0.003 0.773 ± 0.003 0.783 ± 0.013 0.773 ± 0.003 0.773 ± 0.003 0.773 ± 0.003
-NLPD 1.222 ± 0.002 1.176 ± 0.003 1.279 ± 0.002 1.285 ± 0.008 1.279 ± 0.002 1.206 ± 0.004 1.279 ± 0.002
-16 Time 1.876 ± 0.181 3.484 ± 0.034 3.484 ± 0.034 3.484 ± 0.034 3.484 ± 0.034 5.427 ± 0.052 3.484 ± 0.034
-RMSE 0.774 ± 0.003 0.773 ± 0.003 0.773 ± 0.003 0.795 ± 0.022 0.773 ± 0.003 0.774 ± 0.003 0.773 ± 0.003
-NLPD 1.222 ± 0.002 1.171 ± 0.003 1.279 ± 0.002 1.293 ± 0.013 1.279 ± 0.002 1.188 ± 0.005 1.279 ± 0.002
-64 Time 1.876 ± 0.181 1.120 ± 0.024 1.120 ± 0.024 1.120 ± 0.024 1.120 ± 0.024 2.155 ± 0.060 1.120 ± 0.024
-RMSE 0.774 ± 0.003 0.773 ± 0.003 0.773 ± 0.003 0.893 ± 0.100 0.773 ± 0.003 0.776 ± 0.008 0.773 ± 0.003
-NLPD 1.222 ± 0.002 1.168 ± 0.003 1.278 ± 0.002 1.360 ± 0.070 1.278 ± 0.002 1.175 ± 0.008 1.278 ± 0.002
+Figure 8: Boxplots of the optimized hyperparameter estimates on the $n=10^{4}$ dataset.